### v2

# On the singular loci of higher secant varieties of Veronese embeddings

Abstract.
In this paper we study singular loci of higher secant varieties of the image of the $d$-uple Veronese embedding of projective $n$-space, $v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N={n+d\choose d}-1}$. We call the closure of the union of all the $(k-1)$-planes spanned by $k$ points on $X$ the $k$-th secant of $X$ for a given embedded variety $X\subset{\mathbb{P}}^{N}$ and denote it by $\sigma_{k}(X)$. For the singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, it has been known only for $k\leq 3$. As investigating geometry of moving tangents along subvarieties, we determine the (non-)singularity of so-called the $m$-subsecant locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, consisting of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ with any $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$. These loci with arbitrary $k,d,m,n$ show an interesting trichotomy between generic smoothness, non-trivial singularity and trivial singularity. In many cases, they can be served as a new source for the singular locus of the $k$-th secant other than a trivial one, the $(k-1)$-th secant variety. We also consider the case of the $4$-th secant of $v_{d}({\mathbb{P}}^{n})$ by applying main results and computing conormal space via a certain type of Young flattening. In the end, some generalization and discussions for further development are presented as well.
Key words and phrases:
singular locus of higher secant varieties, the subsecant loci, Veronese embedding, symmetric tensors, identifiability
2010 Mathematics Subject Classification:
14M12, 14N05, 14J60, 15A21, 15A69
The first named author was supported by JSPS KAKENHI Grant Number 22K03236. The second named author was supported by a National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT, no. 2017R1E1A1A03070765 and no. 2021R1F1A104818611).
1. Introduction
Throughout the paper, we work over ${\mathbb{C}}$, the field of complex numbers. Let $X\subset{\mathbb{P}}^{N}$ be an embedded projective variety. The $k$-th secant variety of $X$ is defined as:
$\sigma_{k}(X)=\overline{\bigcup_{x_{1},\dots,x_{k}\in X}\langle\,x_{1},\dots,x_{k}\,\rangle}\subset{\mathbb{P}}^{N}~{},$
where $\langle\,x_{1},\dots,x_{k}\,\rangle\subset{\mathbb{P}}^{N}$ denotes the linear span of the points $x_{1},\dots,x_{k}$ and the overline means the Zariski closure. In particular, $\sigma_{1}(X)=X$ and $\sigma_{2}(X)$ has been often called just the secant or secant line variety of $X$ in the literature.
The secant variety construction (more generally, join construction of subvarieties) is not only one of the most famous methods in classical algebraic geometry, but also a very popular object to be considered in recent years, especially in connection with fields of research such as tensor rank and decomposition, algebraic statistics, data science, geometric complexity theory and so on (see [21, 22] for more details).
Despite of a rather long history and its popularity, most of fundamental questions on the higher secant varieties $\sigma_{k}(X)$ still remain open even for many well-known base variety $X$. For instance, the dimension theorem for $\sigma_{k}(X)$ has been only settled completely for $X$ being curves, Veronese varieties or few other cases. Questions about defining equations of $k$-secant variety also have been answered only for very small $k$’s of few cases and seem far from understanding the essence of sources for the equations (see also [21, chapter 5] for a reference).
In this paper we concentrate on the case of $X=v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$, the Veronese variety, which is the image of the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\hookrightarrow{\mathbb{P}}^{N}$ with $N={n+d\choose d}-1$. In particular, we study the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, arbitrary higher secant variety of the Veronese variety.
Knowledge on singularities of higher secant varieties is fundamental and very important for its own sake in the study of algebraic geometry and also can be useful for problems in applications. For example, this can be served as a key condition to determine identifiability of structured tensors (see e.g. the introduction in [8] and references therein).
For any irreducible variety $X\subset{\mathbb{P}}^{N}$, it is classically well-known that
(1) $\sigma_{k-1}(X)\subset\operatorname{Sing}(\sigma_{k}(X))$
unless $\sigma_{k}(X)$ fills the whole linear span $\langle\,X\,\rangle$. In the paper, we call a point $p\in\sigma_{k}(X)$ a non-trivial singular point if $p\notin\sigma_{k-1}(X)$ and $\sigma_{k}(X)$ is singular at $p$, while the points belonging to $\sigma_{k-1}(X)$ are called trivial singular points.
There have been some known results on the singular loci of $\sigma_{k}(X)$ mostly for very small $k$’s. The equality in (1) holds for determinantal varieties defined by minors of a generic matrix. It is completely described for the secant variety of the Segre product of projective spaces (i.e., $k=2$) in [27]. It has been recently generalized to the case of $\sigma_{2}(X)$ where $X$ is a Segre-Veronese embedding by [2] and $X$ is a Grassmannian by [13]. For the $3$-th secant variety of Grassmannian $\mathbb{G}(2,6)$, dimension and some other properties of the singular loci was studied in [4].
For the case of Veronese varieties, it is classical that $\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ holds both for the binary case (i.e., $n=1$) and for the case of quadratic forms (i.e., $d=2$) (see e.g. [19, chapter 1]). When $k=2$, it was proved that the equality above (i.e., having only trivial singular locus) is also true for any $d,n$ by [20]. For $k=3$, by the second author the singular locus was completely determined in [16]. In particular, it was shown that a non-trivial singularity can happen and it is only when $d=4$ and $n\geq 3$.
In the present paper, we explore on the singular loci of any higher secant varieties of Veronese varieties and introduce another main origin for the singularity of $k$-secant variety other than the trivial singular locus. We call this the ‘subsecant locus’ and it shows an interesting trichotomy phenomenon among generic smoothness, non-trivial singular points and trivial singularity as in our main results.
For any given point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, as considering $(1,d-1)$-symmetric flattening (see §3 for details), it is immediate to see that there exists an $m$-plane (i.e., $m$-dimensional linear subvariety) ${\mathbb{P}}^{m}$ of ${\mathbb{P}}^{n}$ with $11\leq m\leq k-$ 1 such that $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$. So, from now on, we call $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ an $m$-subsecant variety of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ (or simply a subsecant variety in case of no confusion) if $m<k-1$ and $m<n$. We also define
(2) $\Sigma_{k,d}(m)=\Sigma_{k,d}(m;{{\mathbb{P}}^{n}}):=\bigcup_{\forall{\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}}\sigma_{k}(v_{d}({\mathbb{P}}^{m})),$
the $m$-subsecant locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. It naturally forms an increasing sequence of loci in the $k$-secant as
$\Sigma_{k,d}(1)\subset\Sigma_{k,d}(2)\subset\cdots\subset\Sigma_{k,d}(\min\{k-1,n\}-1)\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))=\Sigma_{k,d}(\min\{k-1,n\})~{}.$
We call $\Sigma_{k,d}(\min\{k-1,n\}-1)$ (i.e., the union of all proper subsecant varieties) the maximum subsecant locus of the given $k$-th secant $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Any point outside the maximum subsecant locus and the previous secant variety (i.e., $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\big{(}\Sigma_{k,d}(\min\{k-1,n\}-1)\cup\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))\big{)}$) is called a point of full-secant locus. Note that in the case of $k=3$ of [16], the only exceptional case for $\operatorname{Sing}(\sigma_{3}(v_{d}({\mathbb{P}}^{n})))$ (i.e., $d=4$ and $n\geq 3$), the singular locus is given as the maximum subsecant locus $\Sigma_{3,d}(1)$, while the locus is just the trivial one, $\sigma_{2}(v_{d}({\mathbb{P}}^{n}))$, for all the other $d,n$ (see Remark 4 (a)). Most of previously known results on singular loci of secant varieties can be understood in this viewpoint (see Remark 37).
Thus, a basic question for our concern could be stated as follows: For given $k,d,m,n$,
When does an $m$-subsecant locus $\Sigma_{k,d}(m)$ become singular in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$?
In principle, it is in some sense (in spite of its computational complexity) straightforward to check the singularity, once a complete set of equations for a higher secant variety is attained. But, as mentioned above, not much has been known on the defining equations and it seems far from being understood completely at this moment, even for the Veronese case (see [23, 15] for the state of the art). Lacking in knowledge on the equations for higher secant varieties, it is very difficult to determine singular locus in general.
In this paper, without further understanding on equations (!), we introduce a geometric way to pursue it for this kind of problems, which is based on a careful study on behavior of embedded tangent spaces moving along a locus in the Veronese variety . For $m=1$ case, we first present the following result for the $1$-subsecant locus $\Sigma_{k,d}(1)=\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which is a generalization of [16, theorem 2.1] (i.e., $k=3$ case) to any higher $k$-secant of Veronese varieties.
Theorem 1.
Let $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ be the $d$-uple Veronese embedding with $n\geq 2$, $d\geq 3$, and $N={n+d\choose d}-1$. Assume $k\geq 3$. For $(k,d,n)\neq(3,4,2)$, the following holds:
  1. (i)
    If $k\leq\frac{d+1}{2}$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at every point in $\Sigma_{k,d}(1)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
  2. (ii)
    If $k=\frac{d+2}{2}$ , then $\Sigma_{k,d}(1)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(1)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity). This case happens only if $d$ is even.
  3. (iii)
    If $k\geq\frac{d+3}{2}$, then $\Sigma_{k,d}(1)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
For $(k,d,n)=(3,4,2)$, it holds that
  1. (iv)
    $\sigma_{3}(v_{4}({\mathbb{P}}^{2}))$ is smooth at every point in $\Sigma_{3,4}(1)\setminus\sigma_{2}(v_{4}({\mathbb{P}}^{2}))$.
Concerning singular points of arbitrary $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ originated from subsecant loci, we prove the following general theorems on the $m$-subsecant locus $\Sigma_{k,d}(m)$ with $m\geq 2$ and $k\geq 4$ as our main results.
Theorem 2.
Let $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ be the $d$-uple Veronese embedding with $n\geq 3$, $d\geq 3$, and $N={n+d\choose d}-1$. Let $k\geq 4$ and let ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ be an $m$-plane with $2\leq m<\min\{k-1,n\}$. Assume that $(d,m)\notin\{(3,4),(4,2),(4,3),(4,4)\}$.
For $(k,d,n)\neq(4,3,3)$, setting $\mu$ to be ($\frac{{m+d\choose m}}{m+1}$ if $(d,m)\notin\{(6,2),(3,5)\}$ and to be ($\frac{{m+d\choose m}}{m+1}-1$ if $(d,m)\in\{(6,2),(3,5)\}$, we have:
  1. (i)
    If $k\leq\mu$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
  2. (ii)
    If $\mu<k<\mu+1$, then $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity). In particular, this singularity appears if and only if $\mu\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$.
  3. (iii)
    If $k\geq\mu+1$, then $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
For $(k,d,n)=(4,3,3)$, it holds that
  1. (iv)
    $\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ is smooth at every point in $\Sigma_{4,3}(2)\setminus\sigma_{3}(v_{3}({\mathbb{P}}^{3}))$.
For the cases of $(d,m)=(3,4),(4,2),(4,3)$ or $(4,4)$, only in which the dimension of a $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ can be defective, we have the following statement.
Theorem 3.
In the same situation as Theorem 2, for $(d,m)\in\{(3,4),(4,2),(4,3),(4,4)\}$ and $\mu=\frac{{m+d\choose m}}{m+1}$, the following holds:
  1. (i)
    If $k\leq\lfloor\mu\rfloor-1$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
  2. (ii)
    If $\lfloor\mu\rfloor\leq k\leq\lceil\mu\rceil+1$, then $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
  3. (iii)
    If $k\geq\lceil\mu\rceil+2$, then $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
We summarize Theorem 3 in Table 1 with exact ranges of $k$.
$(d,m)$ $\Sigma_{k,d}(m)$ : Generically smooth (i) Non-trivial singular locus (ii) Trivial singularity (iii)
$(3,4)$ $k\leq 6$ $k=7,8$ $9\leq k$
$(4,2)$ $k\leq 4$ $k=5,6$ $7\leq k$
$(4,3)$ $k\leq 7$ $k=8,9,10$ $11\leq k$
$(4,4)$ $k\leq 13$ $k=14,15$ $16\leq k$
Table 1. (Non-)singularity of $\Sigma_{k,d}(m)$ in case of defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s in Theorem 3
Remark 4.
We’d like to make some remarks on the theorems above.
  1. (a)
    In the case of $k=3$ and $n\geq 3$, the condition of (ii) in Theorem 1 holds if and only if $d=4$, and then $\operatorname{Sing}(\sigma_{3}(v_{4}({\mathbb{P}}^{n}))$ is equal to $\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle=\Sigma_{3,4}(1)$ since $\sigma_{3}(v_{4}({\mathbb{P}}^{1}))=\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle$ and $\sigma_{2}(v_{4}({\mathbb{P}}^{n}))\subset\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle$ (see also Corollary 32). This gives a geometric description of the only exceptional case for the singular loci of $3$rd secant variety in [16].
  2. (b)
    (i) of Theorem 1 is stronger than (i) of Theorems 2 and 3, since it claims smoothness for ‘every point’ in the $m$-subsecant locus $\Sigma_{k,d}(m)$ outside $\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$. The ‘general point’ condition of (i) in Theorems 2 and 3 can not be deleted (see Example 27). Except this aspect, (i)-(iii) of Theorem 1 can be also derived as applying Theorem 2 directly to the case $m=1$ (recall that curve case has no defectivity in its higher secants).
  3. (c)
    (i) in both Theorem 2 and Theorem 3 mostly corresponds to the identifiable case of a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$, which means that there is a unique $k$-tuple of points $x_{1},\dots,x_{k}\in v_{d}({\mathbb{P}}^{m})$ such that $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$. In the viewpoint of secant fiber in the incidence (3) in §2, it means that the fiber $p^{-1}(a)$ under $p$ consists of only one element up to permuting $x_{i}$’s. On the other hand, (ii) of Theorems 2 and 3 corresponds to the case of generic non-identifiability of the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ and to the situation of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq\langle v_{d}({\mathbb{P}}^{m})\rangle$. This non-identifiability can happen if $\dim(p^{-1}(a))>0$, or if #$\#p^{-1}(a)\geq 2$ (modulo permutation) in case of $\dim p^{-1}(a)=0$; the former one corresponds to $(d,m)\notin\{(6,2),(3,5),(4,3)\}$ and the latter $(d,m)\in\{(6,2),(3,5),(4,3)\}$.
  4. (d)
    Theorem 3 deals with the case in which the given subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ can possibly be defective due to the famous theorem of Alexander-Hirschowitz [1], whereas Theorem 2 concerns the case of always non-defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s.
  5. (e)
    Two cases $(k,d,m,n)=(3,4,1,2),(4,3,2,3)$ (i.e., (iv) of Theorems 1 and  2), a posteriori, turn out to be the only two exceptional cases which do not follow this trichotomy pattern; in other words, though $k$ belongs to the range of (ii), $\Sigma_{k,d}(m)$ does not provide non-trivial singularity (see also Remark 24). In [15], we in fact show that if $(k,d,n)=(3,4,2),(4,3,3)$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is a del Pezzo $k$-secant variety, that is, a $k$-secant variety of next-to-minimal degree. In this sense, these two cases also belong to a special class with respect to degrees of higher secant varieties. (For basic definitions and results on such varieties, see [11], [10].)
As an application of our main results for $\Sigma_{4,d}(2)$ and $\Sigma_{4,d}(1)$, we obtain the following result on the singular locus of the fourth secant variety of any Veronese embeddings.
Theorem 5 (Singular locus for $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$).
Let $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ be the $d$-uple Veronese embedding with $n\geq 3$, $d\geq 3$, and $N={n+d\choose d}-1$. Then the followings hold.
  1. (i)
    $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is smooth at every point outside $\Sigma_{4,d}(2)\cup\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$.
  2. (ii)
    If $d\geq 4$, then a general point in $\Sigma_{4,d}(2)\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is also a smooth point of $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$. When $d=3$ and $n=3$, all points in $\Sigma_{4,d}(2)\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ are smooth. If $d=3$ and $n\geq 4$, then $\Sigma_{4,d}(2)\subset\operatorname{Sing}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{4,d}(2)\not\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity).
  3. (iii)
    For $d\geq 7$, $\Sigma_{4,d}(1)\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\subset\operatorname{Sm}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$. When $d=6$, $\Sigma_{4,d}(1)\subset\operatorname{Sing}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{4,d}(1)\not\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity). When $d\leq 5$, $\Sigma_{4,d}(1)\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
Example 6 (Cases with a nice description).
The smallest case for singular locus of $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ beyond the classical results for $d=2$ or $n=1$ is $(d,n)=(3,2)$, but there is nothing to check in this case, because $\sigma_{4}(v_{3}({\mathbb{P}}^{2}))$ fills up the ambient space ${\mathbb{P}}^{9}$ (i.e., $\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{2})))=\emptyset$). In next case $(d,n)=(3,3)$, by Theorem 5 (i) and (ii), we have
$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{3})))=\sigma_{3}(v_{3}({\mathbb{P}}^{3}))~{}.$
For the case $(d,n)=(3,4)$, if $V$ denotes a $5$-dimensional ${\mathbb{C}}$-vector space with ${\mathbb{P}}^{4}={\mathbb{P}}V$, Theorem 5 tells us that the singular locus of the $4$-th secant of $v_{3}({\mathbb{P}}V)$ in ${\mathbb{P}}^{34}$ is precisely the locus of cubic hypersurfaces in 5 variables which are cones with the vertex dimension $\geq 1$ as follows:
$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}V)))=\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\Sigma_{4,3}(2;{{\mathbb{P}}V})=\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\sigma_{4}(v_{3}({\mathbb{P}}^{2}))\big{\}}=\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\langle\,v_{3}({\mathbb{P}}^{2})\,\rangle$
=$=\big{\{}f\in{\mathbb{P}}S^{3}V~{}|~{}\textrm{$f$ defines a cubichypersurface in ${\mathbb{P}}V$ which is a cone with $\dim\operatorname{Vertex}(f)\geq 1$}\big{\}}~{},$
which is just the maximum subsecant locus $\Sigma_{4,3}(2;{{\mathbb{P}}V})$, an irreducible $15$-dimensional locus in the $19$-dimensional variety $\sigma_{4}(v_{3}({\mathbb{P}}V))$. In the same argument, we can obtain
$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{n})))=\Sigma_{4,3}(2;{{\mathbb{P}}^{n}})$ for any $n\geq 4$.
This kind of a simple description of the singular locus can be attained in few more cases (see Corollary 32 for details).
The paper is structured as follows. In §2, as preparation, we recall some preliminaries on $k$-th secant varieties and corresponding incidences, and reveal some geometric properties of subsecant varieties in higher secants of Veronese varieties, which are crucial for the proof of main theorems. In §3, as an illustration of the whole picture and our main ideas, we treat the case $m=1$ and prove Theorem 1. In §4 we deal with the general case (i.e., $m\geq 2$) and prove Theorem 2 and Theorem 3 as generalizing the ideas used in the previous section. We’d like to remark that this can be done because the dimension theorem [1] and general identifiability question [9, 12] was settled for the case of Veronese varieties. In §5 we focus on the singular locus of $4$-th secant variety and prove Theorem 5 by dividing the case into two parts; ‘full-secant locus points (i.e., $m=3$)’ treated in Theorem 29 via Young flattening and conormal space computation and ‘the subsecant locus’ by Corollary 30. Finally, we make some generalization and remarks on the materials for further development in §6.
Acknowledgements
The authors would like to give thank to Korea Institute for Advanced Study (KIAS) for inviting them and giving a chance to start the project. We are grateful to Giorgio Ottaviani for a helpful comment on Example 6 and to Luca Chiantini and Jarek Buczyński for related discussions as well. The first author also wishes to express his gratitude to Hajime Kaji and Hiromichi Takagi for useful discussions.
2. Some geometric properties of subsecant varieties
2.1. Projection from the incidence to the secant variety
For a (reduced and irreducible) variety $X$, we denote $X\times\cdots\times X$, the (usual) product of $k$ copies of $X$, by $(X)^{k}$. We denote the $k$-fold symmetric product of $X$, $(X)^{k}/S_{n}$, by $\operatorname{Sym}^{k}(X)$.
For the $d$-uple Veronese embedding $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ with $N={n+d\choose d}-1$, we regard the incidence variety $I=I_{(n)}\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$ to be the Zariski closure of
(3) $I^{0}=I_{(n)}^{0}:=\left\{(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
where we write $x_{i}=v_{d}(x_{i}^{\prime})$ for $x_{i}^{\prime}\in{\mathbb{P}}^{n}$.
[…]
Taking the first projection $p:I\rightarrow{\mathbb{P}}^{N}$, we have $p(I)=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ (see also [28, definition 1.1.3], [29, chapter I, §1]). For any $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, $p^{-1}(a)$ is often called the secant fiber of $a$. Note that $I$ is invariant under permuting factors on $({\mathbb{P}}^{n})^{k}$ from the definition so that both $p$ and $q_{i}$ maps factor through ${\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$. We also have $\dim I={nk+k-1}$ by considering general fibers of $q:I\rightarrow({\mathbb{P}}^{n})^{k}$. For each $1\leq i\leq k$, let
$q_{i}:I\rightarrow({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{n}$
be the composition of $q$ and the projection to the $i$-th factor $({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{n}$.
Remark 7.
We have some remarks on the incidence $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$.
  1. (a)
    $I^{0}$ can be viewed as a ${\mathbb{P}}^{k-1}$-bundle over a non-empty open subset $U$ of $({\mathbb{P}}^{n})^{k}$, consisting of $k$-tuples of points with the expected spanning dimension, so that $I$ is irreducible. Further, $I^{0}=q^{-1}(U)$, since both are irreducible closed subsets in ${\mathbb{P}}^{N}\times U$ having the same dimension. So, for any $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in I\setminus I^{0}$ and for $x_{i}=v_{d}(x_{i}^{\prime})$, it holds that $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ (in other words, there is no $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in I\setminus I^{0}$ such that $\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1$ but $a\notin\langle\,x_{1},\dots,x_{k}\,\rangle$). Finally, note that the Euclidean closure of $I^{0}$ also coincides with $I$ in this case (see e.g. [22, theorem 3.1.6.1]).
  2. (b)
    In the case of $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$, it holds $p(I\setminus I^{0})\neq\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. In addition, if $k\leq\binom{n+d-1}{n}$, then as setting
    $D=\left\{(x_{1}^{\prime},\dots,x_{k}^{\prime})\in({\mathbb{P}}^{n})^{k}\mid\dim(\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle)<k-1\right\},$
    we have $\dim(I\cap({\mathbb{P}}^{N}\times D))<\dim(I)$, and hence $p(I\cap({\mathbb{P}}^{N}\times D))\neq\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
  3. (c)
    (Alternative incidences for the $k$-th secant variety) In the literature, instead of $(X)^{k}$, other spaces such as the symmetric product $\operatorname{Sym}^{k}(X)$ (e.g. in [9]), the Hilbert scheme of degree $k$ 00-dimensional subschemes $\operatorname{Hilb}_{k}(X)$ (e.g. in [7]), and the Grassmannian $\mathbb{G}(k-1,N)$ (e.g. in [26]) also have been used in the incidence to consider the $k$-th secant variety of $X\subset{\mathbb{P}}^{N}$.
Remark 8.
Let us fix an $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ and consider the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$, $v_{d}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta}$ with $\beta={m+d\choose m}-1$. We often use the following notation.
  1. (a)
    Changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}:=n-m$, we may assume that a given ${\mathbb{P}}^{m}$ in ${\mathbb{P}}^{n}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$. For any $x_{i}^{\prime}\in{\mathbb{P}}^{n}$, say
    $x_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:x_{i,m+1}^{\prime}:\dots:x_{i,n}^{\prime}]\in{\mathbb{P}}^{n},$
    then we set $y_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:0:\dots:0]$. Thus, $y_{i}^{\prime}$ gives a point of the $m$-plane if $x_{i}^{\prime}$ is not of the form [0::0:::][0$\dots$:0$\ast$:$\dots$:$\ast$][ 0 : … : 0 : ∗ : … : ∗ ]. By abuse of notations, we denote by $y_{i}^{\prime}$ both corresponding points in ${\mathbb{P}}^{m}$ and in ${\mathbb{P}}^{n}$. Further, as considering linear projections $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ (eliminating the $u$-variables) and $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ (eliminating all the monomials of degree $d$ which involve $u$-variables), we have a natural commuting diagram as $v_{d}\circ\pi_{1}=\pi_{2}\circ v_{d}$
[…]
    In particular, when $d=2$ and $n=m+1$ (i.e., $m^{\prime}=1$), then $\pi_{1}$ is an (inner) projection from one point $a\in{\mathbb{P}}^{n}$ and $\pi_{2}$ corresponds to a tangential projection of ${\mathbb{P}}^{N}$ from $\mathbb{T}_{v_{2}(a)}v_{2}({\mathbb{P}}^{n})$.
  2. (b)
    On the affine open subset $\left\{t_{0}\neq 0\right\}$, the $d$-uple Veronese embedding $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized by monomials in $\mathsf{mono}[t,u]_{\leq d}$, where $\mathsf{mono}[t,u]_{\leq e}~{}\textrm{(resp. $\mathsf{mono}[t]_{\leq e}$)}$ is defined to be the set of monomials in ${\mathbb{C}}[t_{1},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}]$ (resp. in ${\mathbb{C}}[t_{1},\dots,t_{m}]$) of degree $\leq e$ for an integer $e$.
Let us study behaviors of some points in the boundary of $I$, which belong to $I$ but do not belong to $I^{0}$, as follows.
Lemma 9.
Let $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ be any $m$-plane with $m<n$, and consider ${\mathbb{P}}^{\beta}=\langle\,v_{d}(L)\,\rangle\subset{\mathbb{P}}^{N}$. For $v_{d}:L={\mathbb{P}}^{m}\hookrightarrow{\mathbb{P}}^{\beta}$, let us take the incidence variety $I_{(m)}\subset{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k}$ of the same kind as $I=I_{(n)}$ in (3). In this setting, for any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ with $a\in{\mathbb{P}}^{\beta}$, one of the following conditions holds:
  1.  (i)
    $a\in p(I_{(m)}\setminus I_{(m)}^{0})$; or
  2.  (ii)
    there is a subset $C\subset({\mathbb{P}}^{m})^{k}$ of dimension $>0$ such that $\left\{a\right\}\times C\subset I_{(m)}$.
Proof.
Let ${\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ be a linear projection whose center does not intersect with $L$. It induces a natural linear projection ${\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$, and then we have
$\rho_{1}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\dashrightarrow{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k},\quad\rho_{2}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\dashrightarrow{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k},$
where $\overline{\rho_{2}(\rho_{1}(I))}=I_{(m)}$. Taking two Segre embeddings
${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}},\quad{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\hookrightarrow{\mathbb{P}}^{l_{2}},$
we may regard $\rho_{1}$ as the restriction of a linear projection ${\mathbb{P}}^{l_{1}}\dashrightarrow{\mathbb{P}}^{l_{2}}$ whose center is a certain linear subvariety $R\subset{\mathbb{P}}^{l_{1}}$.
[…]
Under the embedding $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}}$, taking $R\cap I$ in ${\mathbb{P}}^{l_{1}}$, we have $\operatorname{codim}(R\cap I,I)\geq 2$, as follows. Let $R^{\prime}$ be the center of the projection ${\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$, where $R^{\prime}\cap L=\emptyset$ and $\operatorname{codim}(R^{\prime},{\mathbb{P}}^{n})=m+1$. For $R^{\prime}(i)=\left\{(x_{1}^{\prime},\dots,x_{k}^{\prime})\in({\mathbb{P}}^{n})^{k}\mid x_{i}^{\prime}\in R^{\prime}\right\}$, we have
$R\cap I=q^{-1}\bigl{(}\bigcup_{i=1}^{k}R^{\prime}(i)\bigr{)},$
where $q:I\rightarrow({\mathbb{P}}^{n})^{k}$. In particular, since dimension of fibers of $q|_{I^{0}}$ is constant, it follows that $\operatorname{codim}(R\cap I^{0},I)=m+1$. Next, let $\left\{W_{j}\right\}$ be the irreducible components of $I\setminus I^{0}$, and let $(a_{j},x_{j,1}^{\prime},\dots,x_{j,k}^{\prime})\in W_{j}$. A general $(n-m-1)$-plane ${\mathbb{P}}^{n-m-1}\subset{\mathbb{P}}^{n}$ satisfies $x_{j,i}^{\prime}\notin{\mathbb{P}}^{n-m-1}$. Since $\left\{W_{j}\right\}$ does not depend on the choice of $R^{\prime}={\mathbb{P}}^{n-m-1}$, we may assume $x_{j,i}^{\prime}\notin R^{\prime}$. Then $W_{j}\not\subset R$ for any $j$. Taking an irreducible component $Q$ of $R\cap I$ such that $Q\subset I\setminus I^{0}$, we have some $j$ such that $Q\subset R\cap W_{j}\subsetneq W_{j}\subsetneq I$. Hence $\operatorname{codim}(Q,I)\geq 2$. As a result, $\operatorname{codim}(R\cap I,I)\geq 2$.
Now, let $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ satisfy $a\in{\mathbb{P}}^{\beta}$, where $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ for $x_{i}=v_{d}(x_{i}^{\prime})$ (note that we do not know $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$). We regard $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$ as a point of ${\mathbb{P}}^{l_{1}}$ under the embedding $I\subset{\mathbb{P}}^{l_{1}}$.
If $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\notin R$, then $\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$ is determined in $\overline{\rho_{1}(I)}\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}$, where $y_{i}^{\prime}\in{\mathbb{P}}^{m}$ is the image of $x_{i}^{\prime}$ under ${\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{m}$. Since $a\in{\mathbb{P}}^{\beta}$, $\rho_{2}(\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime}))=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$ is determined and is contained in $I_{(m)}$. Then, $\dim\langle\,y_{1},\dots,y_{k}\,\rangle<k-1$ for $y_{i}=v_{d}(y_{i}^{\prime})\in{\mathbb{P}}^{\beta}$, which means $(a,y_{1}^{\prime},\dots,y_{1}^{\prime})\in I_{(m)}\setminus I_{(m)}^{0}$.
Assume $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in R$. Let ~$\tilde{{\mathbb{P}}}^{l_{1}}\subset{\mathbb{P}}^{l_{1}}\times{\mathbb{P}}^{l_{2}}$ be the graph of ${\mathbb{P}}^{l_{1}}\dashrightarrow{\mathbb{P}}^{l_{2}}$. In other words, $\widetilde{{{\mathbb{P}}}^{l_{1}}}=Bl_{R}({\mathbb{P}}^{l_{1}})$, the blowing-up of ${\mathbb{P}}^{l_{1}}$ with respect to $R$. Let $r_{1}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{1}}$ and $r_{2}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{2}}$ be projections. Note that, for the strict transformation $S\subset\tilde{{\mathbb{P}}}^{l_{1}}$ of ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\subset{\mathbb{P}}^{l_{1}}$, two composite morphisms $S\xrightarrow{r_{1}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}$ and $S\xrightarrow{r_{2}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$ coincide since it holds on an open subset of $S$. Let $E\subset\tilde{{\mathbb{P}}}^{l_{1}}$ be the exceptional divisor, and let $\widetilde{I}\subset\tilde{{\mathbb{P}}}^{l_{1}}$ be the strict transformation of $I\subset{\mathbb{P}}^{l_{1}}$. Then $r_{2}(\widetilde{I})=\overline{\rho_{1}(I)}$.
Let $E_{1}=r_{1}^{-1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\cap(E\cap\widetilde{I})$, the fiber of $E\cap\widetilde{I}\rightarrow R\cap I$ at $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$. It follows from $\operatorname{codim}(R\cap I,I)\geq 2$ that $\dim(E_{1})\geq 1$. Since $r_{1}^{-1}(z)\simeq r_{2}(r_{1}^{-1}(z))\subset{\mathbb{P}}^{l_{2}}$ for each $z\in{\mathbb{P}}^{l_{1}}$, we have $\dim(r_{2}(E_{1}))\geq 1$. Since the image of $r_{2}(E_{1})$ under ${\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$ is $\left\{a\right\}$, there is $C\subset({\mathbb{P}}^{m})^{k}$ of positive dimension such that $r_{2}(E_{1})=\left\{a\right\}\times C\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}$. Since $a\in{\mathbb{P}}^{\beta}$ and $\rho_{2}(r_{2}(E_{1}))\subset\rho_{2}(r_{2}(\tilde{I}))=\overline{\rho_{2}(\rho_{1}(I))}=I_{(m)}$, we have $\left\{a\right\}\times C=\rho_{2}(\left\{a\right\}\times C)\subset I_{(m)}$. ∎
Lemma 10 (Non-triviality of subsecant varieties).
For an $m$-plane $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$, we have
$\sigma_{k}(v_{d}(L))\cap\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))\subset\sigma_{k-1}(v_{d}(L)).$
In particular, $\sigma_{k}(v_{d}(L))\not\subseteq\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ unless $\sigma_{k-1}(v_{d}(L))=\sigma_{k}(v_{d}(L))$.
Proof.
Let $a\in\sigma_{k}(v_{d}(L))\cap\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (note that $a$ can be in the boundary of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$). For a general point $b_{0}\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, we take an irreducible curve $C\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ such that $a,b_{0}\in C$. Let $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ be the linear projection in Remark 8 (a). Since $a\in{\mathbb{P}}^{\beta}$, we have $a=\pi_{2}(a)\in C^{\prime}:=\overline{\pi_{2}(C)}\subset{\mathbb{P}}^{\beta}$.
Since $b_{0}$ is general, for a general point $b\in C$, we have $b\in\langle\,x_{1},\dots,x_{k-1}\,\rangle$ with $x_{1},\dots,x_{k-1}\in v_{d}({\mathbb{P}}^{n})$. Take $x_{i}^{\prime}\in{\mathbb{P}}^{n}$ with $x_{i}=v_{d}(x_{i}^{\prime})$. As in Remark 8 (a), we take $y_{i}^{\prime}\in L$ for each $i=1,\dots,k-1$ and say $y_{i}=v_{d}(y_{i}^{\prime})$. Then $\pi_{2}(x_{i})=y_{i}$. It follows that $\pi_{2}(b)\in\langle\,y_{1},\dots,y_{k-1}\,\rangle\subset\sigma_{k-1}(v_{d}(L))$. As a result, $a\in C^{\prime}\subset\sigma_{k-1}(v_{d}(L))$ and the assertion follows. ∎
Remark 11.
We have some consequences of Lemma 10.
  1. (a)
    (Border Rank Preserving Pair) For any ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ and for any $k,d>0$, by Lemma 10, we can derive
    (4) $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle=\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$
    as a set. For one inclusion is obvious, let us prove $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle\subset\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$. Suppose that it does not hold. Then, there exists a form $f\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle$ with $f\in\sigma_{k_{0}}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ for some $k_{0}>k$. Then, $f\in\sigma_{k_{0}}(v_{d}({\mathbb{P}}^{m}))\cap\sigma_{k_{0}-1}(v_{d}({\mathbb{P}}^{n}))$ so that $f\in\sigma_{k_{0}-1}(v_{d}({\mathbb{P}}^{m}))$ by Lemma 10. Similarly, as repeating the same ‘descent’ argument, we have $f\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$, which is a contradiction. Thus, the equality in (4) is true.
    In other words, for any $d$-th Veronese embedding $X=v_{d}({\mathbb{P}}^{n})$ and the linear span $L=\langle v_{d}({\mathbb{P}}^{m})\rangle$ of any sub-Veronese variety $v_{d}({\mathbb{P}}^{m})$, we showed that $(X,L)$ is a border rank preserving pair for any $k,d>0$ in the terminology of [21, definition 5.7.3.1] (we’d also like to note that [5, theorem 1.1] can imply the same result for any $d\geq k$ in case of ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$).
  2. (b)
    (Every $\Sigma_{k,d}(m)$ is closed) Recall that the symmetric subspace variety $Sub_{m}(S^{d}V)$ ([21, section 7.1.3]) is defined as
    $\{f\in{\mathbb{P}}S^{d}V~{}|~{}\exists~{}W\subset V,~{}\dim W=m+1~{},~{}f\in{\mathbb{P}}S^{d}W\}~{}.$
    Let $n=\dim{\mathbb{P}}V$. For any $m\leq n$, we have $Sub_{m}(S^{d}V)=\bigcup_{{\mathbb{P}}^{m}\subset{\mathbb{P}}V}\langle v_{d}({\mathbb{P}}^{m})\rangle$. Now, we show that $\Sigma_{k,d}(m)$ is the intersection of the whole $k$-th secant $\sigma_{k}(v_{d}({\mathbb{P}}V))$ and the symmetric subspace variety $Sub_{m}(S^{d}V)$ set-theoretically. See that
    $\sigma_{k}(v_{d}({\mathbb{P}}V))\cap Sub_{m}(S^{d}V)$ $=\sigma_{k}(v_{d}({\mathbb{P}}V))\cap\displaystyle\bigcup_{{\mathbb{P}}^{m}\subset{\mathbb{P}}V}\langle v_{d}({\mathbb{P}}^{m})\rangle=\displaystyle\bigcup_{{\mathbb{P}}^{m}\subset{\mathbb{P}}V}\bigg{(}\sigma_{k}(v_{d}({\mathbb{P}}V))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle\bigg{)}$
[…]
    Therefore, we obtain that $\Sigma_{k,d}(m)=\sigma_{k}(v_{d}({\mathbb{P}}V))\cap Sub_{m}(S^{d}V)$ as a set and in particular every $m$-subsecant $\Sigma_{k,d}(m)$ is a Zariski-closed locus in $\sigma_{k}(v_{d}({\mathbb{P}}V))$.
2.2. General secant fiber of a subsecant variety, entry locus, and its Veronese image
For an $m$-dimensional projective variety $Z\subset{\mathbb{P}}^{\beta}$, we take another incidence variety $J\subset{\mathbb{P}}^{\beta}\times(Z)^{k}$ to be the Zariski closure of
(5) $J^{0}:=\left\{(a,x_{1},\dots,x_{k})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
with the projections $p:J\rightarrow\sigma_{k}(Z)\subset{\mathbb{P}}^{\beta}$ and
$q_{i}:J\rightarrow(Z)^{k}\rightarrow Z,$
where $(Z)^{k}\rightarrow Z$ is the projection to the $i$-th factor for $1\leq i\leq k$. Then we have $\dim J=mk+k-1$ for any $k\leq\dim\langle Z\rangle+1$ by considering general fibers of $J\rightarrow(Z)^{k}$.
For $a\in\sigma_{k}(Z)$, the scheme-theoretic image $q_{i}(p^{-1}(a))$ in $Z$ is called the ($k$-th) entry locus of $Z$ with respect to $a$ in the literature. It is known that for a general $a\in\sigma_{k}(Z)$ the locus $q_{i}(p^{-1}(a))$ is equidimensional and moreover, if $Z$ is smooth and in the characteristic 0, then $p^{-1}(a)$ is smooth so that $q_{i}(p^{-1}(a))$ is reduced (see [28, definition 1.4.5]).
Let $Z,X\subset{\mathbb{P}}^{N}$ be projective varieties of dimensions $m,n$. Let $Z\subset X$ and $Z\subset{\mathbb{P}}^{\beta}$, where ${\mathbb{P}}^{\beta}$ is a $\beta$-plane of ${\mathbb{P}}^{N}$ (i.e., $Z$ is degenerate in ${\mathbb{P}}^{N}$). Now, a general point $a\in Z$ does not have to be general in $X$ any longer. If $\beta<km+k-1$, then the projection $p$ has positive dimensional fibers.
We begin with a consequence of Terracini’s lemma in our setting and add two more lemmas concerning ‘the entry locus’ $q_{i}(p^{-1}(a))$.
Lemma 12.
Assume that $\sigma_{k}(Z)\not\subset\operatorname{Sing}(\sigma_{k}(X))$. Let $F$ be an irreducible component of $p^{-1}(a)$ for a general point $a\in\sigma_{k}(Z)$ in the incidence (5). Then, for a general point $x\in q_{i}(F)$ with $1\leq i\leq k$, we have
$\mathbb{T}_{x}(X)\subset\mathbb{T}_{a}(\sigma_{k}(X))~{},$
where $\mathbb{T}_{x}(X)\subset{\mathbb{P}}^{N}$ means the embedded tangent space to $X$ at $x$.
Proof.
Since $\sigma_{k}(Z)\setminus\operatorname{Sing}(\sigma_{k}(X))$ is non-empty open in $\sigma_{k}(Z)$ and $a$ is general, it is a smooth point of $\sigma_{k}(X)$, and also is contained in the $(k-1)$-plane $\langle\,x_{1},\dots,x_{k}\,\rangle$ for general $x_{1},\dots,x_{k}\in Z$ with $x_{i}=x$. Then the assertion follows by Terracini’s lemma (cf. [28, corollary 1.4.2], [29, Chapter II, 1.10]). ∎
Lemma 13.
For a projective variety $Z$, let $\langle\,Z\,\rangle={\mathbb{P}}^{\beta}$ and consider the incidence $J$ as (5). Suppose that the ($k-1$)-secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$. Then, for a general point $(a,x_{1},\dots,x_{k})\in J$, and for any irreducible component $F$ of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$, $q_{i}|_{F}:F\rightarrow q_{i}(F)$ is generically finite.
Proof.
For simplicity, we set $i=1$. First, since $(a,x_{1},\dots,x_{k})$ is a general point of the incidence $J$, we may assume that $a$ is a general point of $\sigma_{k}(Z)$ and $x_{1},\dots,x_{k}$ are general $k$ points on $Z$.
Let $\pi_{x_{1}}:{\mathbb{P}}^{\beta}\dashrightarrow{\mathbb{P}}^{\beta-1}$ be the linear projection from $x_{1}$. Since $Z$ is non-degenerate in ${\mathbb{P}}^{\beta}$ and $\sigma_{k-1}(Z)\neq{\mathbb{P}}^{\beta}$, the map $\pi_{x_{1}}|_{\sigma_{k-1}(Z)}$ is generically finite onto its image (otherwise, a general point $x_{1}$ is contained in $\operatorname{Vertex}(\sigma_{k-1}(Z))$, a linear subvariety of $\sigma_{k-1}(Z)$, a contradiction).
Let $J^{\prime}\subset{\mathbb{P}}^{\beta}\times(Z)^{k-1}$ be the incidence for $(k-1)$-secant of $Z$, i.e., the closure of
$\left\{(b,\tilde{x}_{2},\dots,\tilde{x}_{k})\mid b\in\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle\text{ and }\dim\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle=k-2\right\}.$
Since $\sigma_{k-1}(Z)$ is not secant defective, the first projection $p_{J^{\prime}}:J^{\prime}\rightarrow\sigma_{k-1}(Z)$ is generically finite. Hence the composite map
(6) $\rho=\pi_{x_{1}}\circ p_{J^{\prime}}:J^{\prime}\rightarrow\pi_{x_{1}}(\sigma_{k-1}(Z))$
is generically finite.
Let $J_{1}=q_{1}^{-1}(x_{1})\subset J$. For the projection $q_{2,\dots,k}:J\rightarrow(Z)^{k-1},(a,\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\mapsto(\tilde{x}_{2},\dots,\tilde{x}_{k})$, a general fiber of $q_{2,\dots,k}|_{J_{1}}:J_{1}\rightarrow(Z)^{k-1}$ at $(\tilde{x}_{2},\dots,\tilde{x}_{k})$ is isomorphic to the ($k-1$)-plane $\langle\,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$. In particular, $\pi_{x_{1}}\circ p|_{J_{1}}:J_{1}\rightarrow\pi_{x_{1}}(\sigma_{k-1}(Z))$ is surjective. Let $\alpha\in\pi_{x_{1}}(\sigma_{k-1}(Z))$ be a general point. Then we may consider the element $(a,x_{1},x_{2},\dots,x_{k})$ as a general point in $J_{1}$ with $\alpha=\pi_{x_{1}}(a)$.
Let $F$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},x_{2},\dots,x_{k})$, and suppose that $q_{1}|_{F}$ has fibers of positive dimensions. Then, the fiber of $q_{1}|_{F}$ at $x_{1}\in q_{1}(F)$,
$F\cap J_{1}$
is of positive dimension. For $(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}$, considering a line $\langle\,a,x_{1}\,\rangle\subset\langle\,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$ and a hyperplane $\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$, we may take a point $\tilde{b}\in\langle\,a,x_{1}\,\rangle\cap\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$. Note that $\pi_{x_{1}}(\tilde{b})=\pi_{x_{1}}(a)=\alpha$. It gives the following locus of positive dimension,
$\left\{(\tilde{b},\tilde{x}_{2},\dots,\tilde{x}_{k})\mid(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}\right\}\subset\rho^{-1}(\alpha)\subset J^{\prime},$
which contradicts that $\rho$, given in (6), is generally finite. ∎
Lemma 14.
For a projective variety $Z\subset\langle\,Z\,\rangle={\mathbb{P}}^{\beta}$, suppose that $\sigma_{k-1}(Z)$ is a hypersurface in ${\mathbb{P}}^{\beta}$ and $\sigma_{k}(Z)={\mathbb{P}}^{\beta}$. Then, we have $q_{i}(p^{-1}(a))=Z$ for a general point $a\in\sigma_{k}(Z)$. In particular, there is an irreducible component $F$ of $p^{-1}(a)$ such that $q_{i}(F)=Z$.
Proof.
Since $p^{-1}(a)$ is invariant under permuting $x_{i}$-factors, we set $i=1$ for simplicity. Take a general $a\in\sigma_{k}(Z)={\mathbb{P}}^{\beta}$. Since $\operatorname{Vertex}(\sigma_{k-1}(Z))$ is a linear subvariety of $\sigma_{k-1}(Z)\subsetneq{\mathbb{P}}^{\beta}$ and $Z$ is non-degenerate in ${\mathbb{P}}^{\beta}$, $\operatorname{Vertex}(\sigma_{k-1}(Z))\cap Z\subsetneq Z$. This implies that for general $b\in Z$, $\dim\langle\,b,\sigma_{k-1}(Z)\,\rangle>\dim(\sigma_{k-1}(Z))$ so that $\langle\,b,\sigma_{k-1}(Z)\,\rangle={\mathbb{P}}^{\beta}$. Thus, any given general point $a\in{\mathbb{P}}^{\beta}$ sits on a line $\langle\,b,c\,\rangle$ for any general $b\in Z$ and for some general $c\in\sigma_{k-1}(Z)$ and $(a,b,\tilde{x_{2}},\dots,\tilde{x_{k}})\in p^{-1}(a)$, which means $q_{1}(p^{-1}(a))=Z$. ∎
Now let us focus on the case $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta=\binom{m+d}{m}-1}$, the image of the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$. Here we prove a very useful proposition, which has an independent interest itself. In Proposition 15, we consider the entry locus of a general point in $Z$ and estimate the dimension of the linear span of its image under ($d-1$)-uple Veronese embedding $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}=\binom{m+d-1}{m}-1}$.
Proposition 15.
Let $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta}$ with $d\geq 3$, $22\leq m\leq k-$ 2, and $\beta=\binom{m+d}{m}-1<km+k-1$. Assume $\dim(\sigma_{k-1}(Z))=(k-1)m+k-2<\beta$ (in other words, the ($k-1$)-th secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$). Let $J\subset{\mathbb{P}}^{\beta}\times(Z)^{k}$ be the Zariski closure of the incidence (5), let $(a,x_{1},\dots,x_{k})\in J$ be a general point, and let $F\subset J$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$. Then the following holds.
  1. (i)
    If $(k-1)m+k<\binom{m+d}{m}$, then we have
    $\dim\langle\,v_{d-1}(A)\,\rangle\geq k+(km+k-1)-\dim\sigma_{k}(Z)$
    for the preimage $A\subset{\mathbb{P}}^{m}$ of $q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\}\subset Z$ under $v_{d}:{\mathbb{P}}^{m}\simeq Z$ and for each $1\leq i\leq k$.
  2. (ii)
    If $(k-1)m+k=\binom{m+d}{m}$, then it holds $q_{i}(F)=Z$. In addition, if $(d,m)\neq(3,2)$, then $\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle\geq k+m$.
Remark 16.
  1. (a)
    Let $\mu=\frac{\binom{m+d}{m}}{m+1}$. Then inequalities $\beta<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ (i.e., $\binom{m+d}{m}<km+k$ and $(k-1)m+k-1<\binom{m+d}{m}$) are equivalent to $\mu<k<\mu+1$; it happens if and only if $\mu\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$.
  2. (b)
    In (ii) of Proposition 15, if $(d,m)=(3,2)$, then the condition $(k-1)m+k=\binom{m+d}{m}$ gives $k=4$. In this case, $q_{i}(F)=Z$ is still true (e.g., by Lemma 14), but $\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle=\dim\langle\,v_{2}({\mathbb{P}}^{2})\,\rangle=5$ is $k+m-1$, not greater or equal to $k+m$.
To prove Proposition 15, we settle two lemmas, Lemmas 17 and 19; the former one is technical and the latter geometric.
Lemma 17.
Let $d,m,k$ be integers such that $d\geq 3$, $22\leq m\leq k-$ 2, and $\binom{m+d}{m}-1<km+k-1$.
  1. (i)
    If $(k-1)m+k<\binom{m+d}{m}$, then $\binom{m+d-1}{m}-1-2m-k\geq 0$.
  2. (ii)
    If $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)\neq(3,2)$, then $\binom{m+d-1}{m}-1-m-k\geq 0$.
Proof.
(i) From Remark 18 (a) below, we may assume $m\geq 3$. Let $\delta:=\binom{m+d-1}{m}-(1+k+2m)$. Since $\binom{m+d-1}{m}=\frac{d}{m+d}\binom{m+d}{m}$, using $(k-1)m+k+1\leq\binom{m+d}{m}$, we have
$\delta=\binom{m+d-1}{m}-(1+k+2m)\geq\frac{1}{m+d}\Bigl{(}d((k-1)m+k+1)-(m+d)(1+k+2m)\Bigr{)},$
Setting $m=k-a$ with $a\geq 2$, we have
(7) $d((k-1)m+k+1)-(m+d)(1+k+2m)=km(d-1)-3dm-m-2m^{2}\\ =(m+a)m(d-1)-3dm-m-2m^{2}=m(dm+(a-3)d-3m-(a+1))\\ \geq m((d-3)(m-1)-6)$
Then, $\delta\geq 0$ holds in the following three cases: $d\geq 6$ and $m\geq 3$; $d=5$ and $m\geq 4$; or $d=4$ and $m\geq 7$. In addition, in Remark 18 (b) (see also Remark 16 (a)), we explicitly check $\delta\geq 0$ if $d=4$ and $m\leq 6$, and check that there is no $k$ in our range if $d=5$ and $m=3$.
On the other hands, when $d=3$, (7) implies
$d((k-1)m+k+1)-(m+d)(1+k+2m)=m((a-3)3-(a+1))=m(2a-10).$
Hence, $\delta\geq 0$ holds if $d=3$ and $a\geq 5$. For $d=3$ and $m=k-a$ with $a=2,3,4$, we have
$\delta=\frac{m^{2}-m-2k}{2}=\frac{m(m-3)-2a}{2}>0$
if $m\geq 5$. In addition, in Remark 18 (b), we explicitly check that there is no $k$ in our range if $d=3$ and $m\leq 4$.
(ii) Let $\delta:=\binom{m+d-1}{m}-(1+k+m)$. In the same way as (i), using $(k-1)m+k=\binom{m+d}{m}$, we have
$\delta=\binom{m+d-1}{m}-(1+k+m)=\frac{1}{m+d}\Bigl{(}d((k-1)m+k)-(m+d)(1+k+m)\Bigr{)},$
and have
$d((k-1)m+k)-(m+d)(1+k+m)=km(d-1)-2dm-m-m^{2}-d\\ \geq(m+2)m(d-1)-2dm-m-m^{2}-d=(d-2)(m^{2}-1)-2-3m.$
If $d\geq 3$ and $m\geq 4$, then since $(d-2)(m^{2}-1)-2-3m\geq m(m-3)-3\geq 1$, it holds $\delta\geq 0$.
If $d\geq 4$ and $m=3$, then since $(d-2)(m^{2}-1)-2-3m\geq 5$, it similarly holds $\delta\geq 0$.
On the other hands, if $d=3$ and $m=3$, then $(k-1)m+k=\binom{m+d}{m}$ implies $204k-3=\frac{(d+3)(d+2)(d+1)}{3!}=$ 20; this case does not occur since $k$ cannot be an integer.
If $d\geq 4$ and $m=2$, then $3k-2=\frac{(d+2)(d+1)}{2}$. In this case, $\delta=\frac{(d+1)d}{2}-k-3\geq 0$, because of
$3\left(\frac{(d+1)d}{2}-k-3\right)\geq\frac{3(d+1)d}{2}-\frac{(d+2)(d+1)}{2}-11=d^{2}-12\geq 4$
Remark 18.
In the proof of Lemma 17, we refer to two calculations as follows.
  1. (a)
    Let $m=2$ and $km+k-1>\binom{m+d}{m}-1$. Then $(k-1)m+k<\binom{m+d}{m}$ does not occur. Otherwise, we have $3k-2<\binom{m+d}{m}=\frac{(d+2)(d+1)}{2}<3k$, and then $\frac{(d+2)(d+1)}{2}=3k-1$. Considering the congruence modulo $3$, we have $(d+2)(d+1)\equiv 6k-2\equiv 1\pmod{3}$. Then the possible values of $(d+2)(d+1)$ are
    $(d+2)(d+1)\equiv\begin{cases}2\cdot 1\equiv 2&(d=0)\\ 0\cdot 2\equiv 0&(d=1)\\ 1\cdot 0\equiv 0&(d=2),\end{cases}$
    modulo $3$, which is absurd.
  2. (b)
    For $d=3,4,5$, we calculate numbers $m$ such that $\mu:=\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$ satisfies $(k-1)m+k<\beta$. For $d=3$, the smallest $m=5$. For $d=4$, the smallest $m=3$ and the next smallest $m=7$. For $d=5$, the smallest $m=9$. The explicit values of $\delta=\binom{m+d-1}{m}-(1+k+2m)$ for them are obtained as follows:
    $(d,m,k,\delta)=(3,5,10,0),(4,3,9,4),(4,7,42,63),(5,9,201,495).$
The next lemma concerns a general fact on linear sections of Veronese varieties, which has an independent interest itself.
Lemma 19.
Let $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ be general $k$ points, let $v_{e}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{e}=\binom{m+e}{m}-1}$ be the $e$-uple Veronese embedding of ${\mathbb{P}}^{m}$, and let $R\subset{\mathbb{P}}^{\beta_{e}}$ be any $k$-plane containing the ($k-1$)-plane $M:=\langle\,v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle$. Then the following holds.
  1. (i)
    Assume $k\leq\beta_{e}-m$, and assume that there is a curve $C\subset{R\cap v_{e}({\mathbb{P}}^{m})}$ passing through $v_{e}(x_{1}^{\prime})$. Then it holds that
    (8) $R\subset\langle\,{v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})},{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}\,\rangle.$
  2. (ii)
    Assume $e\geq 3$ and $k\leq\beta_{e}-2m$. Then we have
    $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0~{},$
    where the left-hand side means dimension of component(s) passing through $v_{e}(x_{1}^{\prime})$. In particular, the set of a point $\left\{v_{e}(x_{1}^{\prime})\right\}$ is an irreducible component of ${R\cap v_{e}({\mathbb{P}}^{m})}$.
  3. (iii)
    Assume $e=2$ and $k-1\leq\beta_{2}-2m$. Assume that there is a curve $C\subset R\cap v_{2}({\mathbb{P}}^{m})$ such that $v_{2}(x_{1}^{\prime})\in C$. Then, for any irreducible subset $D\subset R\cap v_{2}({\mathbb{P}}^{m})$, it holds $D\subset v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$ for some $l=2,\dots,k$, where $v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$ is a conic curve in ${\mathbb{P}}^{\beta_{2}}$ given as the image of the line $\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle\subset{\mathbb{P}}^{m}$,
Proof.
Let $C\subset{R\cap v_{e}({\mathbb{P}}^{m})}$ be a curve passing through $v_{e}(x_{1}^{\prime})$. Let
$\pi=\pi_{\langle\,v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle}:{\mathbb{P}}^{\beta_{e}}\dashrightarrow{\mathbb{P}}^{\beta_{e}-k+1}$
be the linear projection from the ($k-2$)-plane $\langle\,v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle$. If $k\leq\beta_{e}-m$, then the generalized trisecant lemma ([28, proposition 1.4.3]) implies $M\cap v_{e}({\mathbb{P}}^{m})=\left\{v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\right\}$; in particular, $C\not\subset M$. Hence $C$ is not contracted to a point under $\pi$. Then $\overline{\pi(C)}\subset{\mathbb{P}}^{\beta_{e}-k+1}$ is a curve passing through $\pi(v_{e}(x_{1}^{\prime}))$, and then $\overline{\pi(C)}=\overline{\pi(R)}$, which is indeed a line. It follows
$\overline{\pi(C)}=\overline{\pi(R)}\subset\mathbb{T}_{\pi(v_{e}(x_{1}^{\prime}))}\overline{\pi(v_{e}({\mathbb{P}}^{m}))}$
where, by generic smoothness, the right hand side is equal to $\pi({\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})})$ since $x_{1}^{\prime}\in{\mathbb{P}}^{m}$ is general. It follows that $R$ is contained in the preimage of $\pi({\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})})$, which implies the inclusion (8) of (i).
The condition $k\leq\beta_{e}-m$ holds if $k$ or $k-1$ is $\leq\beta_{e}-2m$. Next, we consider a tangential projection
$\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}:{\mathbb{P}}^{\beta_{e}}\dashrightarrow{\mathbb{P}}^{\beta_{e}-m-1}$
from the $m$-plane ${\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}\subset{\mathbb{P}}^{\beta_{e}}$, and its restriction $\tilde{\pi}=\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}|_{v_{e}({\mathbb{P}}^{m})}$ on the Veronese variety. Note that $v_{e}({\mathbb{P}}^{m})\cap{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}=\left\{v_{e}(x_{1}^{\prime})\right\}$. If $R\subset{\mathbb{P}}^{\beta}$ satisfies (8), we have
(9) $\tilde{\pi}(R\cap v_{e}({\mathbb{P}}^{m}))\subset\tilde{\pi}(v_{e}({\mathbb{P}}^{m}))\cap\langle\,\tilde{\pi}(v_{e}(x_{2}^{\prime})),\dots,\tilde{\pi}(v_{e}(x_{k}^{\prime}))\,\rangle.$
Let $e\geq 3$ and $k\leq\beta_{e}-2m$. Suppose that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})>0$, which means the existence of a curve $C$ satisfying the condition of (i). Since $\operatorname{codim}(\tilde{\pi}(v_{e}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{e}-m-1})-(k-1)\geq(\beta_{e}-2m-1)-(k-1)\geq 0$, again the trisecant lemma implies that the right hand side in (9) is only the set of $k-1$ points ${\tilde{\pi}(v_{e}(x_{2}^{\prime})),\dots,\tilde{\pi}(v_{e}(x_{k}^{\prime}))}$. Thus, each irreducible subset $D\subset R\cap v_{e}({\mathbb{P}}^{m})$ satisfies
(10) $\tilde{\pi}(D)=\tilde{\pi}(v_{e}(x_{l}^{\prime}))$
for some $l=2,\dots,k$. (At least, taking $D=C$, we exactly have (10).) From $e\geq 3$, $\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))=2m+1$ (i.e., non-defective). Then $\tilde{\pi}$ must be a generically finite map since general two tangent spaces of $\sigma_{2}(v_{e}({\mathbb{P}}^{m}))$ do not meet each other, and we reach a contradiction since $v_{e}(x_{l}^{\prime})$ is a general point. It implies that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0$.
Finally, we consider the case of $e=2$. Then it is well known that the tangential projection $\tilde{\pi}:v_{2}({\mathbb{P}}^{m})\dashrightarrow\tilde{\pi}(v_{2}({\mathbb{P}}^{m}))$ has fibers of dimension $1$. Moreover, as in Remark 8 (a) we know that $\pi_{\mathbb{T}_{v_{2}(x_{1}^{\prime})}v_{2}({\mathbb{P}}^{m})}:{\mathbb{P}}^{\beta_{2}=\frac{(m+2)(m+1)}{2}-1}\dashrightarrow{\mathbb{P}}^{\beta_{2}-m-1=\frac{(m+1)m}{2}-1}$ satisfies the commutative diagram:
${{\mathbb{P}}^{m}}$${{\mathbb{P}}^{\frac{(m+2)(m+1)}{2}-1}}$${{\mathbb{P}}^{m-1}}$${{\mathbb{P}}^{\frac{(m+1)m}{2}-1}\quad,}$$\scriptstyle{v_{2}}$$\scriptstyle{\pi_{x_{1}^{\prime}}}$$\scriptstyle{\pi_{\mathbb{T}_{v_{2}(x_{1}^{\prime})}v_{2}({\mathbb{P}}^{m})}}$$\scriptstyle{v_{2}}$
where $\pi_{x_{1}^{\prime}}:{\mathbb{P}}^{m}\dashrightarrow{\mathbb{P}}^{m-1}$ is the linear projection from $x_{1}^{\prime}$, and $v_{2}:{\mathbb{P}}^{m-1}\hookrightarrow{\mathbb{P}}^{\frac{(m+1)m}{2}-1}$ is the Veronese embedding of ${\mathbb{P}}^{m-1}$. Then $\dim(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})))=\dim(v_{2}({\mathbb{P}}^{m-1}))=m-1$. If $k-1\leq\beta_{2}-2m$, then $\operatorname{codim}(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{2}-m-1})-(k-1)\geq(\beta_{2}-2m)-(k-1)\geq 0$; hence the trisecant lemma implies $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$ for some $l=2,\dots,k$ as we discussed for (10).
In the above diagram, for any $y^{\prime}\in\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle\subset{\mathbb{P}}^{m}$ with $y^{\prime}\neq x_{1}^{\prime}$, we have $\tilde{\pi}(v_{2}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(x_{l}^{\prime}))=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$; indeed, $\tilde{\pi}^{-1}(\tilde{\pi}(v_{2}(x_{l}^{\prime})))=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$. Since $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$, we have $D\subset v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$. ∎
Now we give a proof for Proposition 15. For simplicity, we set $i=1$. Then $x_{1}\in q_{1}(F)\subset Z=v_{d}({\mathbb{P}}^{m})$.
Proof of Proposition 15.
(i) Set $s=\dim\sigma_{k}(Z)$. Then an irreducible component $F$ of $p^{-1}(a)$ is of dimension $\dim J-s=(km+k-1)-s$. From Lemma 13, we have $\dim q_{1}(F)=(km+k-1)-s$.
Let $q_{1}(F)^{\prime}\subset{\mathbb{P}}^{m}$ be the preimage of $q_{1}(F)\subset Z$ under $v_{d}:{\mathbb{P}}^{m}\simeq Z$, and let $A=q_{1}(F)^{\prime}\cup\left\{x_{1}^{\prime},\dots,x_{k}^{\prime}\right\}\subset{\mathbb{P}}^{m}$. Let $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}}$ be the ($d-1$)-uple Veronese embedding. Then the ($k-1$)-plane $\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset{\mathbb{P}}^{\beta_{d-1}}$ is contained in the linear variety
$\langle\,v_{d-1}(A)\,\rangle=\langle\,v_{d-1}(q_{1}(F)^{\prime})\cup\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}\,\rangle,$
and is of codimension $c=\dim\langle\,v_{d-1}(A)\,\rangle-(k-1)$. By Lemma 17 (i), $\beta_{d-1}-2m-k\geq 0$. So, by the generalized trisecant lemma, $v_{d-1}({\mathbb{P}}^{m})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle=\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$. In particular, $v_{d-1}(q_{1}(F)^{\prime})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$.
Since $\dim q_{1}(F)^{\prime}\geq 1$, we may take a point $y^{\prime}\in q_{1}(F)^{\prime}$ such that
$v_{d-1}(y^{\prime})\notin\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle.$
Assume $d\geq 4$. Applying Lemma 19 (ii) to $R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle\subset\langle\,v_{d-1}(A)\,\rangle$, we have
$\dim_{v_{d-1}(x_{1}^{\prime})}({R\cap v_{d-1}({\mathbb{P}}^{m})})=0.$
In particular, $\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}(q_{1}(F)^{\prime}))=0$. On the other hand, regarding it as an intersection of two irreducible subvarieties in $\langle\,v_{d-1}(A)\,\rangle$, we deduce that every irreducible component of $R\cap v_{d-1}(q_{1}(F)^{\prime})$ is of dimension $\geq\dim(v_{d-1}(q_{1}(F)^{\prime}))-(c-1)$. Hence $\dim(\langle\,v_{d-1}(A)\,\rangle)\geq k+\dim(v_{d-1}(q_{1}(F)^{\prime}))=k+(km+k-1)-s$.
Next, let us consider the case of $d=3$. For $l=2,\dots,k$, since ${v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)}\subset{\mathbb{P}}^{\beta_{2}}$ is a conic, it follows that $\langle\,v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$ is a $2$-plane, which is equal to $\langle\,v_{2}(x_{1}^{\prime}),v_{2}(x_{l}^{\prime}),z\,\rangle$ for some $z\in{\mathbb{P}}^{\beta_{2}}$. Then
$\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),z\,\rangle$
is a linear subvariety of dimension $\leq k$. Since $v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$ is empty or is a set of points, the intersection $v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$ is of dimension $\leq 1$. On the other hand, since $m\neq 2$ by Remark 18 (a), we have $\dim q_{1}(F)^{\prime}=(km+k-1)-\beta\geq 2$ as in Remark 20 below. For the union
$W=\bigcup_{l=2,\dots,k}v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle\subset{\mathbb{P}}^{\beta_{2}},$
we see that $q_{1}(F)^{\prime}\setminus v_{2}^{-1}(W)\neq\emptyset$ and may take $y^{\prime}\in q_{1}(F)^{\prime}\setminus v_{2}^{-1}(W)$.
Let $R=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(y^{\prime})\,\rangle\subset\langle\,v_{2}(A)\,\rangle$ and suppose that $\dim_{v_{2}(x_{1}^{\prime})}(R\cap v_{2}(q_{1}(F)^{\prime}))>0$, that is to say, there is a curve $C\subset R\cap v_{2}(q_{1}(F)^{\prime})$ containing ${v_{2}(x_{1}^{\prime})}$. Taking $D=C$ and applying Lemma 19 (iii), we have $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$ for some $l>1$. If $\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k$, then $R=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$, contradicting to the definition of $W$ and our choice of $y^{\prime}$. Else if, $\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k-1$, then $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\subset v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$, also contradicting that the intersection is of dimension at most 00.
Hence $\dim_{v_{2}(x_{1}^{\prime})}(R\cap v_{2}(q_{1}(F)^{\prime}))=0$. Then, in the same way as above, we have $\dim(\langle\,v_{2}(A)\,\rangle)\geq k+(km+k-1)-s$.
(ii) In the case when $(k-1)m+k=\binom{m+d}{m}$, we have $km+k-1-s\geq m$. It follows from Lemma 13 and ${\mathbb{P}}^{m}\simeq Z$ that $q_{i}(F)=Z$. From Lemma 17 (ii), if $(d,m)\neq(3,2)$, then ${\mathbb{P}}^{\beta_{d-1}}=\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle$ is of dimension $\geq k+m$. ∎
Remark 20.
For $d=3$, $m\neq 2$, if $\mu:=\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$ (as under the conditions of Proposition 15 and Remark 16 (a)), then we have $(km+k-1)-\beta_{d}\neq 1$. The reason is as follows. First, we may write $\mu={(m+3)(m+2)}/6=M/3$ for some $M\in{\mathbb{N}}$ since $(m+3)(m+2)$ is divided by $2$. In addition, dividing $M$ by $3$ with remainder, we have $M=3Q+R$ for a quotient $Q$ and a remainder $R$. Since $\mu=M/3\notin{\mathbb{N}}$, $R$ must be $1$ or $2$. In this setting, $k=\lceil M/3\rceil=Q+1$. It follows that $(km+k-1)-\beta_{d}$ is equal to
$(km+k)-\tbinom{m+3}{3}=(m+1)\left({k-\tfrac{(m+3)(m+2)}{6}}\right)=(m+1)\left((Q+1)-(Q+\tfrac{R}{3})\right)=(m+1)\cdot\tfrac{3-R}{3}~{}.$
If $(km+k-1)-\beta_{d}=1$, then $3=(m+1)(3-R)$. Since $3-R$ is $1$ or $2$ and $m\in{\mathbb{N}}$, we get $m=2$.
We end this subsection by making one more important remark on the case of secant defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s, which do not satisfy the condition $\beta<km+k-1$ in Proposition 15.
Remark 21 (Estimate in defective cases).
For four defective cases $(k,d,m)=(5,4,2)$,$(9,4,3)$,$(14,4,4)$ and $(7,3,4)$, similarly as in Proposition 15, we can have an estimation
$\dim\langle\,v_{d-1}(A)\,\rangle\geq k+\delta~{},$
where $A=v_{d}^{-1}(q_{1}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$, the preimage of the entry locus of $a$, and $\delta$ is the secant defect of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$; here $A$ is $\delta$-equidimensional, the $k$ general points $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ are contained in $A$, and it’s well known that $\delta=2$ when $(k,d,m)=(9,4,3)$ and $\delta=1$ in all the other defective cases.
For 3 cases $(k,d,m)=(5,4,2),(9,4,3)$, and $(14,4,4)$, we see that $\beta_{d-1}-2m\geq k$ and $v_{d-1}(A)\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$ by the trisecant lemma so that we may take a point $y^{\prime}\in A$ such that $v_{d-1}(y^{\prime})\notin\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle$. By Lemma 19 (ii), we get $\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}({\mathbb{P}}^{m}))=0$ where $R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle$. Thus, by the same intersection argument in $\langle\,v_{d-1}(A)\,\rangle$ (as Proposition 15 (i)), we derive the estimation
$\dim\langle\,v_{d-1}(A)\,\rangle\geq\dim R+\dim v_{d-1}(A)=k+\delta~{}.$
For the remaining case $(k,d,m)=(7,3,4)$, it holds $\beta_{d-1}-2m=k-1$, and we still can claim that $\dim(\langle\,v_{2}(A)\,\rangle)\geq k+\delta=7+1=8$ as follows. For the $6$-dimensional subspace $M:=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\,\rangle\subset\langle\,v_{2}(A)\,\rangle$, the trisecant lemma implies $M\cap v_{2}(A)\subset M\cap v_{2}({\mathbb{P}}^{4})=\left\{v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\right\}$, 00-dimensional intersection. Then $\dim(\langle\,v_{2}(A)\,\rangle)\geq 7$ (otherwise, we get $M=\langle\,v_{2}(A)\,\rangle$ so that $M\cap v_{2}(A)=v_{2}(A)$, a contradiction). Suppose that $\dim(\langle\,v_{2}(A)\,\rangle)=7$, and set $R=\langle\,v_{2}(A)\,\rangle$. We take the irreducible components of $1$-equidimensional closed set $A$ as $A=\bigcup_{j=1}^{s}A_{j}$. Note that $v_{2}(A_{j})\subset R\cap v_{2}({\mathbb{P}}^{m})$. Since $x_{1}^{\prime}\in A$, there is a curve $A_{j_{0}}$ containing $x_{1}^{\prime}$. Taking $C=v_{2}(A_{j_{0}})$ and applying Lemma 19 (iii), for any $j$ with $1\leq j\leq s$, we have $v_{2}(A_{j})=v_{2}(\langle x_{1}^{\prime},x_{l_{j}}^{\prime}\rangle)$ for some $l_{j}=2,\dots,k$. It is equivalent to $A_{j}=\langle x_{1}^{\prime},x_{l_{j}}^{\prime}\rangle$, a line in ${\mathbb{P}}^{4}$; in particular, $x_{1}^{\prime}\in A_{j}$. In the same way, $A_{j}$ must contain $x_{1}^{\prime},\dots,x_{7}^{\prime}$. But this is a contradiction, because these points are chosen as general $7$ points in ${\mathbb{P}}^{4}$. Thus, $\dim(\langle\,v_{2}(A)\,\rangle)\geq 8$.
2.3. Estimate for the linear span of tangents moving along a subsecant variety
First, we give an explicit description of the embedded tangent space $\mathbb{T}_{x}v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$ at a point $x$ in $v_{d}({\mathbb{P}}^{n})$ or $v_{d}({\mathbb{P}}^{m})$ as follows (see also [14]).
Recall that $\mathsf{mono}[t]_{\leq e}$ denotes the set of monomials $f\in{\mathbb{C}}[t_{1},\dots,t_{m}]$ with $\deg f\leq e$. Note that $1\in\mathsf{mono}[t]_{\leq e}$ as the monomial of degree 00. As mentioned in Remark 8, as changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}:=n-m$, we may assume that ${\mathbb{P}}^{m}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$. On the affine open subset $\left\{t_{0}\neq 0\right\}$, the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized by monomials of ${\mathbb{C}}[t_{1},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}]$ of degree $\leq d$. So it is expressed as
(11) $[\,\mathsf{mono}[t]_{\leq d}:u_{1}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad\cdots\quad:u_{m^{\prime}}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad*\quad\,]~{},$
where $u_{i}\cdot\mathsf{mono}[t]_{\leq d-1}$ means $\left\{u_{i}f\mid f\in\mathsf{mono}[t]_{\leq d-1}\right\}=(u_{i}:u_{i}t_{1}:u_{i}t_{2}:\dots:u_{i}t_{m}^{d-1})$, and “*” means the remaining monomials.
Let $x=v_{d}(x^{\prime})$ with $x^{\prime}\in\left\{t_{0}\neq 0\right\}\subset{\mathbb{P}}^{n}$. Then $\mathbb{T}_{x}v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$ coincides with
(12) $n$-plane spanned by the $(n+1)$ points corresponding to the row vectors of the matrix
$\begin{bmatrix}v_{d}\\ \partial v_{d}/\partial t_{1}\\ \vdots\\ \partial v_{d}/\partial t_{m}\\ \partial v_{d}/\partial u_{1}\\ \vdots\\ \partial v_{d}/\partial u_{m^{\prime}}\end{bmatrix}(x^{\prime})$ =[$\mathsf{mono}[t]_{\leq d}&u_{1}\cdot\mathsf{mono}[t]_{\leq d-1}&\dots&u_{m^{\prime}}\cdot\mathsf{mono}[t]_{\leq d-1}&&*&\\ (\mathsf{mono}[t]_{\leq d})_{t_{1}}&u_{1}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{1}}&\dots&u_{m^{\prime}}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{1}}&&*&\\ \vdots&\vdots&&\vdots&&\vdots&\\ (\mathsf{mono}[t]_{\leq d})_{t_{m}}&u_{1}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{m}}&\dots&u_{m^{\prime}}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{m}}&&*&\\ \mathsf{O}&\mathsf{mono}[t]_{\leq d-1}&\dots&\mathsf{O}&&*&\\ \vdots&&\ddots&&&\vdots&\\ \mathsf{O}&\mathsf{O}&\dots&\mathsf{mono}[t]_{\leq d-1}&&*&\end{array}\right](x^{\prime})~{}$
using (11), where $(\mathsf{mono}[t]_{\leq e})_{t_{i}}$ means $\left\{\partial f/\partial t_{i}\mid f\in\mathsf{mono}[t]_{\leq e}\right\}$ and $\mathsf{O}$ is a zero matrix with suitable size.
In particular, in case of $x^{\prime}\in{\mathbb{P}}^{m}=\left\{u_{1}=\dots=u_{m^{\prime}}=0\right\}$, we see that the matrix is of the form
(13) [$\mathsf{mono}[t]_{\leq d}&\mathsf{O}&\dots&\mathsf{O}&&&&&\mathsf{O}&&&\\ (\mathsf{mono}[t]_{\leq d})_{t_{1}}&\mathsf{O}&\dots&\mathsf{O}&&&&&\mathsf{O}&&&\\ \vdots&\vdots&&\vdots&&&&&\vdots&&&\\ (\mathsf{mono}[t]_{\leq d})_{t_{m}}&\mathsf{O}&\dots&\mathsf{O}&&&&&\mathsf{O}&&&\\ \mathsf{O}&\mathsf{mono}[t]_{\leq d-1}&\dots&\mathsf{O}&&&&&\mathsf{O}&&&\\ \vdots&&\ddots&&&&&&\vdots&&&\\ \mathsf{O}&\mathsf{O}&\dots&\mathsf{mono}[t]_{\leq d-1}&&&&&\mathsf{O}&&&\end{array}\right](x^{\prime})~{}.$
As a consequence, we settle a key proposition which estimates a lower bound of dimension of linear span of moving embedded tangent spaces along a subset of a given ${\mathbb{P}}^{m}$.
Proposition 22.
Let $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ be the $d$-uple Veronese embedding. For an $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$, for a (possibly reducible) subset $A\subset{\mathbb{P}}^{m}$, and for a linear variety $\Lambda\subset\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle$, the dimension of the linear variety
$\langle\,\Lambda\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle\subset{\mathbb{P}}^{N}$
is greater than or equal to
(14) $\dim\langle\,\Lambda\cup v_{d}(A)\,\rangle+(n-m)\big{\{}1+\dim\langle\,v_{d-1,m}(A)\,\rangle\big{\}},$
where $v_{e,m}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\binom{m+e}{m}-1}$ is the $e$-uple Veronese embedding of ${\mathbb{P}}^{m}$.
Proof of Proposition 22.
For a given $A\subset{\mathbb{P}}^{m}$, we consider
$B_{0}=v_{d}(A),B_{1}=(\partial v_{d}/\partial u_{1})(A),~{}\dots~{},B_{m^{\prime}}=(\partial v_{d}/\partial u_{m^{\prime}})(A)$
as subsets in ${\mathbb{P}}^{N}$, where $B_{i}$ is embedded by the parameterization of $(m+1+i)$-th row of the matrix of (12) for $1\leq i\leq m^{\prime}$. Note that for the homogeneous coordinates $[w_{0}:\dots:w_{N}]$ on ${\mathbb{P}}^{N}$ corresponding to (11), $\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle={\mathbb{P}}^{\beta=\binom{m+d}{m}-1}$ is the zero set of $w_{\beta+1},\dots,w_{N}$, and $\Lambda\cup B_{0}$ is contained in the set.
Since $A\subset\left\{u_{1}=\dots=u_{m^{\prime}}=0\right\}$, it follows from (13) that the linear variety
(15) $\langle\,\Lambda\cup B_{0},B_{1},~{}\dots~{},B_{m^{\prime}}\,\rangle\subset{\mathbb{P}}^{N}$
is of dimension $\dim(\langle\,\Lambda\cup B_{0}\,\rangle)+\dim(\langle\,B_{1}\,\rangle)+\dots\dim(\langle\,B_{m^{\prime}}\,\rangle)+m^{\prime}$.
By (13) again, we see that $B_{0}\simeq v_{d}(A)$ and $B_{i}\simeq v_{d-1,m}(A)$ for $1\leq i\leq m^{\prime}$. Since the linear variety (15) is contained in $\langle\,\Lambda\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle$, we have the assertion. ∎
3. Case of $m=1$
3.1. Symmetric flattening and conormal space computation
For the proof of Theorem 1, we begin with some preliminaries on equations for secant varieties and conormal space computation via known sets of equations, whereas we mainly adopt geometric viewpoint and techniques for $m\geq 2$ case in §4.
Let $V$ be a $(n+1)$-dimensional ${\mathbb{C}}$-vector space ${\mathbb{C}}\langle x_{0},x_{1},\dots,x_{n}\rangle$. For the Veronese variety $v_{d}({\mathbb{P}}V)$, we have a natural 1-1 correspondence between a point of the ambient space $\langle\,v_{d}({\mathbb{P}}V)\,\rangle$ and a homogeneous polynomial $f$ of degree $d$ (or a form) in $S:={\mathbb{C}}[x_{0},x_{1},\dots,x_{n}]$. First of all, let us recall some related notions to this correspondence.
Given a form $f$ of degree $d$, the minimum number of linear forms $l_{i}$ needed to write $f$ as a sum of $d$-th powers is the so-called (Waring) rank of $f$ and let us denote it by $\operatorname{rank}(f)$. The (Waring) border rank is given by the same notion in the limiting sense. In other words, if there is a family $\{f_{\epsilon}\mid\epsilon>0\}$ of polynomials with constant rank $r$ and $\lim_{\epsilon\to 0}f_{\epsilon}=f$, then we say that $f$ has border rank at most $r$. The minimum such $r$ is called the border rank of $f$ and let’s denote it again by $\underline{\mathrm{rank}}(f)$. Note that by definition $\sigma_{k}(v_{d}({\mathbb{P}}V))$ is the variety of homogeneous polynomials $f$ of degree $d$ with border rank $\underline{\mathrm{rank}}(f)\leq k$.
Now, we recall that some part of defining equations for $\sigma_{k}(v_{d}({\mathbb{P}}V))$ comes from so-called symmetric flattenings. Consider the polynomial ring $S=S^{\bullet}V={\mathbb{C}}[x_{0},\ldots,x_{n}]$ and consider another polynomial ring $T=S^{\bullet}V^{\ast}={\mathbb{C}}[y_{0},\ldots,y_{n}]$, where $V^{\ast}$ is the dual ${\mathbb{C}}$-vector space of $V$. Define the differential action of $T$ on $S$ as follows: for any $g\in T_{d-a},f\in S_{d}$, we set
$g\cdot f=g(\partial_{0},\partial_{1},\ldots,\partial_{n})f\in S_{a}~{},$
where $\partial_{i}=\partial/\partial{x_{i}}$ Let us take bases for $S_{a}$ and $T_{d-a}$ as
$\mathbf{X}^{I}=\frac{1}{i_{0}!\cdots i_{n}!}x_{0}^{i_{0}}\cdots x_{n}^{i_{n}}$ $\quad\textrm{and}\quad\mathbf{Y}^{J}=y_{0}^{j_{0}}\cdots y_{n}^{j_{n}}~{},$
with $|I|=i_{0}+\cdots+i_{n}=a$ and $|J|=j_{0}+\cdots+j_{n}=d-a$. For a given $f=\sum_{|I|=d}b_{I}\cdot\mathbf{X}^{I}$ in $S_{d}$, we have a linear map
$\phi_{d-a,a}(f):T_{d-a}\to S_{a},\quad g\mapsto g\cdot f$
for any $a$ with $11\leq a\leq d-$ 1, which can be represented by the following (${a+n\choose n}\times{d-a+n\choose n}$-matrix:
($\ &b_{I,J}&\\ &&\end{array}\right)\quad\textrm{with $b_{I,J}=b_{I+J}$}~{},$
in the bases defined above. We call this ‘$(d-a,a)$-symmetric flattening (or catalecticant) matrix’ of $f$. It is easy to see that the transpose $\phi_{d-a,a}(f)^{T}$ is equal to $\phi_{a,d-a}(f)$.
It is obvious that if $f$ has (border) rank 1, then any symmetric flattening $\phi_{d-a,a}(f)$ has rank 1. By subadditivity of matrix rank, we also know that $\operatorname{rank}~{}\phi_{d-a,a}(f)\leq r$ if $\underline{\mathrm{rank}}(f)\leq r$. So, we obtain a set of defining equations coming from $(k+1)$-minors of the matrix $\phi_{d-a,a}(f)$ for $\sigma_{k}(v_{d}({\mathbb{P}}V))$. For some small values of $k$, it is known that these minors are sufficient to cut the variety $\sigma_{k}(v_{d}({\mathbb{P}}V))$ scheme-theoretically (see [23, theorem 3.2.1 (1)]).
Let us recall another basic terms and facts. Let $Z\subset{\mathbb{P}}W$ be a (reduced and irreducible) variety and $\hat{Z}$ be its affine cone in $W$. Consider a (closed) point $p\in\hat{Z}$ and say $[p]$ the corresponding point in ${\mathbb{P}}W$. We denote the affine tangent space to $Z$ at $[p]$ in $W$ by $\hat{T}_{[p]}Z$ and we define the (affine) conormal space to $Z$ at $[p]$, $\hat{N}^{\ast}_{[p]}Z$ as the annihilator $(\hat{T}_{[p]}Z)^{\perp}\subset W^{\ast}$. Since $\dim\hat{N}^{\ast}_{[p]}Z+\dim\hat{T}_{[p]}Z=\dim W$ and $\dim Z\leq\dim\hat{T}_{[p]}Z-1$, we get that $\dim\hat{N}^{\ast}_{[p]}Z\leq\operatorname{codim}(Z,{\mathbb{P}}W)$ and the equality holds if and only if $Z$ is smooth at $[p]$. This conormal space is quite useful to study the (embedded) tangent space $\mathbb{T}_{[p]}Z$.
For any given form $f\in S^{d}V$, we call $\partial\in T_{t}$ apolar to $f$ if the differentiation $\partial\cdot f$ gives zero (i.e. $\partial\in\ker\phi_{t,d-t}(f)$). And we define the apolar ideal $f^{\perp}\subset T$ as
$f^{\perp}:=\{\partial\in T~{}|~{}\partial\cdot f=0\}~{}.$
It is straightforward to see that $f^{\perp}$ is indeed an ideal of $T$. Moreover, it is well-known that the quotient ring $T_{f}:=T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ (see e.g. [19, chapter 1]). In terms of this apolar ideal, we have a useful description of (a part of) conormal space as follows:
Proposition 23.
Suppose that $f\in S^{d}V$ corresponds to a (closed) point of $\sigma_{k}(v_{d}({\mathbb{P}}V))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}V))$. Then, for any $a$ with $1\leq a\leq{\lfloor\frac{d+1}{2}\rfloor}$ with $\operatorname{rank}\phi_{d-a,a}(f)=k$ we have
$\hat{N}^{\ast}_{[f]}\sigma_{k}(v_{d}({\mathbb{P}}V))\supseteq(f^{\perp})_{a}\cdot(f^{\perp})_{d-a}$
as a subspace of $T_{d}=S^{d}V^{\ast}$.
Proof.
Let $X\subset{\mathbb{P}}W$ be any variety. For any linear embedding $W\hookrightarrow A\otimes B$ and the induced embedding $X\subset{\mathbb{P}}W\hookrightarrow{\mathbb{P}}(A\otimes B)$, it is well-known that for any $f\in\sigma_{k}(X)\subset{\mathbb{P}}(A\otimes B)$, considering $\sigma_{k}(X)$ as a subvariety of ${\mathbb{P}}(A\otimes B)$, we have
$\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\supseteq\ker(f)\otimes\operatorname{im}(f)^{\perp}=\hat{N}^{\ast}_{[f]}\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))\quad\textrm{in $A^{\ast}\otimes B^{\ast}$}$
provided that $X\subseteq\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$, $X\nsubseteq\sigma_{p-1}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$ and $f$ has rank $k\cdot p$ as a linear map in $\operatorname{Hom}(A^{\ast},B)$ (see e.g. [23, section 2.5]). Here, $Seg({\mathbb{P}}A\times{\mathbb{P}}B)$ means the Segre variety in ${\mathbb{P}}(A\otimes B)$.
Further, since $X\subset{\mathbb{P}}W\subset{\mathbb{P}}(A\otimes B)$, then as a subvariety of ${\mathbb{P}}W$ it holds that
$\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\supseteq\pi(\ker(f)\otimes\operatorname{im}(f)^{\perp})=\hat{N}^{\ast}_{[f]}(\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))\cap{\mathbb{P}}W)~{},$
where $\pi:A^{\ast}\otimes B^{\ast}\to W^{\ast}$ is the dual map of the given inclusion $W\hookrightarrow A\otimes B$.
The assertion is immediate when we apply this fact to a partial polarization $S^{d}V\hookrightarrow S^{a}V\otimes S^{d-a}V$, because $X=v_{d}({\mathbb{P}}V)$ is contained in $Seg({\mathbb{P}}S^{a}V\times{\mathbb{P}}S^{d-a}V)$ in ${\mathbb{P}}(S^{a}V\otimes S^{d-a}V)$ (i.e. $p=1$ case) and $\operatorname{rank}\phi_{d-a,a}(f)=k$, $\ker\phi_{d-a,a}(f)=(f^{\perp})_{d-a}$, $\operatorname{im}(\phi_{d-a,a}(f))^{\perp}=(f^{\perp})_{a}$. ∎
3.2. Proof of Theorem 1
Now we study singularity and non-singularity of the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ in each range of $k,d$ as Theorem 1.
Proof of Theorem 1.
(i) Let $f$ be any form belonging to $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$. Set $X=v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$, the Veronese variety. Consider $f$ as a polynomial in ${\mathbb{C}}[x_{0},x_{1}]$ as in §3.1. Then, by [19, theorem 1.44] we know that $T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ and that $f^{\perp}$ is a complete intersection of two homogeneous polynomials $F,G$ of each degree $a$ and $b$ $(a\leq b)$ with $a+b=d+2$ as an ideal of ${\mathbb{C}}[y_{0},y_{1}]$, where the Hilbert function of $T/f^{\perp}$ is $(1,2,\cdots,a-1,a,\cdots,a,a-1,\cdots,2,1)$. We claim that $\operatorname{rank}~{}\phi_{k,d-k}(f)=k$ (i.e. $a=k$). If $a<k$, then $\operatorname{rank}~{}\phi_{t,d-t}(f)<k$ for all $t$. Since it is well-known that the $k$-minors of catalecticant $\phi_{t,d-t}$ for any $k-1\leq t\leq d-(k-1)$ gives the ideal of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))$, we get $f\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ which is contradiction.
Hence, we have that $f^{\perp}=\big{(}F,G,y_{2},\ldots,y_{n}\big{)}$ as an ideal in $T={\mathbb{C}}[y_{0},y_{1},\ldots,y_{n}]$ for some polynomial $F$ of degree $k$ and $G$ of degree $(d-k+2)$ in ${\mathbb{C}}[y_{0},y_{1}]$.
Now, let us show that $\sigma_{k}(X)$ is smooth at $f$ by computing the dimension of conormal space. In general, we have
(16) (${n+d\choose d}-kn-k\geq\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X)~{},$
where the left hand side is given by the expected codimension of the conormal space. By Proposition 23, we also have
(17) $\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\geq\dim_{\mathbb{C}}(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}~{}.$
Thus, $f$ is a smooth point of $\sigma_{k}(X)$ if the lower bound for the dimension of conormal space in (17) is equal to the expected codimension in (16).
Since $k\leq\frac{d+1}{2}$ by the assumption, note that $d-k\geq k$ unless $d$ is odd and $k=\frac{d+1}{2}$, in which $d-k=\frac{d-1}{2}<k$.
a) $d$: odd and $k=\frac{d+1}{2}$ case : In this case, we have
$(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}$ $=\big{(}F,y_{2},\ldots,y_{n}\big{)}_{k}\cdot\big{(}y_{2},\ldots,y_{n}\big{)}_{d-k}$
[…]
So, we obtain
$\dim\hat{N}^{\ast}_{[f]}\sigma_{k}(X)$ $\geq\dim_{\mathbb{C}}(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}$
=($={n+d\choose d}-(d+1)-d(n-1)+(n-1)(d-k)\quad(\textrm{note that $k=\frac{d+1}{2}$})$
=($={n+d\choose d}-kn-k~{},$
which tells us that $\sigma_{k}(X)$ is smooth at $f$.
b) $d$: even or $d$: odd and $k<\frac{d+1}{2}$ case : In the second case, we have $k\leq d-k$ and
$(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}$ $=\big{(}F,y_{2},\ldots,y_{n}\big{)}_{k}\cdot\big{(}F,y_{2},\ldots,y_{n}\big{)}_{d-k}$
[…]
Thus, by a similar dimension counting as case a), we see that
$\dim\hat{N}^{\ast}_{[f]}\sigma_{k}(X)$ ($\geq{n+d\choose d}-(d+1)-d(n-1)+(n-1)(d-k)+(d-2k+1)$
=($={n+d\choose d}-kn-k~{},$
which coincides with the expected codimension as desired. Thus, $f$ is a smooth point of $\sigma_{k}(X)$.
(ii) First note that $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-1,d\right\}$ and the incidence $I$ has dimension $12k-$ 1. In the case $d\leq 2k-2$, each fiber of $p:I\rightarrow{\mathbb{P}}^{d}$ is of dimension $\geq 1$, and then $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$ with $1\leq i\leq k$ in the incidence (5) in §2.2.
Now, let $n\geq 3,k=3$ or $n\geq 2,k\geq 4$ and $d=2k-2$. Suppose $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$. Then a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{1}))={\mathbb{P}}^{d}$ is a smooth point of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Since $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$, it follows from Lemma 12 that
(18) $\langle\,\bigcup_{x\in v_{d}({\mathbb{P}}^{1})}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
Taking $m=1$ and $A={\mathbb{P}}^{1}$ in Proposition 22, the number (14), a lower bound for dimension of left hand side of (18), is equal to $dn$. Thus we have
$(2k-2)n=dn\leq k(n+1)-1,$
which is equivalent to the formula $n\leq(k-1)/(k-2)$. It follows that $n\leq 2$ if $k=3$, and $n=1$ if $k\geq 4$, contrary to our assumption.
Finally, since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subsetneq\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ when $d\geq 2k-2$ (note that $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=2k-3<d$), the $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ is a non-trivial singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which means $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, by Lemma 10.
(iii) By assumption, $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-3,d\right\}=d$, that is to say, $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle={\mathbb{P}}^{d}$; hence the assertion follows.
(iv) For $(n,k)=(2,3)$, smoothness of all points in $\sigma_{3}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{2}(v_{d}({\mathbb{P}}^{2}))$ for $d\geq 4$ was already proved in [16, theorem 2.14]. This is included for completeness of the statement. ∎
Remark 24.
(iv) is the exception to the trichotomy in Theorem 1. Under the condition $(k,d,m,n)=(3,4,1,2)$ of (iv), the arithmetic deduced from the containment assumption (18) of moving tangents in the proof does not make any contradiction. The situation is similar in the other exceptional case to the trichotomy, $(k,d,m,n)=(4,3,2,3)$ (Part (iv) of Theorem 2).
4. Proof of main results
In this section, we prove Theorem 2 and Theorem 3. We will discuss the non-singularity result first and results for the singular loci next.
4.1. Generic smoothness
We begin with a lemma which deals with a secant fiber of a general point in an $m$-subsecant variety $v_{d}({\mathbb{P}}^{m})$ in $v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$.
Lemma 25.
Assume $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$, $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={mk+k-1}$, and $k\leq\tbinom{m+d-1}{m}$. Fix $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ to be an $m$-plane, and take $a\in\sigma_{k}(v_{d}(L))$ to be a general point. Then, in the incidence $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$ with the first projection $p:I\rightarrow{\mathbb{P}}^{N}$ as (3) in §2, we have the following inclusion scheme-theoretically,
$p^{-1}(a)\subset\left\{a\right\}\times(L)^{k}.$
Proof.
Consider any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in p^{-1}(a)\subset I$. Let $I_{(m)}\subset I\subset{\mathbb{P}}^{N}\times(L)^{k}$ be the another incidence as in Lemma 9. Since $a\in\sigma_{k}(v_{d}(L))$ is general, it follows $a\notin\sigma_{k-1}(v_{d}(L))$ and $a\notin p(I_{(m)}\setminus I_{(m)}^{0})$ by Remark 7 (b). Since $\dim\sigma_{k}(v_{d}(L))={mk+k-1}$, the secant fiber of $I_{(m)}\rightarrow{\mathbb{P}}^{N}$ at $a$ also consists of finite points. So, by Lemma 9, we have $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I^{0}$. From Lemma 10, it is also true that $a\notin\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
We may write $a=\sum_{i=1}^{k}c_{i}x_{i}$ for some $c_{i}\in{\mathbb{C}}$ as regarding $a$ and $x_{i}=v_{d}(x_{i}^{\prime})$ as vectors in affine space ${\mathbb{C}}^{N+1}$. As in Remark 8 (a), we set $y_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:0:\dots:0]$. Then $a=\sum_{i=1}^{k}c_{i}y_{i}$, where it holds that $c_{i}\neq 0$ and $y_{i}^{\prime}\neq 0$ for all $1\leq i\leq k$. We may assume $x_{i}^{\prime}\in\left\{t_{0}\neq 0\right\}$ for all $i$.
Since $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized by $\mathsf{mono}[t,u]_{\leq d}$, and $a\in\langle\,v_{d}(L)\,\rangle$, it holds that
$0=\sum_{i=1}^{k}c_{i}\cdot\{(u_{1}\cdot\mathsf{mono}[t]_{\leq d-1})(x_{i}^{\prime})\}=\sum_{i=1}^{k}c_{i}\cdot x_{i,m+1}^{\prime}\cdot\{\mathsf{mono}[t]_{\leq d-1}(y_{i}^{\prime})\}~{},$
where $\{Mono(pt)\}$ means the vector obtained by evaluating monomials in $Mono$ at the value of $pt$.
Applying Remark 7 (b) to $\sigma_{k}(v_{d}(L))$, we have
$\dim(\langle\,v_{d-1}(y_{1}^{\prime}),\dots,v_{d-1}(y_{k}^{\prime})\,\rangle)=k-1.$
It means that $c_{i}\cdot x_{i,m+1}^{\prime}=0$, thus $x_{i,m+1}^{\prime}=0$ for all $1\leq i\leq k$. Similarly, we can obtain $x_{i,j}^{\prime}=0$ for all $1\leq i\leq k$ and $j>m$, which give the linear defining equations for $(L)^{k}$ in $({\mathbb{P}}^{n})^{k}$. Hence, scheme-theoretically, it follows that $x_{1}^{\prime},\dots,x_{k}^{\prime}\in L$. ∎
Remark 26.
We recall some known results on $k$-the secant variety and its incidence in terms of $k$-fold symmetric product of ${\mathbb{P}}^{n}$.
  1. (a)
    It is known that $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ is non-singular at $(x_{1}^{\prime},x_{2}^{\prime},\dots,x_{k}^{\prime})$ if $x_{i}^{\prime}\neq x_{j}^{\prime}$ whenever $i\neq j$. Thus, the subset of all distinct $k$-points of ${\mathbb{P}}^{n}$ is a smooth open subscheme of $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ (see e.g. [6, lemma 7.1.4]). Then we also consider the incidence variety in this setting as
    $\widetilde{I}\subset{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$
    , where $\widetilde{I}$ corresponds to $I$ in (3) under the natural map ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ and $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\subset{\mathbb{P}}^{N}$ is the first projection.
  2. (b)
    Assume $k(n+1)<\binom{n+d}{n}$. Then, from [9, theorem 1.1] we know that the projection $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is birational except $(k,d,n)=(9,6,2),(8,4,3),(9,3,5)$, because it is a dominant and generically injective morphism.
Now, we are ready to prove Part (i) of Theorem 2 and Theorem 3.
Proof of Part (i) of Theorem 2 and Part (i) of Theorem 3.
For an $m$-plane $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$, we take the $m$-subsecant variety $Z=\sigma_{k}(v_{d}(L))$ of $Y=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. By assumptions of Theorems 2 and 3, we know that $\dim Y={nk+k-1}$ and $\dim Z={mk+k-1}\leq\dim\langle Z\rangle=\tbinom{m+d}{m}-1$ (i.e. $Z$ is non-defective) by Alexander-Hirschowitz theorem and $Z$ is also identifiable at its general point by [9, theorem 1.1]. Further, $(k,d,m)\neq(9,6,2),(8,4,3),(9,3,5)$ and note that $k\leq\tbinom{m+d-1}{m}$ for each $k$ of this range.
Let $a\in Z$ be a general point and consider $\widetilde{p}:\widetilde{I}\rightarrow Y$. By Lemma 25 and Remark 26 (b), we may assume that the scheme-theoretic fiber $\widetilde{p}^{-1}(a)$ is one point $\mathbf{x}=(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$ in $\widetilde{I}\cap({\mathbb{P}}^{N}\times\operatorname{Sym}^{k}(L))$ and $\mathbf{x}$ is a non-singular point in $\widetilde{I}$, because $\mathbf{x}$ is contained in a smooth Zariski open subset of $\widetilde{I}$ (i.e. a projective bundle over a smooth open base, see Remarks 7 (a) and 26 (a)).
Now, we restrict the projective morphism $\widetilde{p}:\widetilde{I}\rightarrow Y$ onto a non-empty affine open neighborhood $a\in W=\operatorname{Spec}A\subset Y$ and another open subset $\mathbf{x}\in U=\operatorname{Spec}B\subset\widetilde{I}$ and denote the injective ring homomorphism corresponding to $\widetilde{p}|_{U}:U\to W$ by $f:A\hookrightarrow B$ (let also $m_{a}$ (resp. $m_{\mathbf{x}}$) be the maximal ideal of $a$ in $A$ (resp. of $\mathbf{x}$ in $B$)). Note that we may take $U$ so that $\widetilde{p}|_{U}:U\to W$ is a finite morphism (cf. [18, chap. II, ex.3.22 (d)]).
Since $A/m_{a}\simeq B/m_{\mathbf{x}}\simeq{\mathbb{C}}$ and $\tilde{p}^{-1}(a)\simeq\operatorname{Spec}(B\otimes_{A}A/m_{a})\simeq\operatorname{Spec}(B/f(m_{a})B)$ is isomorphic to $\operatorname{Spec}{A/m_{a}}$, we have $f(m_{a})=m_{\mathbf{x}}$ in $B$ and $m_{a}B_{m_{a}}=m_{\mathbf{x}}B_{m_{a}}$ in the local ring $B_{m_{a}}$. We also know $A_{m_{a}}+m_{a}B_{m_{a}}/m_{a}B_{m_{a}}\simeq A_{m_{a}}/m_{a}A_{m_{a}}\simeq B_{m_{a}}/m_{a}B_{m_{a}}$, thus as $A_{m_{a}}$-module $A_{m_{a}}=B_{m_{a}}$ holds by Nakayama lemma. Note that $A_{m_{a}}=B_{m_{a}}$ is also true as a local ring. Finally, since we have $m_{a}A_{m_{a}}=m_{a}B_{m_{a}}=m_{\mathbf{x}}B_{m_{a}}$, we obtain the desired local ring isomorphism
$A_{m_{a}}=B_{m_{a}}\simeq(B_{m_{a}})_{m_{a}B_{m_{a}}}=(B_{m_{a}})_{m_{\mathbf{x}}B_{m_{a}}}\simeq B_{m_{\mathbf{x}}},$
which implies that $a$ is a smooth point in $Y$. ∎
We present an example which shows that one can not extend this generic smoothness result to arbitrary point in the locus $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
Example 27 (Singularity can happen at a special point in Part (i) of Theorem 3).
Let $V={\mathbb{C}}\langle x,y,z,w\rangle\supset W={\mathbb{C}}\langle x,y,z\rangle$ and $f=x^{2}y^{2}+z^{4}$ be a form of degree $4$. Then, $f$ represents a point in $\sigma_{4}(v_{4}({\mathbb{P}}W))\setminus\sigma_{3}(v_{4}({\mathbb{P}}V))$. Note that $\operatorname{rank}\phi_{2,2}(f)=4>3$ where $\phi_{a,d-a}:S^{d}V\to S^{a}V\otimes S^{d-a}V$ is the symmetric flattening. Theorem 3 (i) shows that a general form in $\sigma_{4}(v_{4}({\mathbb{P}}W))\setminus\sigma_{3}(v_{4}({\mathbb{P}}V))$ is a smooth point. But, here we show that $f$ is a singular point of $\sigma_{4}(v_{4}({\mathbb{P}}V))$. We know that the form $f_{D}=x^{2}y^{2}$ has Waring rank 3 so that $f_{D}=\ell_{1}^{4}+\ell_{2}^{4}+\ell_{3}^{4}$ for some $\ell_{i}\in{\mathbb{C}}[x,y]_{1}$. By [16, theorem 2.1] $f_{D}$ is also a singular point of $\sigma_{3}(v_{4}({\mathbb{P}}V))$. Since $f\in\langle f_{D},z^{4}\rangle$, by Terracini’s lemma, we see that $\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\subset\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))$ and $\mathbb{T}_{\ell_{i}}v_{4}({\mathbb{P}}V)\subset\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))$ for any $i$. Further, because $\sigma_{3}(v_{4}({\mathbb{P}}^{1}))=\langle v_{4}({\mathbb{P}}^{1})\rangle$ and $f_{D}$ has 1-dimensional secant fiber in its incidence, one can move $\ell_{i}$ along this ${\mathbb{P}}^{1}$. Thus, we have
(19) $\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))\supseteq\langle\bigcup_{\ell_{i}\in{\mathbb{P}}^{1}}\mathbb{T}_{\ell_{i}^{4}}v_{4}({\mathbb{P}}V),\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\rangle~{}.$
Note that using the parameterization (11) we can estimate the dimension of right hand side of (19). Take an affine open subset $\{[1:t:u_{1}:u_{2}]\}$ of ${\mathbb{P}}^{3}$ and (with a change of coordinates) let $z^{4}$ be $[1:0:1:1]$ and $\ell_{i}\in{\mathbb{P}}^{1}$ be reprsented by $[1:t:0:0]$ for $t\in{\mathbb{C}}$. Then, by (12) the embedded tangent space to $v_{4}({\mathbb{P}}V)$ at $[1:t:u_{1}:u_{2}]$ is given as the row span of
$\begin{bmatrix}[ccccc:cccc:cccc:ccccccc]1&t&t^{2}&t^{3}&t^{4}&u_{1}&u_{1}t&u_{1}t^{2}&u_{1}t^{3}&u_{2}&u_{2}t&u_{2}t^{2}&u_{2}t^{3}&u_{1}^{2}&u_{1}^{2}t&\cdots&u_{1}^{3}&\cdots&u_{2}^{2}&\cdots\\ &1&2t&3t^{2}&4t^{3}&&u_{1}&2u_{1}t&3u_{1}t^{2}&&u_{2}&2u_{2}t&3u_{2}t^{2}&&u_{1}^{2}&\cdots&&\cdots&&\cdots\\ &&&&&1&t&t^{2}&t^{3}&&&&&2u_{1}&2u_{1}t&\cdots&3u_{1}^{2}&\cdots&&\cdots\\ &&&&&&&&&1&t&t^{2}&t^{3}&&&\cdots&&\cdots&2u_{2}&\cdots\end{bmatrix}.$
On $[1:t:0:0]$ ($\forall t\in{\mathbb{C}}$), this matrix turns into the shape as
$\begin{bmatrix}[ccccc:cccc:cccc:ccccccc]1&t&t^{2}&t^{3}&t^{4}&&&&&&&&&&&&&&\\ &1&2t&3t^{2}&4t^{3}&&&&&&&&&&&&\mathsf{O}&&&\\ &&&&&1&t&t^{2}&t^{3}&&&&&&&&&&&\\ &&&&&&&&&1&t&t^{2}&t^{3}&&&&&&&\end{bmatrix}$
and at $[1:0:1:1]$ it is equal to
$\begin{bmatrix}[ccccc:cccc:cccc:ccccccc]1&&&&&1&&&&1&&&&1&&&1&\cdots&1&\cdots\\ &1&&&&&1&&&&1&&&&1&&&\cdots&&\cdots\\ &&&&&1&&&&&&&&2&&&3&\cdots&&\cdots\\ &&&&&&&&&1&&&&&&&&\cdots&2&\cdots\end{bmatrix}~{},$
which shows that $\dim\langle\bigcup_{\ell_{i}\in{\mathbb{P}}^{1}}\mathbb{T}_{\ell_{i}^{4}}v_{4}({\mathbb{P}}V)\rangle\geq 12$, $\dim\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)=3$ and $\langle\bigcup_{\ell_{i}\in{\mathbb{P}}^{1}}\mathbb{T}_{\ell_{i}^{4}}v_{4}({\mathbb{P}}V)\rangle\cap\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)=\emptyset$. Thus, by (19) we obtain $\dim\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))\geq 16=12+3+1$, greater than the expect dimension. Hence $f$ is a singular point of $\sigma_{4}(v_{4}({\mathbb{P}}V))$, whereas $\sigma_{4}(v_{4}({\mathbb{P}}V))$ is smooth at a general point of $\sigma_{4}(v_{4}({\mathbb{P}}W))$.
4.2. Singularity
In this subsection we will prove Part (ii) and (iii) both in Theorem 2 and Theorem 3, which show singularity of the $m$-subsecant loci $\Sigma_{k,d}(m)$ in the $k$-secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Since $\Sigma_{k,d}(m)$ is the union of all the $m$-subsecant varieties $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ as (2), it is enough to prove the statements for $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
Proof of Part (ii) of Theorem 2 and Part (ii) of Theorem 3.
As we noted above, here it is enough to show that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for each $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. We will first prove $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ under the condition in Theorem 2 with $(k,d,m)\neq(9,6,2),(9,3,5)$, next for Theorem 3 with $(k,d,m)\neq(8,4,3)$, and finally for the remaining $3$ cases $(k,d,m)=(9,6,2),(9,3,5),(8,4,3)$. Note that in the remaining $3$ cases, $(k,d,m)$ satisfy that $\beta=\binom{m+d}{m}-1\not<km+k-1$ and $k=\bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}-1\bigl{\rceil}$. Basically, we use the same idea for the proof, but a detailed way of estimation will be slightly different according to each case (due to defectivity and identifiability). The non-triviality of the singular locus, i.e. $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ can be directly obtained at the end by Lemma 10.
Let us take a general point $(a,x_{1},\dots,x_{k})$ in the incidence $J$ as (5) for $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta}$ and take an irreducible component $F$ of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$. Then, $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is a general (so, smooth) point in $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$.
Suppose $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$. Then, we may assume that $a$ is also a smooth point in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. In particular, we have
$\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))~{}.$
Terracini’s lemma implies $\mathbb{T}_{x_{i}}v_{d}({\mathbb{P}}^{n})\subset\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ for $i=1,\dots,k$, and Lemma 12 implies $\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ for a general point $x\in q_{i}(F)$. Thus, we have
(20) $\langle\,\;\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\cup\bigcup_{x\in q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\}}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\;\,\rangle\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
First of all, let us consider (ii) of Theorem 2 with $(k,d,m)\neq(9,6,2),(9,3,5)$. Then $\beta=\binom{m+d}{m}-1<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ as in Remark 16. We have ${\mathbb{P}}^{\beta}=\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))$, since $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ fills the whole ${\mathbb{P}}^{\beta}$. It is enough to discuss the following cases: (a1) $(k-1)m+k<\binom{m+d}{m}$; (a2) $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)\neq(3,2)$; (a3) $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)=(3,2)$.
For the case (a1) (i.e., $(k-1)m+k<\binom{m+d}{m}$), by applying Proposition 15 (i) and Proposition 22 to $A=v_{d}^{-1}(q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\})$ in ${\mathbb{P}}^{m}$ and $\Lambda={\mathbb{P}}^{\beta}$, the number (14), which is a lower bound for the dimension of the left hand side in (20), is greater than or equal to $\beta+(n-m)(1+k+(km+k-1)-\beta)$. Thus, from the containment (20), we obtain
$\beta+(n-m)(1+k+(km+k-1)-\beta)\leq kn+k-1~{},$
which implies $(n-m)(1+(km+k-1)-\beta)\leq(km+k-1)-\beta$. It is a contradiction, because $n>m$ and $(km+k-1)-\beta>0$.
Now, assume $(k-1)m+k=\binom{m+d}{m}$ (equivalently, $km+k-1-\beta=m$). Then, in the same way as above, using Propositions 15 (ii) and taking $A={\mathbb{P}}^{m}=v_{d}^{-1}(Z)$ and $\Lambda=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle={\mathbb{P}}^{\beta}$, we have
(21) $\beta+(n-m)(1+\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle)\leq kn+k-1.$
For $(d,m)\neq(3,2)$ (i.e., the case of (a2)), by Propositions 15 (ii), it holds $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle\geq k+m$. Then $(n-m)(m+1)\leq km+k-1-\beta=m$, contrary to $n>m$. For $(d,m)=(3,2)$ (i.e., the case (a3)), we get $\beta=9$, $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle=5$, and $k=\lceil\frac{\binom{m+d}{m}}{m+1}\rceil=4$. The condition $(k,d,n)\neq(4,3,3)$ implies $n\geq 4$. Then we also have a contradiction since (21) does not hold. Hence we show that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
Secondly, let us regard (ii) of Theorem 3. Except $(k,d,m)=(8,4,3)$, this part consists of (b1) $(k,d,m)=(7,3,4),(5,4,2),(9,4,3)$, and $(14,4,4)$ (i.e., the case of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ being defective); (b2) $(k,d,m)=(8,3,4),(6,4,2),(10,4,3)$, and $(15,4,4)$ (i.e., just after defective case). By the same reason, we also have the containment (20) for these cases provided that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
For the defective cases (i.e., the case (b1)), it is known that all the $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s are hypersurfaces in ${\mathbb{P}}^{\beta}$. So, as taking $A=v_{d}^{-1}(q_{i}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$ corresponding to the entry locus of $a$, by Proposition 22 a containment of the same kind as (20) implies
(22) $\beta-1+(n-m)(1+\dim\langle\,v_{d-1}(A)\,\rangle)\leq kn+k-1~{},$
where $\beta$ is equal to $km+k-\delta$ and $\delta$ is the secant defect of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$. (22) is equivalent to
$n-m\leq\frac{\delta}{1+\dim\langle\,v_{d-1}(A)\,\rangle-k}~{}\bigg{(}\leq\frac{\delta}{1+\delta}<1\bigg{)}~{},$
which contradicts $n-m\geq 1$, because $\dim\langle\,v_{d-1}(A)\,\rangle\geq k+\delta$ by Remark 21.
For the case (b2) i.e., just after defective cases $(k,d,m)=(8,3,4),(6,4,2),(10,4,3)$, and $(15,4,4)$, we can also get a contradiction in a similar way. Since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))$ is a hypersurface in ${\mathbb{P}}^{\beta}$ in all these cases, by Lemma 14 we have $q_{i}(F)=v_{d}({\mathbb{P}}^{m})$ for any maximal irreducible component $F$ of $p^{-1}(a)$ so that we can take $A={\mathbb{P}}^{m}$. Further we see that $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is non-defective for any $n>m$ in each case and $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ so that $\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))={\mathbb{P}}^{\beta}$. By Proposition 22, the containment (20) implies
$\beta+(n-m)(1+\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle)=\binom{m+d}{m}-1+(n-m)\binom{m+d-1}{m}\leq kn+k-1~{},$
which fails to hold in each 4 just after defective case above, because $n\geq m+1$. Thus, we obtain $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
For the remaining $(k,d,m)=(9,6,2),(9,3,5),(8,4,3)$, we first note that these are the only generic non-identifiability cases of subgeneric rank in which a generic secant fiber $p^{-1}(a)$ consists of not one, but two points (see [9, theorem 1.1]). Even though they do not satisfy ($\frac{{m+d\choose m}}{m+1}<k<\frac{{m+d\choose m}}{m+1}+1$ as above (so, not having a positive dimensional secant fibers), we can still get a proof by contradiction using another estimation.
Similarly, suppose that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and take a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ so that $a$ is a smooth point both in $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. In each 3 cases of $(k,d,m)$, let $x_{1},\dots,x_{k}$ and $y_{1},\dots,y_{k}$ be the two secant fibers of $a$ (i.e., $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$ and $a\in\langle\,y_{1},\dots,y_{k}\,\rangle$) and let $x_{i}^{\prime}\in{\mathbb{P}}^{m}$ (resp. $y_{j}^{\prime}$) be the preimage of $x_{i}$, that is $v_{d}(x_{i}^{\prime})=x_{i}$, (resp. of $y_{j}$ with $v_{d}(y_{j}^{\prime})=y_{j}$). Now, let us set $A=\{x_{1}^{\prime},\dots,x_{k}^{\prime},y_{1}^{\prime},\dots,y_{k}^{\prime}\}$. Then, similarly we have a containment as before
(23) $\big{\langle}\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\big{\rangle}\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
Since $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is general, we may assume that at least one $k$ points set of the secant fiber (say $x_{1}^{\prime},\dots,x_{k}^{\prime}$) are general points in ${\mathbb{P}}^{m}$. Then $\langle\,v_{d-1,m}(A)\,\rangle$ contains the $(k-1)$-plane $\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$. It follows that $\dim\langle\,v_{d-1,m}(A)\,\rangle\geq k$; otherwise, $\dim\langle\,v_{d-1,m}(A)\,\rangle\leq k-1$ implies $\langle\,v_{d-1,m}(A)\,\rangle=\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$ but by the generalized trisecant lemma ([28, proposition 1.4.3]) $\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle\cap v_{d-1}({\mathbb{P}}^{m})=\{v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\}$ as a scheme, which gives a contradiction.
Again by (23) and Proposition 22, we get
$kn+k-1$ $\geq\dim\langle\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\cup v_{d}(A)\big{\rangle}+(n-m)\{1+\dim\langle v_{d-1,m}(A)\rangle\}$
$\geq\dim\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{m}))+(n-m)(1+k)=km+k-1+(n-m)(1+k)$
$=kn+k+(n-m-1)~{},$
which is a contradiction since $n-m-1\geq 0$. Thus, it also holds that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ in these 3 generic non-identifiable cases.
Finally, since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ for the $k$’s of the range in this Part (ii), the $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is a non-trivial singular locus, which means $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, by Lemma 10. ∎
We finish this section as proving Part (iii) of Theorems 2 and 3 and Part (iv) of Theorem 2.
Proof of Part (iii) of Theorem 2 and Part (iii) of Theorem 3.
By the conditions in the Part (iii) of these two Theorems, we see that $k-1\geq\lceil\frac{{m+d\choose m}}{m+1}\rceil$ if $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is never defective or $k-1\geq\lceil\frac{{m+d\choose m}}{m+1}\rceil+1$ else if $(d,m)$ appears in the defective list of Alexander-Hiroschowitz. In any case, we have $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle$. Hence, $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ and the assertion follows. ∎
Proof of Part (iv) of Theorem 2.
It is shown in [15] by giving the defining equations of $\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ explicitly. ∎
5. Case of 4-th secant variety of Veronese embedding
In this section we aim to prove Theorem 5 as an investigation of the singular loci of the fourth secant variety (i.e. $k=4$) of any Veronese variety. This theorem consists of one part dealing with (non-)singularity of points in full-secant loci (i.e. $m=3$) and the other part for points in the maximum subsecant loci $\Sigma_{4,d}(\min\{k-1,n\}-1)$. So, we will obtain Theorem 5 by proving Theorem 29 (Part (i) of Theorem 5) and Corollary 30 (Part (ii) and (iii) of Theorem 5).
5.1. Equations by Young flattening
In [23] another source of equations for secant varieties of Veronese was introduced via so-called Young flattening. Here we briefly review construction of a certain type of Young flattening and exploit it for computation of conormal space of a given form.
Let $V={\mathbb{C}}^{n+1}$ and $d=d_{1}+d_{2}+1$. For $1\leq a\leq n$, we consider a map
$\mathrm{YF}^{a}_{d_{1},d_{2},n}:S^{d}V\rightarrow S^{d_{1}}V\otimes S^{d_{2}}V\otimes\bigwedge^{a}V^{\ast}\otimes\bigwedge^{a+1}V$
which is obtained by first embedding $S^{d}V\hookrightarrow S^{d_{1}}V\otimes S^{d_{2}}V\otimes V$ via co-multiplication, then tensoring with $\mathrm{Id}\in\bigwedge^{a}V\otimes\bigwedge^{a}V^{\ast}$, and finally skew-symmetrizing and permuting.
For any $f\in S^{d}V$, we identify $\mathrm{YF}^{a}_{d_{1},d_{2},n}(f)\in S^{d_{1}}V\otimes S^{d_{2}}V\otimes\bigwedge^{a}V^{\ast}\otimes\bigwedge^{a+1}V$ as a linear map
(24) $S^{d_{1}}V^{*}\otimes\bigwedge^{a}V\rightarrow S^{d_{2}}V\otimes\bigwedge^{a+1}V.$
Let $\alpha_{1},\dots,\alpha_{\binom{n+1}{a}}$ give a basis of $\bigwedge^{a}V$. For a decomposable $w^{d}\in S^{d}V$, $\mathrm{YF}^{a}_{d_{1},d_{2},n}$ maps as
$w^{d}\mapsto\frac{d!}{d_{1}!d_{2}!}w^{d_{1}}\otimes w^{d_{2}}\otimes\big{(}\sum_{I}\alpha_{I}^{\ast}\otimes(\alpha_{I}\wedge w)\big{)}$
and if we take $z_{0},\dots,z_{n}$, a basis of $V$ (now, $w=\sum c_{j}z_{j}\in V$ for some $c_{j}$’s and $\alpha_{I}=z_{i_{1}}\wedge\cdots\wedge z_{i_{a}}$ for some distinct $i_{1},\ldots,i_{a}$), then we have
$\mathrm{YF}^{a}_{d_{1},d_{2},n}(w^{d})={\small\frac{d!}{d_{1}!d_{2}!}}\sum^{n}_{j=0}c_{j}\sum_{i_{1},\dots,i_{a}\neq j}w^{d_{1}}\otimes w^{d_{2}}\otimes(z_{i_{1}}\wedge\dots\wedge z_{i_{a}})^{\ast}\otimes(z_{i_{1}}\wedge\dots\wedge z_{i_{a}}\wedge z_{j}),$
which shows $\mathrm{YF}^{a}_{d_{1},d_{2},n}(w^{d})$ has rank $\binom{n}{a}$ as the linear map (note that the rank does not depend on the choice of $w$ and just consider the case $w=z_{0}$). Further, for $k\leq\binom{n+d^{\prime}}{d^{\prime}}$ with $d^{\prime}=min\left\{d_{1},d_{2}\right\}$ it is also immediate to see that $\operatorname{rank}(\mathrm{YF}^{a}_{d_{1},d_{2},n}(f))=k\binom{n}{a}$ for a general $k$ sum of $d$-th power $f=\sum_{i=1}^{k}w_{i}^{d}$.
Thus, from $k\binom{n}{a}+1$ minors of the matrix $\mathrm{YF}^{a}_{d_{1},d_{2},n}(f)$ we obtain a set of equations for $\sigma_{k}(v_{d}({\mathbb{P}}V))$ for this range of $k$ (for some values of $k,d,d^{\prime},a$, it is known that these minors cut $\sigma_{k}(v_{d}({\mathbb{P}}V))$ as an irreducible component (see [23, theorem 1.2.3])).
We can also use this Young flattening to compute conormal space of secant varieties of Veronese.
Proposition 28.
Let $V={\mathbb{C}}^{n+1}$ and $f$ be any (closed) point of $\sigma_{k}(v_{d}({\mathbb{P}}V))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}V))$ in ${\mathbb{P}}S^{d}V$. Suppose $\mathrm{YF}_{d_{1},d_{2},n}^{a}(f)$ has rank $k{n\choose a}$ as a linear map in $\operatorname{Hom}(S^{d_{1}}V^{\ast}\otimes\bigwedge^{a}V,S^{d_{2}}V\otimes\bigwedge^{a+1}V)$. Then, we have
$\hat{N}^{\ast}_{f}\sigma_{k}(v_{d}({\mathbb{P}}V))\supseteq(\ker\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))\cdot(\operatorname{im}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))^{\perp}~{},$
where the right hand side is to be understood as the image of the multiplication $S^{d_{1}}V^{\ast}\otimes\bigwedge^{a}V\otimes S^{d_{2}}V^{\ast}\otimes\bigwedge^{a+1}V^{\ast}\to S^{d}V^{\ast}$.
Proof.
This proposition comes directly from the same idea as Proposition 23 by applying it to a linear embedding $S^{d}V\hookrightarrow S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V$. Since $\operatorname{rank}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f)=k{n\choose a}$ and, as observed before, $v_{d}({\mathbb{P}}V)$ is contained in $\sigma_{n\choose a}(Seg({\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast})\times{\mathbb{P}}(S^{d_{2}}V\otimes\bigwedge^{a+1}V)))$ in ${\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V)$ and not in the previous secants of the same Segre variety, it is straightforward from the proof of Proposition 23 (i.e. $p={n\choose a}$ case). ∎
5.2. Singularity and non-singularity
Using Proposition 28, we have the non-singularity of $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ at any point outside $\Sigma_{4,d}(2)\cup\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$.
Theorem 29 (from full-secant locus).
Let $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ be the $d$-uple Veronese embedding with $n\geq 3$, $d\geq 3$, and $N={n+d\choose d}-1$. Suppose that $f\in\sigma_{4}(v_{d}({\mathbb{P}}^{n}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ and $f$ does not belong to any $2$-subsecant $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))$ of $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$. Then $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is smooth at every such $f$.
Proof.
First, note that for every $f$ in the statement there exists a unique 4-dimensional subspace $U$ such that $f\in\sigma_{4}(v_{d}({\mathbb{P}}U))$, which is determined by the kernel of the symmetric flattening $\phi_{1,d-1}$. This gives a fibration as
$\pi:\sigma_{4}(v_{d}({\mathbb{P}}^{n}))\setminus\big{(}\Sigma_{4,d}(2;{{\mathbb{P}}^{n}})\cup\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\big{)}\to\operatorname{Gr}(3,{\mathbb{P}}^{n})~{}$
whose fibers $\pi^{-1}({\mathbb{P}}U)$’s are all isomorphic to $\sigma_{4}(v_{d}({\mathbb{P}}U))\setminus(\Sigma_{4,d}(2;{{\mathbb{P}}U})\cup\sigma_{3}(v_{d}({\mathbb{P}}U)))$, where recall that $\Sigma_{4,d}(2;{{\mathbb{P}}^{n}})\subset\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is the maximum subsecant locus, i.e. the union of all $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$’s and $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))$’s in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$. So, we can reduce the proof of theorem to the case of $n=3$.
In case of $n=3$, there is a list of normal forms in $\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\setminus(\Sigma_{4,d}(\min\{k-1,n\}-1)\cup\sigma_{3}(v_{d}({\mathbb{P}}^{3})))$ due to Landsberg-Teitler (see [21, theorem 10.9.3.1]) such as
$f_{1}={x_{0}}^{d}+{x_{1}}^{d}+{x_{2}}^{d}+{x_{3}}^{d}$
$f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$
$f_{3}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d-1}{x_{3}}$
$f_{4}={x_{0}}^{d-2}{x_{1}}^{2}+{x_{0}}^{d-1}x_{2}+{x_{3}}^{d}$
$f_{5}=x_{0}^{d-3}x_{1}^{3}+x_{0}^{d-2}x_{1}x_{2}+x_{0}^{d-1}x_{3}$
Case (i) $f_{1}={x_{0}}^{d}+{x_{1}}^{d}+{x_{2}}^{d}+{x_{3}}^{d}$ (Fermat-type). It is well-known that this Fermat-type $f_{1}$ belongs to an almost transitive $\mathbf{SL}_{4}({\mathbb{C}})$-orbit, which corresponds to a general point of $\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$. Hence, $f_{1}$ is a smooth point of $\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$.
Case (ii) $f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$. Say $U={\mathbb{C}}\langle x_{0},x_{1},x_{2},x_{3}\rangle$. Consider the Young flattening $\mathrm{YF}_{d-2,1,3}^{1}(f_{2})\in S^{d-2}U\otimes U\otimes U^{\ast}\otimes\wedge^{2}U\simeq\operatorname{Hom}(S^{d-2}U^{\ast}\otimes U,U\otimes\wedge^{2}U)$ defined in (24). For simplicity, call this type of Young flattening by $\phi$ throughout the proof. Then, $\phi(f_{2})$ is
$\alpha x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum_{j=0}^{3}y_{j}\otimes x_{j}\wedge x_{1}\big{)}+\beta x_{0}^{d-3}x_{1}\otimes x_{0}\otimes\big{(}\sum_{j=0}^{3}y_{j}\otimes x_{j}\wedge x_{0}\big{)}+\gamma x_{0}^{d-2}\otimes x_{1}\otimes\big{(}\sum_{j=0}^{3}y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+\delta x_{2}^{d-2}\otimes x_{2}\otimes\big{(}\sum_{j=0}^{3}y_{j}\otimes x_{j}\wedge x_{2}\big{)}+\epsilon x_{3}^{d-2}\otimes x_{3}\otimes\big{(}\sum_{j=0}^{3}y_{j}\otimes x_{j}\wedge x_{3}\big{)}$
for some nonzero $\alpha,\beta,\gamma,\delta,\epsilon\in{\mathbb{C}}$. Note that as a linear map $S^{d-2}U^{\ast}\otimes U\to U\otimes\wedge^{2}U$, $\operatorname{rank}\phi(x_{0}^{5})=3$ and $\operatorname{rank}\phi(f_{2})=4\cdot 3=12$. Thus, by Proposition 28 $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ produces a subspace of $\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$.
For $d=3$, the expected dimension of $\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ for the smoothness is (${3+3\choose 3}-16=4$ and the corresponding four can be chosen as $y_{0}y_{2}y_{3},y_{1}^{2}y_{2},y_{1}y_{2}y_{3},y_{1}^{2}y_{3}$ in ${S}^{3}U^{\ast}$ which are given by the product of $\{y_{1}\otimes x_{0},y_{2}\otimes x_{2},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{2})\subset U^{\ast}\otimes U$ and
(25) $\{~{}y_{0}\otimes y_{2}\wedge y_{3},~{}y_{1}\otimes y_{1}\wedge y_{2},~{}y_{1}\otimes y_{1}\wedge y_{3},~{}y_{1}\otimes y_{2}\wedge y_{3},$
$~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{2}\otimes y_{0}\wedge y_{3},~{}y_{2}\otimes y_{1}\wedge y_{3},~{}y_{3}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{2},~{}y_{3}\otimes y_{1}\wedge y_{2}~{}\}~{}$
in $\operatorname{im}\phi(f_{2})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$. So, $\sigma_{4}$ is non-singular at $f_{2}$.
For any $d\geq 4$, in $\ker\phi(f_{2})\subset S^{d-2}U^{\ast}\otimes U$, one can find a subspace generated by $\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\}$, where $J=\langle y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{2}y_{3}\rangle$ is an ideal in ${S}^{\bullet}U^{\ast}$. Also, in $\operatorname{im}\phi(f_{2})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$ there exists the same subspace as in (25). In this case, our $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ contains the subspace of ${S}^{d}U^{\ast}$ generated by
$\{y_{0}^{d-2}y_{2}^{2},y_{0}^{d-2}y_{2}y_{3},y_{0}^{d-2}y_{3}^{2}\}\cup\{y_{0}^{d-3}y_{1}^{2}y_{2},y_{0}^{d-3}y_{1}^{2}y_{3},\ldots,y_{0}^{d-3}y_{3}^{3}\}$
$\cup\{y_{0}^{d-4}y_{1}^{4},y_{0}^{d-4}y_{1}^{3}y_{2},\ldots,y_{0}^{d-4}y_{3}^{4}\}\cup\ldots\cup\{y_{0}^{2}y_{1}^{d-2},y_{0}^{2}y_{1}^{d-3}y_{2},\ldots,y_{0}^{2}y_{3}^{d-2}\}$
$\cup\{y_{0}y_{1}^{d-1},y_{0}y_{1}^{d-2}y_{2},\ldots,y_{0}y_{1}y_{2}^{d-2}\}\cup\{y_{0}y_{2}^{d-2}y_{3},\ldots,y_{0}y_{2}y_{3}^{d-2}\}$
$\cup\{y_{1}^{d},\ldots,y_{1}^{2}y_{2}^{d-2},\ldots,y_{1}^{2}y_{3}^{d-2}\}\cup\{y_{1}y_{2}^{d-2}y_{3},\ldots,y_{1}y_{2}y_{3}^{d-2}\}\cup\{y_{2}^{d-2}y_{3},\ldots,y_{2}^{2}y_{3}^{d-2}\}$
(note that the terms above are listed in the lexicographical order). These monomial generators can be also represented as
($\left(\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{3}^{d-1}\}\right)$
($\cup\left(\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{3}^{d-1},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}y_{3}^{d-1},y_{3}^{d}\}\right)~{},$
which implies that by Proposition 28
$\dim\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$ $\geq\dim(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}~{},$
{($\geq\bigg{\{}{d-1+3\choose 3}-10+{d+2\choose 2}-6\bigg{\}}=\binom{d+3}{3}-16~{}.$
Hence, $f_{2}$ is a smooth point of $\sigma_{4}$.
Case (iii) $f_{3}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d-1}{x_{3}}$. Then, $\phi(f_{3})$ is
$~{}\alpha x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+\beta x_{0}^{d-3}x_{1}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+\gamma x_{0}^{d-2}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+\delta x_{2}^{d-2}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}+\epsilon x_{2}^{d-3}x_{3}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}+\eta x_{2}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}$
for some nonzero $\alpha,\beta,\gamma,\delta,\epsilon,\eta\in{\mathbb{C}}$ so that $\operatorname{rank}\phi(f_{3})=12$. For $d=3$, a subspace $\langle y_{1}\otimes x_{0},y_{3}\otimes x_{2}\rangle$ in $\ker\phi(f_{3})\subset U^{\ast}\otimes U$ and another subspace in $\operatorname{im}\phi(f_{3})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$
(26) $\langle~{}y_{0}\otimes y_{2}\wedge y_{3},~{}y_{1}\otimes y_{1}\wedge y_{2},~{}y_{1}\otimes y_{1}\wedge y_{3},~{}y_{1}\otimes y_{2}\wedge y_{3},$
$~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{3},~{}y_{3}\otimes y_{1}\wedge y_{3}~{}\rangle~{}$
produces a desired 4-dimensional subspace $\langle~{}y_{0}y_{3}^{2},y_{1}^{2}y_{2},y_{1}^{2}y_{3},y_{1}y_{3}^{2}~{}\rangle$ in ${S}^{3}U^{\ast}$, which says that $\sigma_{4}$ is non-singular at $f_{3}$.
Similarly, for the case of $d\geq 4$, $(\ker\phi(f_{3}))\cdot(\operatorname{im}\phi(f_{3}))^{\perp}$ contains a subspace of $\hat{N}^{\ast}_{f_{3}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\subset{S}^{d}U^{\ast}$ which is generated by
($\left(\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{2}^{d-2}y_{3}\}\right)$
($\cup\left(\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{2}^{d-2}y_{3},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}^{d-2}y_{3}^{2},y_{2}^{d-3}y_{3}^{3}\}\right)~{},$
using a subspace $\langle~{}\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\}~{}\rangle$ in $\ker\phi(f_{3})$, where $J$ is an ideal generated by $\{y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{3}^{2}\}$ in ${S}^{\bullet}U^{\ast}$, and the same subspace in $\operatorname{im}\phi(f_{3})^{\perp}$ as (26). Thus, $\dim\hat{N}^{\ast}_{f_{3}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq{d+3\choose 3}-16$, which means that $f_{3}$ is also smooth.
Case (iv) $f_{4}={x_{0}}^{d-2}{x_{1}}^{2}+{x_{0}}^{d-1}x_{2}+{x_{3}}^{d}$. For $d=3$, we have
$\phi(f_{4})$ $=2x_{0}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+2x_{1}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+2x_{1}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+2x_{0}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}+2x_{0}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+2x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+6x_{3}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}$
and $\operatorname{rank}\phi(f_{4})=12$. $\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{2}y_{3}+y_{1}^{2}y_{3},y_{1}y_{2}y_{3},y_{2}^{3},y_{2}^{2}y_{3}\rangle$ which can be spanned by $\{y_{2}\otimes x_{0},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{4})\subset U^{\ast}\otimes U$ and
$\{~{}y_{3}\otimes y_{0}\wedge y_{1},y_{3}\otimes y_{0}\wedge y_{2},-y_{1}\otimes y_{1}\wedge y_{2}+y_{2}\otimes y_{0}\wedge y_{2},-y_{0}\otimes y_{2}\wedge y_{3}+y_{1}\otimes y_{1}\wedge y_{3}\}$
in $\operatorname{im}\phi(f_{4})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$. So, $\sigma_{4}$ is non-singular at $f_{4}$.
For $d\geq 4$, it holds that
$\phi(f_{4})$ $=2x_{0}^{d-2}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+2(d-2)x_{0}^{d-3}x_{1}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}$
$+2(d-2)x_{0}^{d-3}x_{1}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+(d-2)(d-3)x_{0}^{d-4}x_{1}^{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+(d-1)x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}+(d-1)x_{0}^{d-2}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+(d-1)(d-2)x_{0}^{d-3}x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+d(d-1)x_{3}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}~{}.$
In this case, $\operatorname{rank}\phi(f_{4})$ is also $12$ and $\ker\phi(f_{4})$ has a subspace $A_{1}$ which is generated by
$\{~{}\langle y_{0}y_{3},y_{1}y_{2},y_{1}y_{3},y_{2}^{2},y_{2}y_{3}\rangle_{d-2}\otimes x_{i}~{}(i=0,\ldots,3),~{}\langle y_{0}y_{2}\rangle_{d-2}\otimes x_{0},~{}\langle y_{1}^{2}\rangle_{d-2}\otimes x_{0},$
$~{}\langle y_{3}^{2}\rangle_{d-2}\otimes x_{3},~{}\langle-2y_{0}y_{2}+(d-1)y_{1}^{2}\rangle_{d-2}\otimes x_{2}~{}\}$
and $\operatorname{im}\phi(f_{4})^{\perp}$ has a subspace $B_{1}$ spanned by
$\{~{}y_{2}\otimes y_{1}\wedge y_{2},~{}y_{2}\otimes y_{1}\wedge y_{3},~{}y_{2}\otimes y_{2}\wedge y_{3},~{}y_{3}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{2},~{}y_{3}\otimes y_{1}\wedge y_{2},$
$~{}-(d-1)y_{1}\otimes y_{1}\wedge y_{2}+2y_{2}\otimes y_{0}\wedge y_{2},~{}-2y_{0}\otimes y_{2}\wedge y_{3}+(d-1)y_{1}\otimes y_{1}\wedge y_{3}\}~{}.$
Then, one can check that $A_{1}\cdot B_{1}$ produces a subspace of $\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{1}$ generated by 19 quartics
$\{~{}-4\underline{y_{0}^{2}y_{2}^{2}}+(2d+4)y_{0}y_{1}^{2}y_{2},~{}-2\underline{y_{0}^{2}y_{2}y_{3}}+3y_{0}y_{1}^{2}y_{3},~{}\underline{y_{0}^{2}y_{3}^{2}},~{}-2\underline{y_{0}y_{1}y_{2}^{2}}+(d-1)y_{1}^{3}y_{2},~{}-2\underline{y_{0}y_{1}y_{2}y_{3}}+3y_{1}^{3}y_{3},$
$~{}\underline{y_{0}y_{1}y_{3}^{2}},~{}-2\underline{y_{0}y_{2}^{3}}+(d-1)y_{1}^{2}y_{2}^{2},~{}-2\underline{y_{0}y_{2}^{2}y_{3}}+(d-1)y_{1}^{2}y_{2}y_{3},~{}-2\underline{y_{0}y_{2}y_{3}^{2}}+(d-1)y_{1}^{2}y_{3}^{2},~{}\underline{y_{1}^{3}y_{3}},$
$~{}\underline{y_{1}^{2}y_{2}^{2}},~{}\underline{y_{1}^{2}y_{2}y_{3}},~{}\underline{y_{1}^{2}y_{3}^{2}},~{}\underline{y_{1}y_{2}^{3}},~{}\underline{y_{1}y_{2}^{2}y_{3}},~{}\underline{y_{1}y_{2}y_{3}^{2}},~{}\underline{y_{2}^{4}},~{}\underline{y_{2}^{3}y_{3}},~{}\underline{y_{2}^{2}y_{3}^{2}}\}$
(here, the underline means the leading term with respect to the lexicographic order). Say $T={S}^{\bullet}U^{\ast}$. $I_{1}$ has a minimal free resolution as
(27) $0\to T(-7)^{4}\to T(-6)^{22}\to T(-5)^{36}\to T(-4)^{19}\to I\to 0~{},$
which shows that the Hilbert function of $I$ can be computed as
$H(I,d)$ $=19{d-4+3\choose 3}-36{d-5+3\choose 3}+22{d-6+3\choose 3}-4{d-7+3\choose 3}$
=($={d+3\choose 3}-16\quad(d\geq 4)~{}.$
This implies that (${d+3\choose 3}-16\geq\dim\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq H(I,d)={d+3\choose 3}-16$, which means that $\sigma_{4}$ is also smooth at $f_{4}$.
Case (v) The final form $f_{5}=x_{0}^{d-3}x_{1}^{3}+x_{0}^{d-2}x_{1}x_{2}+x_{0}^{d-1}x_{3}$. We begin with $d=3$. We have
$\phi(f_{5})$ $=6x_{1}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+x_{2}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+x_{1}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}+x_{1}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+x_{0}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}$
$+x_{0}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+2x_{3}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+2x_{0}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}$
$+2x_{0}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
and $\operatorname{rank}\phi(f_{5})=12$. The conormal space $\hat{N}^{\ast}_{f_{5}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{3}^{2}+4y_{1}y_{2}y_{3}-24y_{2}^{3},-y_{1}y_{3}^{2}+12y_{2}^{2}y_{3},y_{2}y_{3}^{2},y_{3}^{3}\rangle$ which can be spanned by $\{y_{3}\otimes x_{0},2y_{1}\otimes x_{0}+12y_{2}\otimes x_{1}+y_{3}\otimes x_{2}\}$ in $\ker\phi(f_{5})\subset U^{\ast}\otimes U$ and
(28) $\{~{}y_{3}\otimes y_{1}\wedge y_{2},-2y_{2}\otimes y_{1}\wedge y_{2}+y_{3}\otimes y_{0}\wedge y_{2},-2y_{1}\otimes y_{2}\wedge y_{3}+y_{3}\otimes y_{0}\wedge y_{3}\}$
in $\operatorname{im}\phi(f_{5})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$. So, $\sigma_{4}$ is non-singular at $f_{5}$.
For each $d\geq 4$, the Young flattening is of the form
$\phi(f_{5})$ $=6x_{0}^{d-3}x_{1}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+3(d-3)x_{0}^{d-4}x_{1}^{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}$
$+3(d-3)x_{0}^{d-4}x_{1}^{2}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+(d-3)(d-4)x_{0}^{d-5}x_{1}^{3}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+(d-2)(d-3)x_{0}^{d-4}x_{1}x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+(d-2)x_{0}^{d-3}x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}$
$+(d-2)x_{0}^{d-3}x_{2}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+(d-2)x_{0}^{d-3}x_{1}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}$
$+(d-2)x_{0}^{d-3}x_{1}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+x_{0}^{d-2}\otimes x_{1}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}$
$+x_{0}^{d-2}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+(d-1)(d-2)x_{0}^{d-3}x_{3}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
$+(d-1)x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}+(d-1)x_{0}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}~{}$
and $\operatorname{rank}\phi(f_{5})$ is also $12$. Now, $\ker\phi(f_{5})$ contains a subspace $A_{2}$ which is generated by
$\{~{}\langle~{}y_{1}y_{3},~{}y_{2}^{2},~{}y_{2}y_{3},~{}y_{3}^{2}~{}\rangle_{d-2}\otimes x_{i}~{}(i=0,\ldots,3),~{}\langle-6y_{0}y_{2}+(d-2)y_{1}^{2}\rangle_{d-2}\otimes x_{3},$
$~{}\langle-y_{0}y_{3}+(d-1)y_{1}y_{2}\rangle_{d-2}\otimes x_{3}~{}\}$
and $\operatorname{im}\phi(f_{5})^{\perp}$ has a subspace $B_{2}$ spanned by
$\{~{}y_{2}\otimes y_{2}\wedge y_{3},~{}y_{3}\otimes y_{1}\wedge y_{2},~{}y_{3}\otimes y_{1}\wedge y_{3},~{}y_{3}\otimes y_{2}\wedge y_{3},~{}-y_{1}\otimes y_{2}\wedge y_{3}+y_{2}\otimes y_{1}\wedge y_{3},$
$~{}-(d-1)y_{2}\otimes y_{1}\wedge y_{2}+y_{3}\otimes y_{0}\wedge y_{2},~{}-(d-1)y_{1}\otimes y_{2}\wedge y_{3}+y_{3}\otimes y_{0}\wedge y_{3},$
$~{}-y_{0}\otimes y_{1}\wedge y_{2}+y_{1}\otimes y_{0}\wedge y_{2}-y_{2}\otimes y_{0}\wedge y_{1},~{}-y_{0}\otimes y_{1}\wedge y_{3}+y_{1}\otimes y_{0}\wedge y_{3}-y_{3}\otimes y_{0}\wedge y_{1},$
$~{}-6y_{0}\otimes y_{2}\wedge y_{3}+(d-2)y_{1}\otimes y_{1}\wedge y_{3}-6(d-1)y_{2}\otimes y_{1}\wedge y_{2}~{}\}~{}.$
Then, one can check that $A_{2}\cdot B_{2}$ produces a subspace of $\hat{N}^{\ast}_{f_{5}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{2}$ generated by 19 quartics
$\{~{}36y_{0}^{2}y_{2}^{2}-12(d-2)y_{0}y_{1}^{2}y_{2}+(d-2)^{2}y_{1}^{4},~{}6y_{0}^{2}y_{2}y_{3}-(d-2)y_{0}y_{1}^{2}y_{3}-6(d-1)y_{0}y_{1}y_{2}^{2}+(d-1)(d-2)y_{1}^{3}y_{2},$
$~{}-y_{0}^{2}y_{3}^{2}+2(d-1)y_{0}y_{1}y_{2}y_{3}-(d-1)^{2}y_{1}^{2}y_{2}^{2},~{}-6y_{0}y_{1}y_{2}y_{3}+(d-2)y_{1}^{3}y_{3},~{}-y_{0}y_{1}y_{3}^{2}+(d-1)y_{1}^{2}y_{2}y_{3},$
$~{}-6y_{0}y_{2}^{3}+(d-2)y_{1}^{2}y_{2}^{2},~{}y_{0}y_{2}^{2}y_{3}-(d-1)y_{1}y_{2}^{3},~{}y_{0}y_{2}y_{3}^{2}-(d-1)y_{1}y_{2}^{2}y_{3},~{}y_{0}y_{3}^{3}-(d-1)y_{1}y_{2}y_{3}^{2},$
$~{}(d-2)y_{1}^{2}y_{2}y_{3}-6(d-1)y_{1}y_{2}^{3},~{}y_{1}^{2}y_{3}^{2},~{}y_{1}y_{2}^{2}y_{3},~{}y_{1}y_{2}y_{3}^{2},~{}y_{1}y_{3}^{3},~{}y_{2}^{4},~{}y_{2}^{3}y_{3},y_{2}^{2}y_{3}^{2},~{}y_{2}y_{3}^{3},~{}y_{3}^{4}\}~{}.$
Note that $I_{2}$ has the same minimal free resolution as $I_{1}$ in (27). Therefore, by the same argument, we conclude that $f_{5}$ is also a smooth point when $d\geq 4$. ∎
As a direct consequence of main results in the paper, we also obtain the following corollary on (non-)singularity of subsecant loci in the 4-th secant variety.
Corollary 30 (from subsecant loci).
Let $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ be the $d$-uple Veronese embedding with $n\geq 3$, $d\geq 3$, and $N={n+d\choose d}-1$. Then the following holds.
  1. (i)
    A general point in $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is smooth for $d\geq 4$. For $d=3$, $\sigma_{4}(v_{3}({\mathbb{P}}^{2}))$ is a non-trivial singular locus for any $n\geq 4$, while all points in $\sigma_{4}(v_{3}({\mathbb{P}}^{2}))\setminus\sigma_{3}(v_{3}({\mathbb{P}}^{3}))$ are smooth for $n=3$.
  2. (ii)
    $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\subset\operatorname{Sm}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$ if $d\geq 7$. $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$ is non-trivial singular locus when $d=6$ and $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ when $d\leq 5$.
Proof.
Since $k=4~{}\&~{}n\geq 3$, the relevant range for an $m$-subsecant locus in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is $1\leq m\leq 2$.
(i) For $m=2$, Theorem 2 (ii) tells $\sigma_{4}(v_{d}({\mathbb{P}}^{m}))$ is non-trivial singular locus in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ if $d=3,n\geq 4$. The case $(d,n)=(3,3)$ is also discussed in Theorem 2 (iv).
When $d=4$, we can say that a general point in $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is smooth by Theorem 3 (i). For any $d\geq 5$, the same conclusion comes from Theorem 2 (i).
(ii) This is given by Theorem 1 for the case $k=4,m=1$. ∎
We add some remarks on Corollary 30.
Remark 31.
(a) For $d=2$, a subsecant variety $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))$ in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is a trivial singular locus, because $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))=\sigma_{3}(v_{d}({\mathbb{P}}^{2}))\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$.
(b) As pointed out in Example 27, a singularity can occur at a special point in $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ even for $d\geq 4$.
Finally, we end this section by listing cases in which the same nice description for the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ as in Example 6 can be made.
Corollary 32.
Let $V$ be a $(n+1)$-dimensional complex vector space $(n\geq 1)$ and $v_{d}({\mathbb{P}}V)\subset{\mathbb{P}}^{N}$ be the image of the $d$-uple $(d\geq 2)$ Veronese embedding of ${\mathbb{P}}V$. Assume that $(k,d,n)$ satisfies one of the following conditions:
  1. (i)
    $d=2$, $\forall~{}n\geq k-1$,
  2. (ii)
    $k=2$, $\forall~{}d,n$,
  3. (iii)
    $k=3$, $d=3$ and $\forall~{}n\geq 2$ or $k=3$, $d=4$ and $\forall~{}n\geq 3$,
  4. (iv)
    $k=4$, $d=3$ and $\forall~{}n\geq 4$.
Then the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}V))$ is given exactly as
(29) $\big{\{}f\in{\mathbb{P}}S^{d}V~{}|~{}\textrm{$f$ is any form which can be expressed using at most $k-1$ variables}\big{\}}~{},$
which is an irreducible locus of dimension $(k-1)(n-k+2)+{d+k-2\choose d}-1$ and is equal to the maximum subsecant locus $\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$.
Proof.
For case (i), the assertion is immediate since it corresponds to symmetric matrices. In case (ii), we draw the conclusion from the fact $\operatorname{Sing}(\sigma_{2}(v_{d}({\mathbb{P}}V)))=v_{d}({\mathbb{P}}V)$ for every $d,n$ [20].
For the remaining cases, we first claim that for any $13\leq k\leq n+$ 1 it holds
(30) $\sigma_{k-1}(v_{d}({\mathbb{P}}V))\subset\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle~{}.$
We note that the right hand side of (30) is an irreducible and closed subvariety of $\sigma_{k}(v_{d}({\mathbb{P}}V))$, since it coincides with a subvariety $\bigcup_{\Lambda\in\mathrm{Im}\Phi}\Lambda$ where a map $\Phi:\mathbb{G}(k-2,n)\to\mathbb{G}({d+k-2\choose d}-1,N)$ sending each subspace $L$ of dimension $k-2$ to the linear span $\langle\,v_{d}(L)\,\rangle$ in ${\mathbb{P}}^{N}$ is regular (see e.g. [17, example 6.10, proposition 6.13]). Then, because a general element of the left hand side is of the form $\ell_{1}^{d}+\cdots+\ell_{k-1}^{d}$ for some linear forms $\ell_{i}$’s, it belongs to $\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$ for some ${\mathbb{P}}^{k-2}\subset{\mathbb{P}}V$ so that the closure is also contained in the subvariety $\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$.
For case (iii), by [16, theorem 2.1, remark 2.4(a), and corollary 2.11] and Theorem 1, and for case (iv), by Theorem 5, we know that
$\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}V)))=\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))\big{\}}~{},$
which can be also written as $\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$. In both case (iii) and (iv), we have $\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))=\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$. Thus, by the claim above, the singular locus is equal to $\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle=\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$, which is irreducible and can be described as written in the statement. The formula for dimension is immediate from dimension counting. ∎
6. Concluding remark
So far, we reported results on singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ coming from the subsecant loci. To the best of our knowledge, there has been no general idea or clear consensus on the singular locus of arbitrary higher secant variety of any Veronese variety yet. From this point of view, the present paper contributes to it as making a more visible picture on the singular locus via showing a generic smoothness of the subsecant loci for relatively low $k$’s and confirming singularity of the same loci for other $k$’s.
As we mentioned in the introduction, each point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ is located in $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for some $1\leq m\leq\min\{k-1,n\}$. To make the picture more complete, we have two future issues: (i) On the subsecant loci (i.e. $m<\min\{k-1,n\}$), one needs to check (non-)singularity not only at a general point but also at every point and (ii) Points in the full-secant locus (i.e. $m=\min\{k-1,n\}$) should be treated.
Issue (i) is expected to be very complicated, because at some special point a singularity can also happen even for a low $k$ as shown in Example 27 (in fact, we can generate more examples using a similar idea). For the points in the subsecant loci, in general, one could not hope to find some nice ‘normal forms’ and the situation is expected to be wild (in other words, the subsecant loci may not be covered with finitely many nice families of $\mathbf{SL}$-orbits). But, still we can push on our viewpoint a bit further and, along the same spirit, we can refine a main result of this paper in the following manner. Based on the singularity results in Theorems 2 and 3 and using a similar estimation as §2.3, more generally we have:
Theorem 33.
Suppose that $k>\mu$ in case of Theorem 2 or $k\geq\lfloor\mu\rfloor$ in case of Theorem 3 (in other words, the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is a singular locus in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$). Let $r\leq n-m$. Then, the following does hold
(31) $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))\subset\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n})))~{},$
where $J(X,Y)$ denotes the ‘(embedded) join’ of two subvarieties $X,Y$ in some ${\mathbb{P}}^{N}$.
Proof.
Suppose that the inclusion (31) does not hold. Then, taking $x_{1},\dots,x_{k}$ to be general points of $v_{d}({\mathbb{P}}^{m})$ and $x_{k+1},\dots,x_{k+r}$ to be general points of $v_{d}({\mathbb{P}}^{n})$, we may assume that $x\notin\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n})))$ for a general $x\in\langle\,x_{1},\dots,x_{k+r}\,\rangle$. By Terracini lemma, we have
(32) $L_{1}:=\langle\,\mathbb{T}_{x_{1}}v_{d}({\mathbb{P}}^{n}),\dots,\mathbb{T}_{x_{k}}v_{d}({\mathbb{P}}^{n})\,\rangle\subset\mathbb{T}_{x}\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$
and by the assumption on $k$ we know $\dim L_{1}>kn+k-1$.
On the other hand, since $x_{k+1},\dots,x_{k+r}$ are general points of $v_{d}({\mathbb{P}}^{n})$,
$L_{2}=\langle\,\mathbb{T}_{x_{k+1}}v_{d}({\mathbb{P}}^{n}),\dots,\mathbb{T}_{x_{k+r}}v_{d}({\mathbb{P}}^{n})\,\rangle\subset\mathbb{T}_{x}\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$
and $L_{2}$ has dimension at least $rn+r-1$.
Moreover, we may assume $L_{1}\cap L_{2}=\emptyset$ as follows. Taking ${\mathbb{P}}^{n-m-1}\subset{\mathbb{P}}^{n}$ such that $x_{k+1},\dots,x_{k+r}\in{\mathbb{P}}^{r-1}\subset{\mathbb{P}}^{n-m-1}$ and ${\mathbb{P}}^{m}\cap{\mathbb{P}}^{n-m-1}=\emptyset$, and changing coordinates $t_{0},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ as in §2.3, we may say that ${\mathbb{P}}^{n-m-1}$ is the zero set of $t_{0}=\dots=t_{m}=0$ and ${\mathbb{P}}^{m}$ is the zero set of $u_{1}=\dots=u_{m^{\prime}}=0$. For a point $x^{\prime}\in{\mathbb{P}}^{m}$, using the parametrization (11), the tangent space $\mathbb{T}_{v_{d}(x^{\prime})}v_{d}({\mathbb{P}}^{n})$ is spanned by the rows of the matrix of the form $[\ast:\mathsf{O}]$ as (13). On the other hand, for a point $x^{\prime\prime}\in{\mathbb{P}}^{n-m-1}$ and for an affine open set containing $x^{\prime\prime}$ we may take $u_{m^{\prime}}=1$ instead of $t_{0}=1$. Then, the only part in the parametrization of $v_{d}$ which contributes $\mathbb{T}_{v_{d}(x^{\prime\prime})}v_{d}({\mathbb{P}}^{n})$ is
$\ t_{0}\cdot\mathsf{mono}[u]_{\leq d-1},\ \dots,\ t_{m}\cdot\mathsf{mono}[u]_{\leq d-1},\mathsf{mono}[u]_{\leq d}$
which corresponds to the tailing “*” part in (11) (recall that $\mathsf{mono}[u]_{\leq e}$ is the set of monomials of ${\mathbb{C}}[u_{1},\dots,u_{m^{\prime}}]$ of degree $\leq e$). Thus, a similar matrix whose rows span the other tangent space $\mathbb{T}_{v_{d}(x^{\prime\prime})}v_{d}({\mathbb{P}}^{n})$ has a form $[\mathsf{O}:\ast]$. This implies $L_{1}\cap L_{2}=\emptyset$.
Hence $\dim\langle\,L_{1},L_{2}\,\rangle>(k+r)n+(k+r)-1$, which is contrary to $\langle\,L_{1},L_{2}\,\rangle\subset\mathbb{T}_{x}\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$. ∎
Remark 34 (Partial subsecant locus).
This new singular locus $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))$ in (31) can be seen as a ‘partial version’ of subsecant locus in this paper. In particular, it contains $m$-subsecant variety $\sigma_{k+r}(v_{d}({\mathbb{P}}^{m}))=J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{m})))$. So, let us call such a locus a partial subsecant locus of $\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$. We note that the singularity of a specific form $f=x^{2}y^{2}+z^{4}$ in Example 27 can be explained using this notion; $f$ is a point of $\Sigma_{4,4}(2;{\mathbb{P}}^{3})$ where only a generic smoothness is known by Theorems 3 (i), but $f$ also belongs to a partial subsecant locus $J(\sigma_{3}(v_{4}({\mathbb{P}}^{1})),\sigma_{1}(v_{4}({\mathbb{P}}^{3})))$ which is singular by Theorem 33.
Therefore, one proper question on the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ here is probably such as
Question 35.
Let $k-1\leq n$ and $\mathcal{D}$ be the union of all possible (partial) subsecant loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Are the points of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\big{(}\mathcal{D}\cup\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))\big{)}$ all smooth in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$?
Note that the answer to Question 35 is affirmative in cases of $k=2$ classically and $k=3$ (by [16]) and $k=4$ (by Theorem 29). For a large value $k$ compared to $n$ (e.g. $n<k-1$), Question 35 may be answered negatively as in the following example.
Example 36.
Let us consider $\sigma_{14}(v_{8}({\mathbb{P}}^{2}))$, the $14$-th secant variety of the Veronese variety $v_{8}({\mathbb{P}}^{2})$. Take general 14 points on $v_{8}({\mathbb{P}}^{2})$. In [3, remark 4.10], the authors presented a concrete point in the linear span of the 14 points which is a non-normal point to $\sigma_{14}(v_{8}({\mathbb{P}}^{2}))$. Note that one can also check this singular point does not belong to $\mathcal{D}$, the locus of all partial subsecants.
Remark 37.
Finally, we would like to remark that the approach based on the same spirit of trichotomy pattern of (non-)singularity on subsecant loci still can be applied to the study of singular loci of higher secant varieties of other classical varieties such as Segre embeddings, Segre-Veronese varieties and Grassmannians. For instance, we can have a conjectural result like
“ For $\vec{\mathbf{n}}:=(n_{1},n_{2},\ldots,n_{r})$, let $X$ be the Segre embedding ${\mathbb{P}}^{n_{1}}\times{\mathbb{P}}^{n_{2}}\times\cdots\times{\mathbb{P}}^{n_{r}}\subset{\mathbb{P}}^{\prod(n_{i}+1)-1}=:{\mathbb{P}}^{\beta(\vec{\mathbf{n}})}$ and denote $\sigma_{k}(X)$ by $\sigma_{k}(\vec{\mathbf{n}})$, $\dim\sigma_{k}(X)$ by $s_{k}(\vec{\mathbf{n}})$. For every $\vec{\mathbf{m}}=(m_{1},m_{2},\ldots,m_{r})$ with $0\leq m_{i}\leq\min\{k-1,n_{i}-1\}$, besides a few exceptional cases, we have that the following holds
  1. (i)
    $\sigma_{k}(\vec{\mathbf{n}})$ is smooth at a general point in $\sigma_{k}(\vec{\mathbf{m}})\setminus\sigma_{k-1}(\vec{\mathbf{n}})$ if $\beta(\vec{\mathbf{m}})\geq s_{k}(\vec{\mathbf{m}})$
  2. (ii)
    $\sigma_{k}(\vec{\mathbf{m}})$ is singular in $\sigma_{k}(\vec{\mathbf{n}})$, but $\sigma_{k}(\vec{\mathbf{m}})\not\subset\sigma_{k-1}(\vec{\mathbf{n}})$ (i.e., non-trivial singular locus) if $s_{k-1}(\vec{\mathbf{m}})<\beta(\vec{\mathbf{m}})<s_{k}(\vec{\mathbf{m}})$
  3. (iii)
    $\sigma_{k}(\vec{\mathbf{m}})\subset\sigma_{k-1}(\vec{\mathbf{n}})$ if $\beta(\vec{\mathbf{m}})\leq s_{k-1}(\vec{\mathbf{m}})$ ” ,
which can recover the result on the singular locus of the secant varieties of Segre embeddings ([27, corollary 7.17]) for $k=2$. Note that, if we assume that everything is non-defective, then the ranges above can be computed as
[…]
We plan to deal with these cases in a forthcoming paper.

### v3

@@
 Abstract.
-In this paper we study singular loci of higher secant varieties of the image of the $d$-uple Veronese embedding of projective $n$-space, $v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N={n+d\choose d}-1}$.
-We call the closure of the union of all the $(k-1)$-planes spanned by $k$ points on $X$ the $k$-th secant of $X$ for a given embedded variety $X\subset{\mathbb{P}}^{N}$ and denote it by $\sigma_{k}(X)$.
-For the singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, it has been known only for $k\leq 3$.
-As investigating geometry of moving tangents along subvarieties, we determine the (non-)singularity of so-called the $m$-subsecant locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, consisting of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ with any $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$.
-These loci with arbitrary $k,d,m,n$ show an interesting trichotomy between generic smoothness, non-trivial singularity and trivial singularity.
-In many cases, they can be served as a new source for the singular locus of the $k$-th secant other than a trivial one, the $(k-1)$-th secant variety.
-We also consider the case of the $4$-th secant of $v_{d}({\mathbb{P}}^{n})$ by applying main results and computing conormal space via a certain type of Young flattening.
-In the end, some generalization and discussions for further development are presented as well.
+The $k$-th secant variety of a projective variety $X\subset\mathbb{P}^{N}$, denoted by $\sigma_{k}(X)$, is defined to be the closure of the union of $(k-1)$-planes spanned by $k$ points on $X$.
+In this paper, we examine the $k$-th secant variety $\sigma_{k}(v_{d}(\mathbb{P}^{n}))\subset\mathbb{P}^{N}$ of the image of the $d$-uple Veronese embedding $v_{d}$ of $\mathbb{P}^{n}$ to $\mathbb{P}^{N}$ with $N=\binom{n+d}{d}-1$, and focus on the singular locus of $\sigma_{k}(v_{d}(\mathbb{P}^{n}))$, which is only known for $k\leq 3$.
+To study the singularity for arbitrary $k,d,n$, we define the $m$-subsecant locus of $\sigma_{k}(v_{d}(\mathbb{P}^{n}))$ to be the union of $\sigma_{k}(v_{d}(\mathbb{P}^{m}))$ with any $m$-plane $\mathbb{P}^{m}\subset\mathbb{P}^{n}$.
+By investigating the projective geometry of moving embedded tangent spaces along subvarieties and using known results on the secant defectivity and the identifiability of symmetric tensors, we determine whether the $m$-subsecant locus is contained in the singular locus of $\sigma_{k}(v_{d}(\mathbb{P}^{n}))$ or not.
+Depending on the value of $k$, these subsecant loci show an interesting trichotomy between generic smoothness, non-trivial singularity, and trivial singularity.
+In many cases, they can be used as a new source for the singularity of the $k$-th secant variety of $v_{d}(\mathbb{P}^{n})$ other than the trivial one, the $(k-1)$-th secant variety of $v_{d}(\mathbb{P}^{n})$.
+We also consider the case of the $4$-th secant variety of $v_{d}(\mathbb{P}^{n})$ by applying main results and computing conormal space via a certain type of Young flattening.
+Finally, we present some generalizations and discussions for further developments.
 Key words and phrases:
-singular locus of higher secant varieties, the subsecant loci, Veronese embedding, symmetric tensors, identifiability
+singular locus of higher secant varieties, the subsecant loci, Veronese embedding, symmetric tensors, secant defectivity, identifiability
 2010 Mathematics Subject Classification:
@@
 Throughout the paper, we work over ${\mathbb{C}}$, the field of complex numbers. Let $X\subset{\mathbb{P}}^{N}$ be an embedded projective variety. The $k$-th secant variety of $X$ is defined as:
-$\sigma_{k}(X)=\overline{\bigcup_{x_{1},\dots,x_{k}\in X}\langle\,x_{1},\dots,x_{k}\,\rangle}\subset{\mathbb{P}}^{N}~{},$
+(1) $\sigma_{k}(X)=\overline{\bigcup_{x_{1},\dots,x_{k}\in X}\langle\,x_{1},\dots,x_{k}\,\rangle}\subset{\mathbb{P}}^{N},$
 where $\langle\,x_{1},\dots,x_{k}\,\rangle\subset{\mathbb{P}}^{N}$ denotes the linear span of the points $x_{1},\dots,x_{k}$ and the overline means the Zariski closure.
-In particular, $\sigma_{1}(X)=X$ and $\sigma_{2}(X)$ has been often called just the secant or secant line variety of $X$ in the literature.
-The secant variety construction (more generally, join construction of subvarieties) is not only one of the most famous methods in classical algebraic geometry, but also a very popular object to be considered in recent years, especially in connection with fields of research such as tensor rank and decomposition, algebraic statistics, data science, geometric complexity theory and so on (see [21, 22] for more details).
-Despite of a rather long history and its popularity, most of fundamental questions on the higher secant varieties $\sigma_{k}(X)$ still remain open even for many well-known base variety $X$.
-For instance, the dimension theorem for $\sigma_{k}(X)$ has been only settled completely for $X$ being curves, Veronese varieties or few other cases.
-Questions about defining equations of $k$-secant variety also have been answered only for very small $k$’s of few cases and seem far from understanding the essence of sources for the equations (see also [21, chapter 5] for a reference).
-In this paper we concentrate on the case of $X=v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$, the Veronese variety, which is the image of the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\hookrightarrow{\mathbb{P}}^{N}$ with $N={n+d\choose d}-1$.
-In particular, we study the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, arbitrary higher secant variety of the Veronese variety.
-Knowledge on singularities of higher secant varieties is fundamental and very important for its own sake in the study of algebraic geometry and also can be useful for problems in applications.
-For example, this can be served as a key condition to determine identifiability of structured tensors (see e.g.
-the introduction in [8] and references therein).
+In particular, $\sigma_{1}(X)=X$ and $\sigma_{2}(X)$ is often simply called the secant or secant line variety of $X$ in the literature.
+The construction of secant varieties (or more generally, join construction of subvarieties) is not only one of the most famous methods in classical algebraic geometry, but also a very popular subject in recent years, especially in connection with fields of research such as tensor rank and decomposition, algebraic statistics, data science, geometric complexity theory and so on (see [Land12, Land17] for more details).
+Despite of a rather long history and the popularity, most of the fundamental questions on the higher secant varieties $\sigma_{k}(X)$ still remain open even for many well-known base varieties $X$.
+For instance, one can ask the secant defectivity question, which concerns the dimension of $\sigma_{k}(X)$.
+We say that $\sigma_{k}(X)$ is secant defective (or simply defective) if the dimension of $\sigma_{k}(X)$ is less than the expected one, $\min\{N,k\cdot\dim(X)+k-1\}$.
+It is classically known that higher secant varieties of curves are non-defective (e.g.
+[Rus16, Corollary 1.2.3]).
+Due to the famous theorem of Alexander-Hirschowitz [AH95], we know the dimensions of higher secant varieties of all Veronese varieties.
+In other research, there are only a few cases where the dimension theorem for $\sigma_{k}(X)$ is fully determined (see [Rus16], [Zak93] for references).
+Questions about defining equations of $k$-th secant varieties have also only been answered for very small $k$’s of a few cases and seem still far from understanding the essence of the sources for the equations (see also [Land12, Chapter 5] for a reference).
+In this paper, we concentrate on the case of $X=v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$, the Veronese variety, which is the image of the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\hookrightarrow{\mathbb{P}}^{N}$ with $N={n+d\choose d}-1$.
+In particular, we study the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, an arbitrary higher secant variety of the Veronese variety.
+The knowledge on singularities of higher secant varieties is fundamental and very important for its own sake in the study of algebraic geometry and also can be useful for problems in applications.
+For example, it can be used as a key condition to establish the identifiability of structured tensors (e.g.
+the introduction in [COV14] and references therein).
 For any irreducible variety $X\subset{\mathbb{P}}^{N}$, it is classically well-known that
 (2) $\sigma_{k-1}(X)\subset\operatorname{Sing}(\sigma_{k}(X))$
-unless $\sigma_{k}(X)$ fills the whole linear span $\langle\,X\,\rangle$.
-In the paper, we call a point $p\in\sigma_{k}(X)$ a non-trivial singular point if $p\notin\sigma_{k-1}(X)$ and $\sigma_{k}(X)$ is singular at $p$, while the points belonging to $\sigma_{k-1}(X)$ are called trivial singular points.
-There have been some known results on the singular loci of $\sigma_{k}(X)$ mostly for very small $k$’s.
+unless $\sigma_{k}(X)$ fills up the whole linear span $\langle\,X\,\rangle$ (see [Rus16, Proposition 1.2.2]).
+In the paper, we say that a point $p\in\sigma_{k}(X)$ is a non-trivial singular point if $p\notin\sigma_{k-1}(X)$ and $\sigma_{k}(X)$ is singular at $p$, while the points belonging to $\sigma_{k-1}(X)$ are called trivial singular points.
+There are some known results on the singular loci of $k$-th secant varieties $\sigma_{k}(X)$, mostly for very small $k$’s.
 The equality in (2) holds for determinantal varieties defined by minors of a generic matrix.
-It is completely described for the secant variety of the Segre product of projective spaces (i.e., $k=2$) in [27].
-It has been recently generalized to the case of $\sigma_{2}(X)$ where $X$ is a Segre-Veronese embedding by [2] and $X$ is a Grassmannian by [13].
-For the $3$-th secant variety of Grassmannian $\mathbb{G}(2,6)$, dimension and some other properties of the singular loci was studied in [4].
+It is completely described for the $2$-nd secant variety of the Segre product of projective spaces in [MOZ15].
+It has recently been generalized to the case of $\sigma_{2}(X)$, where $X$ is a Segre-Veronese embedding by [AMZ20] and $X$ is a Grassmannian by [GS24].
+For the $3$-rd secant variety of the Grassmannian $\mathbb{G}(2,6)$, the dimension and some other properties of the singularity have been studied in [AOP12].
 For the case of Veronese varieties, it is classical that $\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ holds both for the binary case (i.e., $n=1$) and for the case of quadratic forms (i.e., $d=2$) (see e.g.
-[19, chapter 1]).
-When $k=2$, it was proved that the equality above (i.e., having only trivial singular locus) is also true for any $d,n$ by [20].
-For $k=3$, by the second author the singular locus was completely determined in [16].
-In particular, it was shown that a non-trivial singularity can happen and it is only when $d=4$ and $n\geq 3$.
-In the present paper, we explore on the singular loci of any higher secant varieties of Veronese varieties and introduce another main origin for the singularity of $k$-secant variety other than the trivial singular locus.
-We call this the ‘subsecant locus’ and it shows an interesting trichotomy phenomenon among generic smoothness, non-trivial singular points and trivial singularity as in our main results.
-For any given point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, as considering $(1,d-1)$-symmetric flattening (see §3 for details), it is immediate to see that there exists an $m$-plane (i.e., $m$-dimensional linear subvariety) ${\mathbb{P}}^{m}$ of ${\mathbb{P}}^{n}$ with $11\leq m\leq k-$ 1 such that $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$.
-So, from now on, we call $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ an $m$-subsecant variety of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ (or simply a subsecant variety in case of no confusion) if $m<k-1$ and $m<n$.
-We also define
-(2) $\Sigma_{k,d}(m)=\Sigma_{k,d}(m;{{\mathbb{P}}^{n}}):=\bigcup_{\forall{\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}}\sigma_{k}(v_{d}({\mathbb{P}}^{m})),$
-the $m$-subsecant locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. It naturally forms an increasing sequence of loci in the $k$-secant as
-$\Sigma_{k,d}(1)\subset\Sigma_{k,d}(2)\subset\cdots\subset\Sigma_{k,d}(\min\{k-1,n\}-1)\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))=\Sigma_{k,d}(\min\{k-1,n\})~{}.$
-We call $\Sigma_{k,d}(\min\{k-1,n\}-1)$ (i.e., the union of all proper subsecant varieties) the maximum subsecant locus of the given $k$-th secant $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
-Any point outside the maximum subsecant locus and the previous secant variety (i.e., $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\big{(}\Sigma_{k,d}(\min\{k-1,n\}-1)\cup\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))\big{)}$) is called a point of full-secant locus.
-Note that in the case of $k=3$ of [16], the only exceptional case for $\operatorname{Sing}(\sigma_{3}(v_{d}({\mathbb{P}}^{n})))$ (i.e., $d=4$ and $n\geq 3$), the singular locus is given as the maximum subsecant locus $\Sigma_{3,d}(1)$, while the locus is just the trivial one, $\sigma_{2}(v_{d}({\mathbb{P}}^{n}))$, for all the other $d,n$ (see Remark 4 (a)).
-Most of previously known results on singular loci of secant varieties can be understood in this viewpoint (see Remark 37).
+[IK99, Chapter 1]).
+For $k=2$, it was proved in [Kan99] that the above equality holds also for any $d,n$.
+In these cases, the $k$-th secant variety has only the trivial singularity.
+For $k=3$, the singular locus was completely determined by the second author in [Han18]; in particular, the non-trivial singularity occur if and only if $d=4$ and $n\geq 3$.
+In the present paper, we explore the singular locus of any higher secant variety of the Veronese variety, and introduce a new main origin for the singularity other than the trivial singularity.
+We call this the ‘subsecant locus’.
+As in our main results, these loci show an interesting trichotomy phenomenon among generic smoothness, non-trivial singularity, and trivial singularity.
+For any given point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, there exists an $m$-plane (i.e., $m$-dimensional linear subvariety) ${\mathbb{P}}^{m}$ of ${\mathbb{P}}^{n}$ with $11\leq m\leq k-$ 1 such that $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$; it immediately follows for a general $p$, and even if $p$ is in the boundary of the closure in (1), it is also true by considering $(1,d-1)$-symmetric flattening (see §3 for details).
+So, from now on, we say that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is an $m$-subsecant variety of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ if $m<k-1$ and $m<n$, and simply call it a subsecant variety in case there is no confusion.
+We also define the $m$-subsecant locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$,
+[…]
+It naturally forms an increasing sequence of loci in the $k$-th secant variety as
+$\Sigma_{k,d}(1)\subset\Sigma_{k,d}(2)\subset\cdots\subset\Sigma_{k,d}(\min\{k-1,n\}-1)\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))=\Sigma_{k,d}(\min\{k-1,n\}).$
+In particular, $\Sigma_{k,d}(\min\{k-1,n\}-1)$ is the union of all proper subsecant varieties, which we call the maximum subsecant locus of the given $k$-th secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+Any point of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ outside the maximum subsecant locus and the previous secant variety $\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ is called a point of the full-secant locus.
+Note that for $k=3$ of [Han18], when $d=4$ and $n\geq 3$, the singular locus of $\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is given as the maximum subsecant locus $\Sigma_{3,d}(1)$, which is the only case where the singularity pattern of $3$-rd secant varieties becomes exceptional, while for all the other $d,n$ the singularity is just the trivial one, $\sigma_{2}(v_{d}({\mathbb{P}}^{n}))$ (see Remark 4 (a)).
+Most of the previously known results on singular loci of secant varieties can be understood in this viewpoint (see Remark 37).
 Thus, a basic question for our concern could be stated as follows: For given $k,d,m,n$,
-When does an $m$-subsecant locus $\Sigma_{k,d}(m)$ become singular in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$?
-In principle, it is in some sense (in spite of its computational complexity) straightforward to check the singularity, once a complete set of equations for a higher secant variety is attained.
-But, as mentioned above, not much has been known on the defining equations and it seems far from being understood completely at this moment, even for the Veronese case (see [23, 15] for the state of the art).
-Lacking in knowledge on the equations for higher secant varieties, it is very difficult to determine singular locus in general.
-In this paper, without further understanding on equations (!), we introduce a geometric way to pursue it for this kind of problems, which is based on a careful study on behavior of embedded tangent spaces moving along a locus in the Veronese variety .
-For $m=1$ case, we first present the following result for the $1$-subsecant locus $\Sigma_{k,d}(1)=\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which is a generalization of [16, theorem 2.1] (i.e., $k=3$ case) to any higher $k$-secant of Veronese varieties.
+When is $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ singular at points of an $m$-subsecant locus $\Sigma_{k,d}(m)$?
+In principle, it is somewhat straightforward (despite the computational complexity) to check the singularity, once a complete set of equations for a higher secant variety is attained.
+But, as mentioned above, not much is known about the defining equations and they seem quite far from being fully understood at this moment, even for the Veronese case (see [LO13, FH24] for the state of the art).
+Due to the lack of knowledge on the equations for the higher secant variety, it is very difficult to determine the singular locus in general.
+In this paper, without further understanding on the equations (!), we introduce a geometric way to pursue it for this kind of problems, which is based on a careful study on the behavior of embedded tangent spaces moving along a locus in the Veronese variety .
+For the case $m=1$, we first present the following result for the $1$-subsecant locus $\Sigma_{k,d}(1)=\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which is a generalization of [Han18, Theorem 2.1] (i.e., $k=3$ case) to any higher $k$-th secant varieties of Veronese varieties.
 Theorem 1.
@@
     If $k=\frac{d+2}{2}$ , then $\Sigma_{k,d}(1)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(1)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity).
-This case happens only if $d$ is even.
+This case occurs only if $d$ is even.
   3. (iii)
-    If $k\geq\frac{d+3}{2}$, then $\Sigma_{k,d}(1)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
+    If $k\geq\frac{d+3}{2}$, then $\Sigma_{k,d}(1)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity, unless $\sigma_{k}(v_{d}(\mathbb{P}^{n}))=\mathbb{P}^{N}$).
 For $(k,d,n)=(3,4,2)$, it holds that
@@
     $\sigma_{3}(v_{4}({\mathbb{P}}^{2}))$ is smooth at every point in $\Sigma_{3,4}(1)\setminus\sigma_{2}(v_{4}({\mathbb{P}}^{2}))$.
-Concerning singular points of arbitrary $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ originated from subsecant loci, we prove the following general theorems on the $m$-subsecant locus $\Sigma_{k,d}(m)$ with $m\geq 2$ and $k\geq 4$ as our main results.
+Concerning singular points of arbitrary $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ originated from subsecant loci, we prove the following general theorems on the $m$-subsecant locus $\Sigma_{k,d}(m)$ with $m\geq 2$ and $k\geq 4$ as the main results.
 Theorem 2.
@@
 Let $k\geq 4$ and let ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ be an $m$-plane with $2\leq m<\min\{k-1,n\}$.
-Assume that $(d,m)\notin\{(3,4),(4,2),(4,3),(4,4)\}$.
-For $(k,d,n)\neq(4,3,3)$, setting $\mu$ to be ($\frac{{m+d\choose m}}{m+1}$ if $(d,m)\notin\{(6,2),(3,5)\}$ and to be ($\frac{{m+d\choose m}}{m+1}-1$ if $(d,m)\in\{(6,2),(3,5)\}$, we have:
+Assume that $(d,m)\notin\mathcal{E}=\{(3,3),(3,4),(3,5),(4,2),(4,3),(4,4),(5,2),(6,2)\}$.
+For $(k,d,n)\neq(4,3,3)$, setting $\mu=\Bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}\Bigl{\rceil}$, we have:
   1. (i)
-    If $k\leq\mu$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
+    If $k<\mu$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
   2. (ii)
-    If $\mu<k<\mu+1$, then $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity).
-In particular, this singularity appears if and only if $\mu\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$.
+    If $k=\mu$, then $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity).
   3. (iii)
-    If $k\geq\mu+1$, then $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
+    If $k>\mu$, then $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity, unless $\sigma_{k}(v_{d}(\mathbb{P}^{n}))=\mathbb{P}^{N}$).
 For $(k,d,n)=(4,3,3)$, it holds that
@@
     $\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ is smooth at every point in $\Sigma_{4,3}(2)\setminus\sigma_{3}(v_{3}({\mathbb{P}}^{3}))$.
-For the cases of $(d,m)=(3,4),(4,2),(4,3)$ or $(4,4)$, only in which the dimension of a $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ can be defective, we have the following statement.
 Theorem 3.
-In the same situation as Theorem 2, for $(d,m)\in\{(3,4),(4,2),(4,3),(4,4)\}$ and $\mu=\frac{{m+d\choose m}}{m+1}$, the following holds:
+In the same situation as Theorem 2, for
+$(d,m)\in\mathcal{E}=\{(3,3),(3,4),(3,5),(4,2),(4,3),(4,4),(5,2),(6,2)\},$
+if $k$ is in one of the ranges named (i), (ii), (iii) in Table 1 below, then the following property corresponding to the name of the range holds.
   1. (i)
-    If $k\leq\lfloor\mu\rfloor-1$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
+    $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
   2. (ii)
-    If $\lfloor\mu\rfloor\leq k\leq\lceil\mu\rceil+1$, then $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
+    $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
   3. (iii)
-    If $k\geq\lceil\mu\rceil+2$, then $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
-We summarize Theorem 3 in Table 1 with exact ranges of $k$.
-$(d,m)$ $\Sigma_{k,d}(m)$ : Generically smooth (i) Non-trivial singular locus (ii) Trivial singularity (iii)
-$(3,4)$ $k\leq 6$ $k=7,8$ $9\leq k$
-$(4,2)$ $k\leq 4$ $k=5,6$ $7\leq k$
-$(4,3)$ $k\leq 7$ $k=8,9,10$ $11\leq k$
-$(4,4)$ $k\leq 13$ $k=14,15$ $16\leq k$
-Table 1. (Non-)singularity of $\Sigma_{k,d}(m)$ in case of defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s in Theorem 3
+    $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
+$(d,m)$ $\frac{\binom{m+d}{m}}{m+1}$ (i) (ii) (iii)
+$(3,3)$ $5$ $k\leq 5$ None $6\leq k$
+$(3,4)$ $7$ $k\leq 6$ $k=7,8$ $9\leq k$
+$(3,5)$ ${28}/{3}$ $k\leq 8$ $k=9,10$ $11\leq k$
+$(4,2)$ $5$ $k\leq 4$ $k=5,6$ $7\leq k$
+$(4,3)$ ${35}/{4}$ $k\leq 7$ $k=8,9,10$ $11\leq k$
+$(4,4)$ $14$ $k\leq 13$ $k=14,15$ $16\leq k$
+$(5,2)$ $7$ $k\leq 7$ None $8\leq k$
+$(6,2)$ ${28}/{3}$ $k\leq 8$ $k=9,10$ $11\leq k$
+Table 1. (Non-)singularity of $\Sigma_{k,d}(m)$ in Theorem 3
+To understand the reason for considering the conditions that $(d,m)$ is in $\mathcal{E}$ or not, we discuss the secant defectivity and the generic identifiability of an $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+Set ${\mathbb{P}}^{\beta}=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle\subset{\mathbb{P}}^{N}$ with $\beta=\binom{m+d}{m}-1$.
+where $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset{\mathbb{P}}^{\beta}$ is of dimension $\leq km+k-1$.
+From [AH95], the codimension of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ in ${\mathbb{P}}^{\beta}$ is greater than $\max\{\beta-(km+k-1),0\}$ (i.e., $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ does not fill ${\mathbb{P}}^{\beta}$ and is secant defective) if and only if $d=2$ and $2\leq k\leq m$, or $(k,d,m)=(7,3,4),(5,4,2),(9,4,3),(14,4,4)$; indeed, in the latter case, the four $k$-th secant varieties $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s are hypersurfaces of ${\mathbb{P}}^{\beta}$.
+Except these defective cases, we have $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ if $km+k-1\geq\beta$, or equivalently $k\geq{\binom{m+d}{m}}/({m+1})$.
+We say that a point $a\in{\mathbb{P}}^{\beta}$ is $k$-identifiable if there is a unique $k$-tuple of points $x_{1},\dots,x_{k}\in v_{d}({\mathbb{P}}^{m})$ such that $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$.
+We also say that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is generically identifiable if a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is $k$-identifiable.
+From [GM17, Theorem 1], a general point $a\in{\mathbb{P}}^{\beta}$ is $k$-identifiable (or $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is generically identifiable in the case of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$) if and only if $m=1$ and $k=(d+1)/2$, or $(k,d,m)=(5,3,3),(7,5,2)$.
+From [COV17, Theorem 1.1], in the case when $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}$ is not secant defective, $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is not generically identifiable if and only if $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$.
 Remark 4.
-We’d like to make some remarks on the theorems above.
+We make some remarks on the theorems above.
   1. (a)
-    In the case of $k=3$ and $n\geq 3$, the condition of (ii) in Theorem 1 holds if and only if $d=4$, and then $\operatorname{Sing}(\sigma_{3}(v_{4}({\mathbb{P}}^{n}))$ is equal to $\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle=\Sigma_{3,4}(1)$ since $\sigma_{3}(v_{4}({\mathbb{P}}^{1}))=\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle$ and $\sigma_{2}(v_{4}({\mathbb{P}}^{n}))\subset\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle$ (see also Corollary 32).
-This gives a geometric description of the only exceptional case for the singular loci of $3$rd secant variety in [16].
+    In the case of $k=3$ and $n\geq 3$, the condition of (ii) in Theorem 1 holds if and only if $d=4$, and then $\operatorname{Sing}(\sigma_{3}(v_{4}({\mathbb{P}}^{n}))$ is equal to $\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle=\Sigma_{3,4}(1)$ since $\sigma_{3}(v_{4}({\mathbb{P}}^{1}))=\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle$ and $\sigma_{2}(v_{4}({\mathbb{P}}^{n}))\subset\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle$ (see also Corollary 32).
+This gives a geometric description of the only exceptional case for the singular loci of the $3$-rd secant varieties in [Han18].
   2. (b)
@@
 The ‘general point’ condition of (i) in Theorems 2 and 3 cannot be deleted (see Example 27).
-Except this aspect, (i)-(iii) of Theorem 1 can be also derived as applying Theorem 2 directly to the case $m=1$ (recall that curve case has no defectivity in its higher secants).
+We also note that, the ranges of $k$ of (i) and (ii) are slightly different between Theorems 1 and 2.
+If $m=1$ and $d$ is even, then the case $k=\lceil{\binom{m+d}{m}}/({m+1})\rceil=(d+2)/2$ is (ii) of Theorem 1, which is similar to that of Theorem 2; for this $k$, $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ is not generically identifiable.
+However, if $d$ is odd, then the case $k={\binom{m+d}{m}}/({m+1})=(d+1)/2$ belongs to (i) of Theorem 1.
   3. (c)
-    (i) in both Theorem 2 and Theorem 3 mostly corresponds to the identifiable case of a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$, which means that there is a unique $k$-tuple of points $x_{1},\dots,x_{k}\in v_{d}({\mathbb{P}}^{m})$ such that $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$.
-In the viewpoint of secant fiber in the incidence (3) in §2, it means that the fiber $p^{-1}(a)$ under $p$ consists of only one element up to permuting $x_{i}$’s.
-On the other hand, (ii) of Theorems 2 and 3 corresponds to the case of generic non-identifiability of the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ and to the situation of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq\langle v_{d}({\mathbb{P}}^{m})\rangle$.
-This non-identifiability can happen if $\dim(p^{-1}(a))>0$, or if #$\#p^{-1}(a)\geq 2$ (modulo permutation) in case of $\dim p^{-1}(a)=0$; the former one corresponds to $(d,m)\notin\{(6,2),(3,5),(4,3)\}$ and the latter $(d,m)\in\{(6,2),(3,5),(4,3)\}$.
+    (i) in Theorems 1, 2, and 3 corresponds to the $k$-identifiable case of a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$.
+From the viewpoint of the secant fiber in the incidence (4) in §2, it means that the fiber $p^{-1}(a)$ under the projection $p$ consists of only one element up to permuting $x_{i}$’s.
+On the other hand, (ii) of Theorems 1, 2, and 3 corresponds to the case of generic non-identifiability of the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ with the situation of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle$, except $(k,d,m,n)=(3,4,1,2),(4,3,2,3)$.
+This non-identifiability of a general $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ occurs if $\dim p^{-1}(a)>0$, or if $\dim p^{-1}(a)=0$ and #$\#p^{-1}(a)\geq 2$ (modulo permutation).
+For $m=1$, only the former occurs in the case $k=(d+2)/2$.
+For $m\geq 2$, the former corresponds to the case where $k$ is the ceiling of ${\binom{m+d}{m}}/({m+1})\notin{\mathbb{N}}$ with $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$, to the case where $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}$ is secant defective, or to the case where $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}$ is secant defective (i.e., $(k,d,m)=(8,3,4),(6,4,2),(10,4,3),(15,4,4)$), and the latter corresponds to the case where $k$ is the number ${\binom{m+d}{m}}/({m+1})\in{\mathbb{N}}$ with $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ except $(k,d,m)=(5,3,3),(7,5,2)$, or to the case $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$.
   4. (d)
-    Theorem 3 deals with the case in which the given subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ can possibly be defective due to the famous theorem of Alexander-Hirschowitz [1], whereas Theorem 2 concerns the case of always non-defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s.
-  5. (e)
-    Two cases $(k,d,m,n)=(3,4,1,2),(4,3,2,3)$ (i.e., (iv) of Theorems 1 and  2), a posteriori, turn out to be the only two exceptional cases which do not follow this trichotomy pattern; in other words, though $k$ belongs to the range of (ii), $\Sigma_{k,d}(m)$ does not provide non-trivial singularity (see also Remark 24).
-In [15], we in fact show that if $(k,d,n)=(3,4,2),(4,3,3)$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is a del Pezzo $k$-secant variety, that is, a $k$-secant variety of next-to-minimal degree.
-In this sense, these two cases also belong to a special class with respect to degrees of higher secant varieties.
-(For basic definitions and results on such varieties, see [11], [10].)
-As an application of our main results for $\Sigma_{4,d}(2)$ and $\Sigma_{4,d}(1)$, we obtain the following result on the singular locus of the fourth secant variety of any Veronese embeddings.
+    $(k,d,m,n)=(3,4,1,2),(4,3,2,3)$ (i.e., (iv) of Theorems 1 and  2), a posteriori, turn out to be the only two exceptional cases which do not follow this trichotomy pattern; in other words, though $k$ belongs to the range of (ii) and the generic non-identifiability of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ holds, $\Sigma_{k,d}(m)$ does not provide non-trivial singular points (see also Remark 24).
+Indeed, in [FH24] we show that if $(k,d,n)=(3,4,2),(4,3,3)$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is a del Pezzo $k$-th secant variety, that is, a $k$-th secant variety of next-to-minimal degree.
+In this sense, these two cases also belong to a special class with respect to the degrees of higher secant varieties.
+(For basic definitions and results on such varieties, see [CR], [CK].)
+As an application of our main results for $\Sigma_{4,d}(2)$ and $\Sigma_{4,d}(1)$, we obtain the following result on the singularity of the fourth secant variety of any Veronese variety.
 Theorem 5 (Singular locus for $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$).
@@
   3. (iii)
-    For $d\geq 7$, $\Sigma_{4,d}(1)\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\subset\operatorname{Sm}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$.
+    For $d\geq 7$, all points in $\Sigma_{4,d}(1)\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\subset\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ are smooth.
 When $d=6$, $\Sigma_{4,d}(1)\subset\operatorname{Sing}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{4,d}(1)\not\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity).
 When $d\leq 5$, $\Sigma_{4,d}(1)\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
+For a projective variety $X\subset{\mathbb{P}}^{N}$, we denote by $\operatorname{Vertex}(X)$ the set of vertices of $X$. Then $X$ is a cone if and only if $\operatorname{Vertex}(X)\neq\emptyset$.
 Example 6 (Cases with a nice description).
-The smallest case for singular locus of $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ beyond the classical results for $d=2$ or $n=1$ is $(d,n)=(3,2)$, but there is nothing to check in this case, because $\sigma_{4}(v_{3}({\mathbb{P}}^{2}))$ fills up the ambient space ${\mathbb{P}}^{9}$ (i.e., $\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{2})))=\emptyset$).
-In next case $(d,n)=(3,3)$, by Theorem 5 (i) and (ii), we have
-$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{3})))=\sigma_{3}(v_{3}({\mathbb{P}}^{3}))~{}.$
-For the case $(d,n)=(3,4)$, if $V$ denotes a $5$-dimensional ${\mathbb{C}}$-vector space with ${\mathbb{P}}^{4}={\mathbb{P}}V$, Theorem 5 tells us that the singular locus of the $4$-th secant of $v_{3}({\mathbb{P}}V)$ in ${\mathbb{P}}^{34}$ is precisely the locus of cubic hypersurfaces in 5 variables which are cones with the vertex dimension $\geq 1$ as follows:
-$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}V)))=\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\Sigma_{4,3}(2;{{\mathbb{P}}V})=\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\sigma_{4}(v_{3}({\mathbb{P}}^{2}))\big{\}}=\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\langle\,v_{3}({\mathbb{P}}^{2})\,\rangle$
-=$=\big{\{}f\in{\mathbb{P}}S^{3}V~{}|~{}\textrm{$f$ defines a cubichypersurface in ${\mathbb{P}}V$ which is a cone with $\dim\operatorname{Vertex}(f)\geq 1$}\big{\}}~{},$
+The smallest case for the singular locus of $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ beyond the classical results for $d=2$ or $n=1$ is $(d,n)=(3,2)$, but in this case there is nothing to check, because $\sigma_{4}(v_{3}({\mathbb{P}}^{2}))$ fills up the ambient space ${\mathbb{P}}^{9}$ and then $\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{2})))=\emptyset$.
+In the case $(d,n)=(3,3)$, by Theorem 5 (i) and (ii), we have
+$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{3})))=\sigma_{3}(v_{3}({\mathbb{P}}^{3})).$
+For the case $(d,n)=(3,4)$, if $V$ denotes a $5$-dimensional ${\mathbb{C}}$-vector space with ${\mathbb{P}}^{4}={\mathbb{P}}V$, Theorem 5 tells us that the singular locus of the $4$-th secant variety of $v_{3}({\mathbb{P}}V)$ in ${\mathbb{P}}^{34}$ is precisely the locus of cubic hypersurfaces in $5$ variables which are cones with the vertex dimension $\geq 1$ as follows:
+$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}V)))=\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\Sigma_{4,3}(2;{{\mathbb{P}}V})\\ =\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\sigma_{4}(v_{3}({\mathbb{P}}^{2}))\big{\}}=\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\langle\,v_{3}({\mathbb{P}}^{2})\,\rangle\\ =\big{\{}f\in{\mathbb{P}}S^{3}V~{}|~{}\textrm{the cubic hypersurface $X\subset{\mathbb{P}}V$ defined by $f$ is a cone with $\dim\operatorname{Vertex}(X)\geq 1$}\big{\}},$
 which is just the maximum subsecant locus $\Sigma_{4,3}(2;{{\mathbb{P}}V})$, an irreducible $15$-dimensional locus in the $19$-dimensional variety $\sigma_{4}(v_{3}({\mathbb{P}}V))$.
@@
 $\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{n})))=\Sigma_{4,3}(2;{{\mathbb{P}}^{n}})$ for any $n\geq 4$.
-This kind of a simple description of the singular locus can be attained in few more cases (see Corollary 32 for details).
+Such a simple description of the singular locus can be attained in a few more cases (see Corollary 32 for details).
 The paper is structured as follows.
-In §2, as preparation, we recall some preliminaries on $k$-th secant varieties and corresponding incidences, and reveal some geometric properties of subsecant varieties in higher secants of Veronese varieties, which are crucial for the proof of main theorems.
+In §2, as preparation, we first recall some preliminaries on $k$-th secant varieties and corresponding incidences.
+Then, using projective techniques, such as Terracini’s lemma, the trisecant lemma, descriptions of embedding tangent spaces, and tangential projections, we reveal several geometric properties of $m$-subsecant varieties in higher secant varieties of Veronese varieties, which are crucial for the proof of main theorems.
 In §3, as an illustration of the whole picture and our main ideas, we treat the case $m=1$ and prove Theorem 1.
 In §4 we deal with the general case (i.e., $m\geq 2$) and prove Theorem 2 and Theorem 3 as generalizing the ideas used in the previous section.
-We’d like to remark that this can be done because the dimension theorem [1] and general identifiability question [9, 12] was settled for the case of Veronese varieties.
-In §5 we focus on the singular locus of $4$-th secant variety and prove Theorem 5 by dividing the case into two parts; ‘full-secant locus points (i.e., $m=3$)’ treated in Theorem 29 via Young flattening and conormal space computation and ‘the subsecant locus’ by Corollary 30.
-Finally, we make some generalization and remarks on the materials for further development in §6.
+We would like to remark that this can be done because the dimension theorem [AH95] and the generic identifiability question [COV17, GM17] were settled for the case of Veronese varieties.
+In §5 we focus on the singular locus of the $4$-th secant variety and prove Theorem 5 by dividing the case into two parts; ‘full-secant locus points (i.e., $m=3$)’ treated in Theorem 29 via Young flattening and conormal space computation and ‘the subsecant locus’ by Corollary 30.
+Finally, we make some generalizations and remarks on the material for further developments in §6.
 Acknowledgements
@@
 The first author also wishes to express his gratitude to Hajime Kaji and Hiromichi Takagi for useful discussions.
+Finally, we would like to thank the anonymous referees very much for their careful reading, pointing out some error in the first version, and many suggestions which help us to improve the final exposition of the paper.
 2. Some geometric properties of subsecant varieties
@@
 For the $d$-uple Veronese embedding $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ with $N={n+d\choose d}-1$, we regard the incidence variety $I=I_{(n)}\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$ to be the Zariski closure of
-(3) $I^{0}=I_{(n)}^{0}:=\left\{(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
+(4) $I^{0}=I_{(n)}^{0}=\left\{(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
 where we write $x_{i}=v_{d}(x_{i}^{\prime})$ for $x_{i}^{\prime}\in{\mathbb{P}}^{n}$.
-[…]
-Taking the first projection $p:I\rightarrow{\mathbb{P}}^{N}$, we have $p(I)=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ (see also [28, definition 1.1.3], [29, chapter I, §1]).
+Taking the first projection $p:I\rightarrow{\mathbb{P}}^{N}$, we have $p(I)=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ (see also [Rus16, Definition 1.1.3], [Zak93, Chapter I, §1, Chapter V]).
 For any $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, $p^{-1}(a)$ is often called the secant fiber of $a$.
 Note that $I$ is invariant under permuting factors on $({\mathbb{P}}^{n})^{k}$ from the definition so that both $p$ and $q_{i}$ maps factor through ${\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$.
-We also have $\dim I={nk+k-1}$ by considering general fibers of $q:I\rightarrow({\mathbb{P}}^{n})^{k}$.
-For each $1\leq i\leq k$, let
+We also have $\dim I={nk+k-1}$ by considering general fibers of $q:I\rightarrow({\mathbb{P}}^{n})^{k}$. For each $1\leq i\leq k$, let
 $q_{i}:I\rightarrow({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{n}$
-be the composition of $q$ and the projection to the $i$-th factor $({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{n}$.
+be the composition of $q$ and the projection to the $i$-th factor $({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{n}$. Then the following commutative diagram is obtained:
+${I}$${{\mathbb{P}}^{N}}$${({\mathbb{P}}^{n})^{k}}$${{\mathbb{P}}^{n}.}$$\scriptstyle{p}$$\scriptstyle{q}$$\scriptstyle{q_{i}}$
 Remark 7.
-We have some remarks on the incidence $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$.
+We have some remarks on the incidence variety $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$.
   1. (a)
     $I^{0}$ can be viewed as a ${\mathbb{P}}^{k-1}$-bundle over a non-empty open subset $U$ of $({\mathbb{P}}^{n})^{k}$, consisting of $k$-tuples of points with the expected spanning dimension, so that $I$ is irreducible.
-Further, $I^{0}=q^{-1}(U)$, since both are irreducible closed subsets in ${\mathbb{P}}^{N}\times U$ having the same dimension.
+Further, for $q^{-1}(U)=I\cap({\mathbb{P}}^{n}\times U)$, we have $I^{0}=q^{-1}(U)$ since both are irreducible closed subsets in ${\mathbb{P}}^{N}\times U$ having the same dimension.
 So, for any $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in I\setminus I^{0}$ and for $x_{i}=v_{d}(x_{i}^{\prime})$, it holds that $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ (in other words, there is no $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in I\setminus I^{0}$ such that $\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1$ but $a\notin\langle\,x_{1},\dots,x_{k}\,\rangle$).
 Finally, note that the Euclidean closure of $I^{0}$ also coincides with $I$ in this case (see e.g.
-[22, theorem 3.1.6.1]).
+[Land17, Theorem 3.1.6.1]).
   2. (b)
-    In the case of $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$, it holds $p(I\setminus I^{0})\neq\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. In addition, if $k\leq\binom{n+d-1}{n}$, then as setting
+    In the case of $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$, it holds $p(I\setminus I^{0})\neq\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, and then $p^{-1}(a)\subset I^{0}$ for general $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+In addition, if $k\leq\binom{n+d-1}{n}$, then as setting
     $D=\left\{(x_{1}^{\prime},\dots,x_{k}^{\prime})\in({\mathbb{P}}^{n})^{k}\mid\dim(\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle)<k-1\right\}\!,$
@@
     (Alternative incidences for the $k$-th secant variety) In the literature, instead of $(X)^{k}$, other spaces such as the symmetric product $\operatorname{Sym}^{k}(X)$ (e.g.
-in [9]), the Hilbert scheme of degree $k$ 00-dimensional subschemes $\operatorname{Hilb}_{k}(X)$ (e.g.
-in [7]), and the Grassmannian $\mathbb{G}(k-1,N)$ (e.g.
-in [26]) also have been used in the incidence to consider the $k$-th secant variety of $X\subset{\mathbb{P}}^{N}$.
+in [COV17]), the Hilbert scheme of degree $k$ 00-dimensional subschemes $\operatorname{Hilb}_{k}(X)$ (e.g.
+in [BB14]), and the Grassmannian $\mathbb{G}(k-1,N)$ (e.g.
+in [Mel06]) have also been used in the incidence to consider the $k$-th secant variety of $X\subset{\mathbb{P}}^{N}$.
 Remark 8.
-Let us fix an $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ and consider the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$, $v_{d}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta}$ with $\beta={m+d\choose m}-1$.
+Let us fix an $m$-plane $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ and consider the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$, $v_{d}:L={\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta}$ with $\beta={m+d\choose m}-1$.
 We often use the following notation.
   1. (a)
-    Changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}:=n-m$, we may assume that a given ${\mathbb{P}}^{m}$ in ${\mathbb{P}}^{n}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$.
+    Let $\hat{L}\subset{\mathbb{P}}^{n}$ be any $(n-m-1)$-plane not intersecting $L$.
+Changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}=n-m$, we may assume that $L\subset{\mathbb{P}}^{n}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$ and $\hat{L}\subset{\mathbb{P}}^{n}$ is the zero set of $t_{0},\dots,t_{m}$.
 For any $x_{i}^{\prime}\in{\mathbb{P}}^{n}$, say
@@
 By abuse of notations, we denote by $y_{i}^{\prime}$ both corresponding points in ${\mathbb{P}}^{m}$ and in ${\mathbb{P}}^{n}$.
-Further, as considering linear projections $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ (eliminating the $u$-variables) and $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ (eliminating all the monomials of degree $d$ which involve $u$-variables), we have a natural commuting diagram as $v_{d}\circ\pi_{1}=\pi_{2}\circ v_{d}$
+Further, as considering linear projections $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ from the center $\hat{L}$ (eliminating the $u$-variables), and $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ (eliminating all the monomials of degree $d$ which involve $u$-variables), we have a natural commuting diagram as $v_{d}\circ\pi_{1}=\pi_{2}\circ v_{d}$
 […]
@@
 in ${\mathbb{C}}[t_{1},\dots,t_{m}]$) of degree $\leq e$ for an integer $e$.
-Let us study behaviors of some points in the boundary of $I$, which belong to $I$ but do not belong to $I^{0}$, as follows.
+Let us study the behavior of some points in the boundary of $I$, which belong to $I$ but do not belong to $I^{0}$, as follows.
 Lemma 9.
 Let $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ be any $m$-plane with $m<n$, and consider ${\mathbb{P}}^{\beta}=\langle\,v_{d}(L)\,\rangle\subset{\mathbb{P}}^{N}$.
-For $v_{d}:L={\mathbb{P}}^{m}\hookrightarrow{\mathbb{P}}^{\beta}$, let us take the incidence variety $I_{(m)}\subset{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k}$ of the same kind as $I=I_{(n)}$ in (3).
+For $v_{d}:L={\mathbb{P}}^{m}\hookrightarrow{\mathbb{P}}^{\beta}$, let us take $I_{(m)}\subset{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k}$ to be the closure of the set of points $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k}$ such that $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$ and $\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1$ (i.e., the incidence variety of the same kind as $I=I_{(n)}$ in (4)).
 In this setting, for any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ with $a\in{\mathbb{P}}^{\beta}$, one of the following conditions holds:
@@
 Proof.
-Let ${\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ be a linear projection whose center does not intersect with $L$.
-It induces a natural linear projection ${\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$, and then we have
+Let $\left\{W_{j}\right\}$ be the irreducible components of $I\setminus I^{0}$. For each $j$, we take
+(6) $(a_{j},x_{j,1}^{\prime},\dots,x_{j,k}^{\prime})\in W_{j}\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}.$
+Let $\hat{L}\subset{\mathbb{P}}^{n}$ be a general $(n-m-1)$-plane such that all $x_{j,i}^{\prime}\notin\hat{L}$ and $L\cap\hat{L}=\emptyset$.
+For the linear projection $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ from the center $\hat{L}$, we have a natural linear projection $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ as in Remark 8 with the diagram (5), and then define the projections
 $\rho_{1}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\dashrightarrow{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k},\quad\rho_{2}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\dashrightarrow{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k},$
 where $\overline{\rho_{2}(\rho_{1}(I))}=I_{(m)}$. Taking two Segre embeddings
-${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}},\quad{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\hookrightarrow{\mathbb{P}}^{l_{2}},$
-we may regard $\rho_{1}$ as the restriction of a linear projection ${\mathbb{P}}^{l_{1}}\dashrightarrow{\mathbb{P}}^{l_{2}}$ whose center is a certain linear subvariety $R\subset{\mathbb{P}}^{l_{1}}$.
+$Seg_{1}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}},\quad Seg_{2}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\hookrightarrow{\mathbb{P}}^{l_{2}},$
+we may regard $\rho_{1}$ as the restriction of a linear projection $\pi_{R}:{\mathbb{P}}^{l_{1}}\dashrightarrow{\mathbb{P}}^{l_{2}}$ whose center $R$ is a certain linear subvariety of ${\mathbb{P}}^{l_{1}}$.
+Let ~$\tilde{{\mathbb{P}}}^{l_{1}}\subset{\mathbb{P}}^{l_{1}}\times{\mathbb{P}}^{l_{2}}$ be the graph of $\pi_{R}$, which coincides with the blowing-up of ${\mathbb{P}}^{l_{1}}$ with respect to $R$.
+Let $r_{1}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{1}}$ and $r_{2}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{2}}$ be projections.
+Then we have the following commutative diagram:
 […]
 Under the embedding $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}}$, taking $R\cap I$ in ${\mathbb{P}}^{l_{1}}$, we have $\operatorname{codim}(R\cap I,I)\geq 2$, as follows.
-Let $R^{\prime}$ be the center of the projection ${\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$, where $R^{\prime}\cap L=\emptyset$ and $\operatorname{codim}(R^{\prime},{\mathbb{P}}^{n})=m+1$.
-For $R^{\prime}(i)=\left\{(x_{1}^{\prime},\dots,x_{k}^{\prime})\in({\mathbb{P}}^{n})^{k}\mid x_{i}^{\prime}\in R^{\prime}\right\}$, we have
-$R\cap I=q^{-1}\bigl{(}\bigcup_{i=1}^{k}R^{\prime}(i)\bigr{)},$
-where $q:I\rightarrow({\mathbb{P}}^{n})^{k}$.
-In particular, since dimension of fibers of $q|_{I^{0}}$ is constant, it follows that $\operatorname{codim}(R\cap I^{0},I)=m+1$.
-Next, let $\left\{W_{j}\right\}$ be the irreducible components of $I\setminus I^{0}$, and let $(a_{j},x_{j,1}^{\prime},\dots,x_{j,k}^{\prime})\in W_{j}$.
-A general $(n-m-1)$-plane ${\mathbb{P}}^{n-m-1}\subset{\mathbb{P}}^{n}$ satisfies $x_{j,i}^{\prime}\notin{\mathbb{P}}^{n-m-1}$.
-Since $\left\{W_{j}\right\}$ does not depend on the choice of $R^{\prime}={\mathbb{P}}^{n-m-1}$, we may assume $x_{j,i}^{\prime}\notin R^{\prime}$.
-Then $W_{j}\not\subset R$ for any $j$.
-Taking an irreducible component $Q$ of $R\cap I$ such that $Q\subset I\setminus I^{0}$, we have some $j$ such that $Q\subset R\cap W_{j}\subsetneq W_{j}\subsetneq I$.
-Hence $\operatorname{codim}(Q,I)\geq 2$.
-As a result, $\operatorname{codim}(R\cap I,I)\geq 2$.
-Now, let $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ satisfy $a\in{\mathbb{P}}^{\beta}$, where $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ for $x_{i}=v_{d}(x_{i}^{\prime})$ (note that we do not know $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$).
+For the center $\hat{L}$ of $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$, we set
+$\hat{L}(i)=\{\,(x_{1}^{\prime},\dots,x_{k}^{\prime})\in({\mathbb{P}}^{n})^{k}\mid x_{i}^{\prime}\in\hat{L}\,\}.$
+We consider the projection $\bar{q}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow({\mathbb{P}}^{n})^{k}$.
+Then a point $z\in{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$ belongs to the indeterminacy locus of $\rho_{1}$ if and only if $\bar{q}(z)\in\hat{L}(i)$ for some $i$ with $1\leq i\leq k$.
+Since $\pi_{R}|_{I}(=\pi_{R}\circ Seg_{1}|_{I})$ coincides with $Seg_{2}\circ\rho_{1}|_{I}$, and since $R\cap I$ is the indeterminacy locus of $\pi_{R}|_{I}$, we have
+$R\cap I=q^{-1}\bigl{(}\bigcup_{i=1}^{k}\hat{L}(i)\bigr{)},$
+where $q$ is equal to $\bar{q}|_{I}:I\rightarrow({\mathbb{P}}^{n})^{k}$.
+In particular, since the dimension of the fibers of $q|_{I^{0}}$ is constant, and since $\operatorname{codim}(\hat{L}(i),({\mathbb{P}}^{n})^{k})=m+1$, it follows that $\operatorname{codim}(R\cap I^{0},I)=m+1$.
+Now, let $\left\{Q_{s}\right\}$ be the irreducible components of $R\cap I$.
+Then $\operatorname{codim}(Q_{s},I)=m+1$ in the case $Q_{s}\cap I^{0}\neq\emptyset$, i.e., $Q_{s}\not\subset I\setminus I^{0}$.
+In order to consider the remaining case $Q_{s}\subset I\setminus I^{0}$, we use the irreducible components $\left\{W_{j}\right\}$ of $I\setminus I^{0}$.
+For $(a_{j},x_{j,1}^{\prime},\dots,x_{j,k}^{\prime})\in W_{j}$ given in (6), since $x_{j,i}^{\prime}\notin\hat{L}$, it follows $W_{j}\not\subset R$ for any $j$.
+If an irreducible component $Q_{s}$ of $R\cap I$ satisfies $Q_{s}\subset I\setminus I^{0}$, then there is some $j$ such that $Q_{s}\subset R\cap W_{j}\subsetneq W_{j}\subsetneq I$, and then $\operatorname{codim}(Q_{s},I)\geq 2$.
+As a result, in any case, each irreducible component of $R\cap I$ is of codimension $m+1$ or at least $\geq 2$, which implies $\operatorname{codim}(R\cap I,I)\geq 2$.
+Now, let $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ satisfy $a\in{\mathbb{P}}^{\beta}$, where $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ for $x_{i}=v_{d}(x_{i}^{\prime})$ as in Remark 7 (a) (note that we do not know $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$).
 We regard $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$ as a point of ${\mathbb{P}}^{l_{1}}$ under the embedding $I\subset{\mathbb{P}}^{l_{1}}$.
 If $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\notin R$, then $\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$ is determined in $\overline{\rho_{1}(I)}\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}$, where $y_{i}^{\prime}\in{\mathbb{P}}^{m}$ is the image of $x_{i}^{\prime}$ under ${\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{m}$.
-Since $a\in{\mathbb{P}}^{\beta}$, $\rho_{2}(\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime}))=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$ is determined and is contained in $I_{(m)}$.
+Since $a\in{\mathbb{P}}^{\beta}$,
+$\rho_{2}(\rho_{1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime}))=(a,y_{1}^{\prime},\dots,y_{1}^{\prime})$
+is determined and is contained in $I_{(m)}$.
 Then, $\dim\langle\,y_{1},\dots,y_{k}\,\rangle<k-1$ for $y_{i}=v_{d}(y_{i}^{\prime})\in{\mathbb{P}}^{\beta}$, which means $(a,y_{1}^{\prime},\dots,y_{1}^{\prime})\in I_{(m)}\setminus I_{(m)}^{0}$.
 Assume $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in R$.
-Let ~$\tilde{{\mathbb{P}}}^{l_{1}}\subset{\mathbb{P}}^{l_{1}}\times{\mathbb{P}}^{l_{2}}$ be the graph of ${\mathbb{P}}^{l_{1}}\dashrightarrow{\mathbb{P}}^{l_{2}}$.
-In other words, $\widetilde{{{\mathbb{P}}}^{l_{1}}}=Bl_{R}({\mathbb{P}}^{l_{1}})$, the blowing-up of ${\mathbb{P}}^{l_{1}}$ with respect to $R$.
-Let $r_{1}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{1}}$ and $r_{2}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{2}}$ be projections.
-Note that, for the strict transformation $S\subset\tilde{{\mathbb{P}}}^{l_{1}}$ of ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\subset{\mathbb{P}}^{l_{1}}$, two composite morphisms $S\xrightarrow{r_{1}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}$ and $S\xrightarrow{r_{2}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$ coincide since it holds on an open subset of $S$.
+We consider the blowing-up ~$\tilde{{\mathbb{P}}}^{l_{1}}$ of ${\mathbb{P}}^{l_{1}}$ with respect to $R$, and the projections $r_{1},r_{2}$ in the diagram (7).
+Note that, for the strict transformation $S\subset\tilde{{\mathbb{P}}}^{l_{1}}$ of ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\subset{\mathbb{P}}^{l_{1}}$, two composite morphisms
+$S\xrightarrow{r_{1}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}\,\text{ and }\,S\xrightarrow{r_{2}|_{S}}{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$
+coincide since it holds on an open subset of $S$.
 Let $E\subset\tilde{{\mathbb{P}}}^{l_{1}}$ be the exceptional divisor, and let $\widetilde{I}\subset\tilde{{\mathbb{P}}}^{l_{1}}$ be the strict transformation of $I\subset{\mathbb{P}}^{l_{1}}$.
 Then $r_{2}(\widetilde{I})=\overline{\rho_{1}(I)}$.
-Let $E_{1}=r_{1}^{-1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\cap(E\cap\widetilde{I})$, the fiber of $E\cap\widetilde{I}\rightarrow R\cap I$ at $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$.
+Let
+$E_{1}=r_{1}^{-1}(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\cap(E\cap\widetilde{I}),$
+the fiber of $E\cap\widetilde{I}\rightarrow R\cap I$ at $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$.
 It follows from $\operatorname{codim}(R\cap I,I)\geq 2$ that $\dim(E_{1})\geq 1$.
@@
 Since the image of $r_{2}(E_{1})$ under ${\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\rightarrow{\mathbb{P}}^{N}$ is $\left\{a\right\}$, there is $C\subset({\mathbb{P}}^{m})^{k}$ of positive dimension such that $r_{2}(E_{1})=\left\{a\right\}\times C\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}$.
-Since $a\in{\mathbb{P}}^{\beta}$ and $\rho_{2}(r_{2}(E_{1}))\subset\rho_{2}(r_{2}(\tilde{I}))=\overline{\rho_{2}(\rho_{1}(I))}=I_{(m)}$, we have $\left\{a\right\}\times C=\rho_{2}(\left\{a\right\}\times C)\subset I_{(m)}$.
-∎
+Since $a\in{\mathbb{P}}^{\beta}$ and
+$\rho_{2}(r_{2}(E_{1}))\subset\rho_{2}(r_{2}(\tilde{I}))=\overline{\rho_{2}(\rho_{1}(I))}=I_{(m)},$
+we have $\left\{a\right\}\times C=\rho_{2}(\left\{a\right\}\times C)\subset I_{(m)}$. ∎
 Lemma 10 (Non-triviality of subsecant varieties).
@@
 For a general point $b_{0}\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, we take an irreducible curve $C\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ such that $a,b_{0}\in C$.
-Let $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ be the linear projection in Remark 8 (a).
-Since $a\in{\mathbb{P}}^{\beta}$, we have $a=\pi_{2}(a)\in C^{\prime}:=\overline{\pi_{2}(C)}\subset{\mathbb{P}}^{\beta}$.
+Let $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ be the linear projection in Remark 8 (a), and let $C^{\prime}=\overline{\pi_{2}(C)}\subset{\mathbb{P}}^{\beta}$.
+Since $a\in{\mathbb{P}}^{\beta}$, we have $a=\pi_{2}(a)\in C^{\prime}$.
 Since $b_{0}$ is general, for a general point $b\in C$, we have $b\in\langle\,x_{1},\dots,x_{k-1}\,\rangle$ with $x_{1},\dots,x_{k-1}\in v_{d}({\mathbb{P}}^{n})$.
 Take $x_{i}^{\prime}\in{\mathbb{P}}^{n}$ with $x_{i}=v_{d}(x_{i}^{\prime})$.
-As in Remark 8 (a), we take $y_{i}^{\prime}\in L$ for each $i=1,\dots,k-1$ and say $y_{i}=v_{d}(y_{i}^{\prime})$.
-Then $\pi_{2}(x_{i})=y_{i}$.
+Setting $y_{i}=\pi_{2}(x_{i})$, we have $y_{i}=v_{d}(y_{i}^{\prime})$ with $y_{i}^{\prime}\in L$ for each $i=1,\dots,k-1$ as in Remark 8 (a).
 It follows that $\pi_{2}(b)\in\langle\,y_{1},\dots,y_{k-1}\,\rangle\subset\sigma_{k-1}(v_{d}(L))$.
@@
     as a set.
-For one inclusion is obvious, let us prove $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle\subset\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$.
+Since one inclusion is obvious, let us prove $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\cap\langle v_{d}({\mathbb{P}}^{m})\rangle\subset\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$.
 Suppose that it does not hold.
@@
 Thus, the equality in (8) is true.
-    In other words, for any $d$-th Veronese embedding $X=v_{d}({\mathbb{P}}^{n})$ and the linear span $L=\langle v_{d}({\mathbb{P}}^{m})\rangle$ of any sub-Veronese variety $v_{d}({\mathbb{P}}^{m})$, we showed that $(X,L)$ is a border rank preserving pair for any $k,d>0$ in the terminology of [21, definition 5.7.3.1] (we’d also like to note that [5, theorem 1.1] can imply the same result for any $d\geq k$ in case of ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$).
+    In other words, for any $d$-th Veronese embedding $X=v_{d}({\mathbb{P}}^{n})$ and the linear span $L=\langle v_{d}({\mathbb{P}}^{m})\rangle$ of any sub-Veronese variety $v_{d}({\mathbb{P}}^{m})$, we showed that $(X,L)$ is a border rank preserving pair for any $k,d>0$ in the terminology of [Land12, Definition 5.7.3.1] (we would also like to note that [BGL13, Theorem 1.1] can imply the same result for any $d\geq k$ in case of ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$).
   2. (b)
-    (Every $\Sigma_{k,d}(m)$ is closed) Recall that the symmetric subspace variety $Sub_{m}(S^{d}V)$ ([21, section 7.1.3]) is defined as
+    (Every $\Sigma_{k,d}(m)$ is closed) Recall that the symmetric subspace variety $Sub_{m}(S^{d}V)$ ([Land12, section 7.1.3]) is defined as
     $\{f\in{\mathbb{P}}S^{d}V~{}|~{}\exists~{}W\subset V,~{}\dim W=m+1~{},~{}f\in{\mathbb{P}}S^{d}W\}~{}.$
@@
 For an $m$-dimensional projective variety $Z\subset{\mathbb{P}}^{\beta}$, we take another incidence variety $J\subset{\mathbb{P}}^{\beta}\times(Z)^{k}$ to be the Zariski closure of
-(5) $J^{0}:=\left\{(a,x_{1},\dots,x_{k})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
-with the projections $p:J\rightarrow\sigma_{k}(Z)\subset{\mathbb{P}}^{\beta}$ and
-$q_{i}:J\rightarrow(Z)^{k}\rightarrow Z,$
+(9) $J^{0}=\left\{(a,x_{1},\dots,x_{k})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
+with the projections
+$p:J\rightarrow\sigma_{k}(Z)\subset{\mathbb{P}}^{\beta},\quad q_{i}:J\rightarrow(Z)^{k}\rightarrow Z,$
 where $(Z)^{k}\rightarrow Z$ is the projection to the $i$-th factor for $1\leq i\leq k$.
@@
 For $a\in\sigma_{k}(Z)$, the scheme-theoretic image $q_{i}(p^{-1}(a))$ in $Z$ is called the ($k$-th) entry locus of $Z$ with respect to $a$ in the literature.
-It is known that for a general $a\in\sigma_{k}(Z)$ the locus $q_{i}(p^{-1}(a))$ is equidimensional and moreover, if $Z$ is smooth and in the characteristic 0, then $p^{-1}(a)$ is smooth so that $q_{i}(p^{-1}(a))$ is reduced (see [28, definition 1.4.5]).
+It is known that for a general $a\in\sigma_{k}(Z)$ the locus $q_{i}(p^{-1}(a))$ is equidimensional and moreover, if $Z$ is smooth and in the characteristic 0, then $p^{-1}(a)$ is smooth so that $q_{i}(p^{-1}(a))$ is reduced (see [Rus16, Definition 1.4.5]).
 Let $Z,X\subset{\mathbb{P}}^{N}$ be projective varieties of dimensions $m,n$.
@@
 Then, for a general point $x\in q_{i}(F)$ with $1\leq i\leq k$, we have
-$\mathbb{T}_{x}(X)\subset\mathbb{T}_{a}(\sigma_{k}(X))~{},$
+$\mathbb{T}_{x}(X)\subset\mathbb{T}_{a}(\sigma_{k}(X)),$
 where $\mathbb{T}_{x}(X)\subset{\mathbb{P}}^{N}$ means the embedded tangent space to $X$ at $x$.
 Proof.
-Since $\sigma_{k}(Z)\setminus\operatorname{Sing}(\sigma_{k}(X))$ is non-empty open in $\sigma_{k}(Z)$ and $a$ is general, it is a smooth point of $\sigma_{k}(X)$, and also is contained in the $(k-1)$-plane $\langle\,x_{1},\dots,x_{k}\,\rangle$ for general $x_{1},\dots,x_{k}\in Z$ with $x_{i}=x$.
+Since $\sigma_{k}(Z)\setminus\operatorname{Sing}(\sigma_{k}(X))$ is non-empty open in $\sigma_{k}(Z)$ and $a$ is general, it is a smooth point of $\sigma_{k}(X)$; hence the embedded tangent space $\mathbb{T}_{a}(\sigma_{k}(X))$ is defined.
+In addition, $a$ is contained in the $(k-1)$-plane $\langle\,x_{1},\dots,x_{k}\,\rangle$ for general $x_{1},\dots,x_{k}\in Z$ with $x_{i}=x$.
 Then the assertion follows by Terracini’s lemma (cf.
-[28, corollary 1.4.2], [29, Chapter II, 1.10]).
+[Rus16, Corollary 1.4.2], [Zak93, Chapter II, 1.10, Chapter V, 1.4]).
 ∎
 Lemma 13.
-For a projective variety $Z$, let $\langle\,Z\,\rangle={\mathbb{P}}^{\beta}$ and consider the incidence $J$ as (5).
+For a projective variety $Z$, let $\langle\,Z\,\rangle={\mathbb{P}}^{\beta}$ and consider the incidence $J$ as (9) with $k\geq 2$.
 Suppose that the ($k-1$)-secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$.
@@
 Let $J_{1}=q_{1}^{-1}(x_{1})\subset J$.
-For the projection $q_{2,\dots,k}:J\rightarrow(Z)^{k-1},(a,\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\mapsto(\tilde{x}_{2},\dots,\tilde{x}_{k})$, a general fiber of $q_{2,\dots,k}|_{J_{1}}:J_{1}\rightarrow(Z)^{k-1}$ at $(\tilde{x}_{2},\dots,\tilde{x}_{k})$ is isomorphic to the ($k-1$)-plane $\langle\,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$.
-In particular, $\pi_{x_{1}}\circ p|_{J_{1}}:J_{1}\rightarrow\pi_{x_{1}}(\sigma_{k-1}(Z))$ is surjective.
-Let $\alpha\in\pi_{x_{1}}(\sigma_{k-1}(Z))$ be a general point.
-Then we may consider the element $(a,x_{1},x_{2},\dots,x_{k})$ as a general point in $J_{1}$ with $\alpha=\pi_{x_{1}}(a)$.
-Let $F$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},x_{2},\dots,x_{k})$, and suppose that $q_{1}|_{F}$ has fibers of positive dimensions.
+Then $\pi_{x_{1}}\circ p|_{J_{1}}:J_{1}\rightarrow\pi_{x_{1}}(\sigma_{k-1}(Z))$ is dominant (this is because, for general $b\in\sigma_{k-1}(Z)$, we take a general point $c\in\langle\,x_{1},b\,\rangle$ and take $k-1$ points $\tilde{x}_{2},\dots,\tilde{x}_{k}\in Z$ such that $b\in\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$; then $(c,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in J_{1}$, whose image under $\pi_{x_{1}}\circ p$ is $\pi_{x_{1}}(c)=\pi_{x_{1}}(b)$).
+Since $(a,x_{1},\dots,x_{k})\in J_{1}$ is general in $J$ and by the dominance of $\pi_{x_{1}}\circ p|_{J_{1}}$, we may consider $\alpha=\pi_{x_{1}}(a)$ as a general point in $\pi_{x_{1}}(\sigma_{k-1}(Z))$.
+Let $F$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},x_{2},\dots,x_{k})$, and suppose that $q_{1}|_{F}$ has general fibers of positive dimensions.
 Then, the fiber of $q_{1}|_{F}$ at $x_{1}\in q_{1}(F)$,
 $F\cap J_{1}$
-is of positive dimension.
-For $(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}$, considering a line $\langle\,a,x_{1}\,\rangle\subset\langle\,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$ and a hyperplane $\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$, we may take a point $\tilde{b}\in\langle\,a,x_{1}\,\rangle\cap\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$.
-Note that $\pi_{x_{1}}(\tilde{b})=\pi_{x_{1}}(a)=\alpha$.
-It gives the following locus of positive dimension,
-$\left\{(\tilde{b},\tilde{x}_{2},\dots,\tilde{x}_{k})\mid(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}\right\}\subset\rho^{-1}(\alpha)\subset J^{\prime},$
+is of positive dimension, which means that we have $(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}$ for fixed $a,x_{1}$ and moving $(\tilde{x}_{2},\dots,\tilde{x}_{k})$ with positive dimension.
+For general $(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}$, we have the intersection point $b(\tilde{x}_{2},\dots,\tilde{x}_{k})$ of the line $\langle\,a,x_{1}\,\rangle$ and the hyperplane $\langle\,\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle$ in $\langle\,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k}\,\rangle={\mathbb{P}}^{k-1}$.
+Note that
+$\pi_{x_{1}}(b(\tilde{x}_{2},\dots,\tilde{x}_{k}))=\pi_{x_{1}}(a)=\alpha.$
+The $k$-tuple $(b(\tilde{x}_{2},\dots,\tilde{x}_{k}),\tilde{x}_{2},\dots,\tilde{x}_{k})$ moves with positive dimension since so does ($k-1$)-tuple $(\tilde{x}_{2},\dots,\tilde{x}_{k})$.
+In other words, the following locus is of positive dimension,
+$\left\{(b(\tilde{x}_{2},\dots,\tilde{x}_{k}),\tilde{x}_{2},\dots,\tilde{x}_{k})\mid(a,x_{1},\tilde{x}_{2},\dots,\tilde{x}_{k})\in F\cap J_{1}\right\}\subset\rho^{-1}(\alpha)\subset J^{\prime},$
 which contradicts that $\rho$, given in (10), is generally finite. ∎
@@
 Take a general $a\in\sigma_{k}(Z)={\mathbb{P}}^{\beta}$.
-Since $\operatorname{Vertex}(\sigma_{k-1}(Z))$ is a linear subvariety of $\sigma_{k-1}(Z)\subsetneq{\mathbb{P}}^{\beta}$ and $Z$ is non-degenerate in ${\mathbb{P}}^{\beta}$, $\operatorname{Vertex}(\sigma_{k-1}(Z))\cap Z\subsetneq Z$.
-This implies that for general $b\in Z$, $\dim\langle\,b,\sigma_{k-1}(Z)\,\rangle>\dim(\sigma_{k-1}(Z))$ so that $\langle\,b,\sigma_{k-1}(Z)\,\rangle={\mathbb{P}}^{\beta}$.
-Thus, any given general point $a\in{\mathbb{P}}^{\beta}$ sits on a line $\langle\,b,c\,\rangle$ for any general $b\in Z$ and for some general $c\in\sigma_{k-1}(Z)$ and $(a,b,\tilde{x_{2}},\dots,\tilde{x_{k}})\in p^{-1}(a)$, which means $q_{1}(p^{-1}(a))=Z$.
+Since $\operatorname{Vertex}(\sigma_{k-1}(Z))$ is a linear subvariety of $\sigma_{k-1}(Z)\subsetneq{\mathbb{P}}^{\beta}$ and $Z$ is non-degenerate in ${\mathbb{P}}^{\beta}$,
+$\operatorname{Vertex}(\sigma_{k-1}(Z))\cap Z\subsetneq Z.$
+This implies that for general $x\in Z$, $\dim\langle\,x,\sigma_{k-1}(Z)\,\rangle>\dim(\sigma_{k-1}(Z))$ so that $\langle\,x,\sigma_{k-1}(Z)\,\rangle={\mathbb{P}}^{\beta}$.
+Thus, any given general point $a\in{\mathbb{P}}^{\beta}$ sits on a line $\langle\,x,b\,\rangle$ for any general $x\in Z$ and for some general $b\in\sigma_{k-1}(Z)$.
+Taking $k-1$ points $\tilde{x_{2}},\dots,\tilde{x_{k}}\in Z$ such that $b\in\langle\,\tilde{x_{2}},\dots,\tilde{x_{k}}\,\rangle$, we have $(a,x,\tilde{x_{2}},\dots,\tilde{x_{k}})\in p^{-1}(a)$, which means $q_{1}(p^{-1}(a))=Z$.
 ∎
 Now let us focus on the case $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta=\binom{m+d}{m}-1}$, the image of the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$.
-Here we prove a very useful proposition, which has an independent interest itself.
+Here we prove a very useful proposition, which is of independent interest itself.
 In Proposition 15, we consider the entry locus of a general point in $Z$ and estimate the dimension of the linear span of its image under ($d-1$)-uple Veronese embedding $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}=\binom{m+d-1}{m}-1}$.
@@
 Let $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta}$ with $d\geq 3$, $22\leq m\leq k-$ 2, and $\beta=\binom{m+d}{m}-1<km+k-1$.
-Assume $\dim(\sigma_{k-1}(Z))=(k-1)m+k-2<\beta$ (in other words, the ($k-1$)-th secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$).
+Assume $\dim(\sigma_{k-1}(Z))=(k-1)m+k-2<\beta$ (in other words, the ($k-1$)-th secant of $Z$ is not defective and not equal to ${\mathbb{P}}^{\beta}$), where we have $(k-1)m+k\leq\binom{m+d}{m}$.
 Let $J\subset{\mathbb{P}}^{\beta}\times(Z)^{k}$ be the Zariski closure of the incidence (9), let $(a,x_{1},\dots,x_{k})\in J$ be a general point, and let $F\subset J$ be an irreducible component of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$.
@@
   2. (ii)
-    If $(k-1)m+k=\binom{m+d}{m}$, then it holds $q_{i}(F)=Z$. In addition, if $(d,m)\neq(3,2)$, then $\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle\geq k+m$.
+    If $(k-1)m+k=\binom{m+d}{m}$, then $q_{i}(F)=Z$. In addition, if $(d,m)\neq(3,2)$, then $\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle\geq k+m$.
 Remark 16.
   1. (a)
-    Let $\mu=\frac{\binom{m+d}{m}}{m+1}$.
-Then inequalities $\beta<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ (i.e., $\binom{m+d}{m}<km+k$ and $(k-1)m+k-1<\binom{m+d}{m}$) are equivalent to $\mu<k<\mu+1$; it happens if and only if $\mu\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$.
+    Two inequalities $\beta=\binom{m+d}{m}-1<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ are equivalent to $\frac{\binom{m+d}{m}}{m+1}<k<\frac{\binom{m+d}{m}}{m+1}+1$; it occurs if and only if $\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\Bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}\Bigl{\rceil}$.
   2. (b)
@@
 Lemma 17.
-Let $d,m,k$ be integers such that $d\geq 3$, $22\leq m\leq k-$ 2, and $\binom{m+d}{m}-1<km+k-1$.
+Let $d,m,k$ be integers such that $d\geq 3$, $22\leq m\leq k-$ 2.
   1. (i)
-    If $(k-1)m+k<\binom{m+d}{m}$, then $\binom{m+d-1}{m}-1-2m-k\geq 0$.
+    If $(k-1)m+k<\binom{m+d}{m}<km+k$, then $\binom{m+d-1}{m}-1-2m-k\geq 0$.
   2. (ii)
     If $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)\neq(3,2)$, then $\binom{m+d-1}{m}-1-m-k\geq 0$.
-Proof.
-(i) From Remark 18 (a) below, we may assume $m\geq 3$. Let $\delta:=\binom{m+d-1}{m}-(1+k+2m)$. Since $\binom{m+d-1}{m}=\frac{d}{m+d}\binom{m+d}{m}$, using $(k-1)m+k+1\leq\binom{m+d}{m}$, we have
-$\delta=\binom{m+d-1}{m}-(1+k+2m)\geq\frac{1}{m+d}\Bigl{(}d((k-1)m+k+1)-(m+d)(1+k+2m)\Bigr{)},$
-Setting $m=k-a$ with $a\geq 2$, we have
-(7) $d((k-1)m+k+1)-(m+d)(1+k+2m)=km(d-1)-3dm-m-2m^{2}\\ =(m+a)m(d-1)-3dm-m-2m^{2}=m(dm+(a-3)d-3m-(a+1))\\ \geq m((d-3)(m-1)-6)$
-Then, $\delta\geq 0$ holds in the following three cases: $d\geq 6$ and $m\geq 3$; $d=5$ and $m\geq 4$; or $d=4$ and $m\geq 7$.
-In addition, in Remark 18 (b) (see also Remark 16 (a)), we explicitly check $\delta\geq 0$ if $d=4$ and $m\leq 6$, and check that there is no $k$ in our range if $d=5$ and $m=3$.
-On the other hands, when $d=3$, (7) implies
-$d((k-1)m+k+1)-(m+d)(1+k+2m)=m((a-3)3-(a+1))=m(2a-10).$
-Hence, $\delta\geq 0$ holds if $d=3$ and $a\geq 5$. For $d=3$ and $m=k-a$ with $a=2,3,4$, we have
-$\delta=\frac{m^{2}-m-2k}{2}=\frac{m(m-3)-2a}{2}>0$
-if $m\geq 5$. In addition, in Remark 18 (b), we explicitly check that there is no $k$ in our range if $d=3$ and $m\leq 4$.
-(ii) Let $\delta:=\binom{m+d-1}{m}-(1+k+m)$. In the same way as (i), using $(k-1)m+k=\binom{m+d}{m}$, we have
-$\delta=\binom{m+d-1}{m}-(1+k+m)=\frac{1}{m+d}\Bigl{(}d((k-1)m+k)-(m+d)(1+k+m)\Bigr{)},$
-and have
-$d((k-1)m+k)-(m+d)(1+k+m)=km(d-1)-2dm-m-m^{2}-d\\ \geq(m+2)m(d-1)-2dm-m-m^{2}-d=(d-2)(m^{2}-1)-2-3m.$
-If $d\geq 3$ and $m\geq 4$, then since $(d-2)(m^{2}-1)-2-3m\geq m(m-3)-3\geq 1$, it holds $\delta\geq 0$.
-If $d\geq 4$ and $m=3$, then since $(d-2)(m^{2}-1)-2-3m\geq 5$, it similarly holds $\delta\geq 0$.
-On the other hands, if $d=3$ and $m=3$, then $(k-1)m+k=\binom{m+d}{m}$ implies $204k-3=\frac{(d+3)(d+2)(d+1)}{3!}=$ 20; this case does not occur since $k$ cannot be an integer.
-If $d\geq 4$ and $m=2$, then $3k-2=\frac{(d+2)(d+1)}{2}$. In this case, $\delta=\frac{(d+1)d}{2}-k-3\geq 0$, because of
-$3\left(\frac{(d+1)d}{2}-k-3\right)\geq\frac{3(d+1)d}{2}-\frac{(d+2)(d+1)}{2}-11=d^{2}-12\geq 4$
+  3. (iii)
+    If $\binom{m+d}{m}=km+k$, then $\binom{m+d-1}{m}-1-m-k\geq 0$.
+Note that (iii) of Lemma 17 is applied in a discussion of the proof of Part (ii) of Theorem 2, though it is not used in this section.
+To show the lemma, we need some calculations as follows.
 Remark 18.
-In the proof of Lemma 17, we refer to two calculations as follows.
-  1. (a)
-    Let $m=2$ and $km+k-1>\binom{m+d}{m}-1$.
+(a) Let $m=2$ and $km+k>\binom{m+d}{m}$.
 Then $(k-1)m+k<\binom{m+d}{m}$ does not occur.
@@
 modulo $3$, which is absurd.
-  2. (b)
-    For $d=3,4,5$, we calculate numbers $m$ such that $\mu:=\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$ satisfies $(k-1)m+k<\beta$.
+(b) For $d=3,4,5$, we calculate numbers $m$ satisfying the conditions $\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$, $k=\Bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}\Bigl{\rceil}$, and $(k-1)m+k<\binom{m+d}{m}$.
 For $d=3$, the smallest $m=5$.
@@
 $(d,m,k,\delta)=(3,5,10,0),(4,3,9,4),(4,7,42,63),(5,9,201,495).$
+Proof of Lemma 17.
+(i) From Remark 18 (a), we may assume $m\geq 3$. Let $\delta=\binom{m+d-1}{m}-(1+k+2m)$. Since $\binom{m+d-1}{m}=\frac{d}{m+d}\binom{m+d}{m}$, using $(k-1)m+k+1\leq\binom{m+d}{m}$, we have
+$\delta=\binom{m+d-1}{m}-(1+k+2m)\geq\frac{1}{m+d}\Bigl{(}d((k-1)m+k+1)-(m+d)(1+k+2m)\Bigr{)}.$
+Setting $k=m+a$ with $a\geq 2$, we have
+(11) $d((k-1)m+k+1)-(m+d)(1+k+2m)=m(k(d-1)-3d-1-2m)\\ =m((m+a)(d-1)-3d-1-2m)\geq m((m+2)(d-1)-3d-1-2m)\\ =m(dm-3m-d-3)=m((d-3)(m-1)-6).$
+Then, $\delta\geq 0$ holds in the following three cases: $d\geq 6$ and $m\geq 3$; $d=5$ and $m\geq 4$; or $d=4$ and $m\geq 7$.
+In addition, in Remark 18 (b) (see also Remark 16 (a)), we explicitly check that $\delta\geq 0$ if $d=4$ and $m\leq 6$, and check that there is no $k$ in our range if $d=5$ and $m=3$.
+On the other hand, when $d=3$, (11) implies
+$d((k-1)m+k+1)-(m+d)(1+k+2m)=m(2(m+a)-10-2m)=m(2a-10).$
+Hence, $\delta\geq 0$ holds if $d=3$ and $a\geq 5$. For $d=3$ and $k=m+a$ with $a=2,3,4$, we have
+$\delta=\frac{m^{2}-m-2k}{2}=\frac{m(m-3)-2a}{2}>0$
+if $m\geq 5$. In addition, in Remark 18 (b), we explicitly check that there is no $k$ in our range if $d=3$ and $m\leq 4$.
+(ii) Let $\delta=\binom{m+d-1}{m}-(1+k+m)$. In the same way as (i), using $(k-1)m+k=\binom{m+d}{m}$, we have
+$\delta=\binom{m+d-1}{m}-(1+k+m)=\frac{1}{m+d}\Bigl{(}d((k-1)m+k)-(m+d)(1+k+m)\Bigr{)},$
+and
+$d((k-1)m+k)-(m+d)(1+k+m)=km(d-1)-2dm-m-m^{2}-d\\ \geq(m+2)m(d-1)-2dm-m-m^{2}-d=(d-2)(m^{2}-1)-2-3m.$
+If $d\geq 3$ and $m\geq 4$, then since $(d-2)(m^{2}-1)-2-3m\geq m(m-3)-3\geq 1$, it holds $\delta\geq 0$.
+If $d\geq 4$ and $m=3$, then since $(d-2)(m^{2}-1)-2-3m\geq 5$, it similarly holds $\delta\geq 0$.
+On the other hand, if $d=3$ and $m=3$, then $(k-1)m+k=\binom{m+d}{m}$ implies
+$4k-3=\frac{(d+3)(d+2)(d+1)}{3!}=20;$
+this case does not occur since $k$ cannot be an integer.
+If $d\geq 4$ and $m=2$, then $3k-2=\frac{(d+2)(d+1)}{2}$. In this case, $\delta=\frac{(d+1)d}{2}-k-3\geq 0$, because of
+$3\left(\frac{(d+1)d}{2}-k-3\right)\geq\frac{3(d+1)d}{2}-\frac{(d+2)(d+1)}{2}-11=d^{2}-12\geq 4$
+(iii) Next, we assume $\binom{m+d}{m}=km+k$. Then,
+$\delta=\binom{m+d-1}{m}-(1+k+m)=\frac{1}{m+d}\Bigl{(}d(km+k)-(m+d)(1+k+m)\Bigr{)}.$
+Using $k=m+a=(m+1)+(a-1)$ with $a\geq 2$, we have
+$d(km+k)-(m+d)(1+k+m)\\ =(d-1)km-(m+d)(m+1)=(d-1)m(m+1)+(d-1)m(a-1)-(m+d)(m+1)\\ =(m+1)((d-1)m-(m+d))+(d-1)m(a-1)\\ =(m+1)((d-2)(m-1)-2)+(d-1)m(a-1).$
+If $d\geq 3,m\geq 2$ and $(d,m)\neq(3,2)$, then we have $\delta\geq 0$. If $(d,m)=(3,2)$, then since
+$(m+1)((d-2)(m-1)-2)+(d-1)m(a-1)=-3+4(a-1)\geq 1,$
+we also have $\delta\geq 0$. ∎
 The next lemma concerns a general fact on linear sections of Veronese varieties, which has an independent interest itself.
 Lemma 19.
-Let $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ be general $k$ points, let $v_{e}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{e}=\binom{m+e}{m}-1}$ be the $e$-uple Veronese embedding of ${\mathbb{P}}^{m}$, and let $R\subset{\mathbb{P}}^{\beta_{e}}$ be any $k$-plane containing the ($k-1$)-plane $M:=\langle\,v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle$.
-Then the following holds.
+Let $k,m\geq 2$.
+Let $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ be general $k$ points, let $v_{e}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{e}=\binom{m+e}{m}-1}$ be the $e$-uple Veronese embedding of ${\mathbb{P}}^{m}$, and let $M=\langle\,v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle\subset{\mathbb{P}}^{\beta_{e}}$.
+Then, for any $k$-plane $R\subset{\mathbb{P}}^{\beta_{e}}$ containing the ($k-1$)-plane $M$, the following holds.
   1. (i)
@@
     Assume $e\geq 3$ and $k\leq\beta_{e}-2m$. Then we have
-    $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0~{},$
-    where the left-hand side means dimension of component(s) passing through $v_{e}(x_{1}^{\prime})$.
+    $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0,$
+    where the left hand side means dimension of component(s) passing through $v_{e}(x_{1}^{\prime})$.
 In particular, the set of a point $\left\{v_{e}(x_{1}^{\prime})\right\}$ is an irreducible component of ${R\cap v_{e}({\mathbb{P}}^{m})}$.
@@
 be the linear projection from the ($k-2$)-plane $\langle\,v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\,\rangle$.
-If $k\leq\beta_{e}-m$, then the generalized trisecant lemma ([28, proposition 1.4.3]) implies $M\cap v_{e}({\mathbb{P}}^{m})=\left\{v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\right\}$; in particular, $C\not\subset M$.
-Hence $C$ is not contracted to a point under $\pi$.
-Then $\overline{\pi(C)}\subset{\mathbb{P}}^{\beta_{e}-k+1}$ is a curve passing through $\pi(v_{e}(x_{1}^{\prime}))$, and then $\overline{\pi(C)}=\overline{\pi(R)}$, which is indeed a line.
-It follows
-$\overline{\pi(C)}=\overline{\pi(R)}\subset\mathbb{T}_{\pi(v_{e}(x_{1}^{\prime}))}\overline{\pi(v_{e}({\mathbb{P}}^{m}))}$
+If $k\leq\beta_{e}-m$, then the generalized trisecant lemma ([Rus16, Proposition 1.4.3]) implies
+$M\cap v_{e}({\mathbb{P}}^{m})=\left\{v_{e}(x_{1}^{\prime}),v_{e}(x_{2}^{\prime}),\dots,v_{e}(x_{k}^{\prime})\right\}.$
+In particular, $\dim(M\cap v_{e}({\mathbb{P}}^{m}))=0$ and $C\not\subset M$.
+We have $\pi(v_{e}(x_{1}^{\prime}))\in\overline{\pi(C)}\subset{\mathbb{P}}^{\beta_{e}-k+1}$ because of $v_{e}(x_{1}^{\prime})\in C$.
+If $C$ is contracted to a point under $\pi$, then $\overline{\pi(C)}=\pi(v_{e}(x_{1}^{\prime}))$, which means that $C\subset M$, a contradiction.
+Hence $\overline{\pi(C)}$ must be a curve.
+For the $k$-plane $R\subset{\mathbb{P}}^{\beta_{e}}$, which contains the ($k-2$)-dimensional center of $\pi$, the image $\overline{\pi(R)}$ is a line in ${\mathbb{P}}^{\beta_{e}-k+1}$.
+Thus $\overline{\pi(C)}=\overline{\pi(R)}$.
+Moreover, it follows
+$\overline{\pi(C)}=\overline{\pi(R)}=\mathbb{T}_{\pi(v_{e}(x_{1}^{\prime}))}\overline{\pi(R)}\subset\mathbb{T}_{\pi(v_{e}(x_{1}^{\prime}))}\overline{\pi(v_{e}({\mathbb{P}}^{m}))}$
 where, by generic smoothness, the right hand side is equal to $\pi({\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})})$ since $x_{1}^{\prime}\in{\mathbb{P}}^{m}$ is general.
@@
 $\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}:{\mathbb{P}}^{\beta_{e}}\dashrightarrow{\mathbb{P}}^{\beta_{e}-m-1}$
-from the $m$-plane ${\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}\subset{\mathbb{P}}^{\beta_{e}}$, and its restriction $\tilde{\pi}=\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}|_{v_{e}({\mathbb{P}}^{m})}$ on the Veronese variety.
-Note that $v_{e}({\mathbb{P}}^{m})\cap{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}=\left\{v_{e}(x_{1}^{\prime})\right\}$.
+from the $m$-plane ${\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}\subset{\mathbb{P}}^{\beta_{e}}$, and its restriction $\tilde{\pi}=\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}|_{v_{e}({\mathbb{P}}^{m})}$ on $v_{e}({\mathbb{P}}^{m})$.
+Note that the Veronese variety $v_{e}({\mathbb{P}}^{m})$ and any embedded tangent space to $v_{e}({\mathbb{P}}^{m})$ intersect only at one point; in particular, $v_{e}({\mathbb{P}}^{m})\cap{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}=\left\{v_{e}(x_{1}^{\prime})\right\}$.
 If $R\subset{\mathbb{P}}^{\beta}$ satisfies (12), we have
 (13) $\tilde{\pi}(R\cap v_{e}({\mathbb{P}}^{m}))\subset\tilde{\pi}(v_{e}({\mathbb{P}}^{m}))\cap\langle\,\tilde{\pi}(v_{e}(x_{2}^{\prime})),\dots,\tilde{\pi}(v_{e}(x_{k}^{\prime}))\,\rangle.$
-Let $e\geq 3$ and $k\leq\beta_{e}-2m$.
-Suppose that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})>0$, which means the existence of a curve $C$ satisfying the condition of (i).
-Since $\operatorname{codim}(\tilde{\pi}(v_{e}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{e}-m-1})-(k-1)\geq(\beta_{e}-2m-1)-(k-1)\geq 0$, again the trisecant lemma implies that the right hand side in (9) is only the set of $k-1$ points ${\tilde{\pi}(v_{e}(x_{2}^{\prime})),\dots,\tilde{\pi}(v_{e}(x_{k}^{\prime}))}$.
+By Terracini’s lemma, for general $z\in v_{e}({\mathbb{P}}^{m})$, the linear variety $\langle\,\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m}),\mathbb{T}_{z}v_{e}({\mathbb{P}}^{m})\,\rangle$ coincides with an embedded tangent space to $\sigma_{2}(v_{e}({\mathbb{P}}^{m}))$, and is of dimension $\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))$.
+Then,
+$\pi_{\mathbb{T}_{v_{e}(x_{1}^{\prime})}v_{e}({\mathbb{P}}^{m})}(\mathbb{T}_{z}v_{e}({\mathbb{P}}^{m}))=\mathbb{T}_{\tilde{\pi}(z)}\tilde{\pi}(v_{e}({\mathbb{P}}^{m}))\subset{\mathbb{P}}^{\beta_{e}-m-1}$
+is of dimension $\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))-m-1$. It follows
+(14) $\dim\tilde{\pi}(v_{e}({\mathbb{P}}^{m}))=\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))-m-1.$
+Let $e\geq 3$ and $k\leq\beta_{e}-2m$. Suppose that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})>0$, which means the existence of a curve $C$ satisfying the condition of (i). Since
+$\operatorname{codim}(\tilde{\pi}(v_{e}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{e}-m-1})-(k-1)\geq(\beta_{e}-2m-1)-(k-1)\geq 0,$
+again the trisecant lemma implies that the right hand side in (13) is only the set of $k-1$ points ${\tilde{\pi}(v_{e}(x_{2}^{\prime})),\dots,\tilde{\pi}(v_{e}(x_{k}^{\prime}))}$.
 Thus, each irreducible subset $D\subset R\cap v_{e}({\mathbb{P}}^{m})$ satisfies
@@
 (At least, taking $D=C$, we exactly have (15).) From $e\geq 3$, $\dim\sigma_{2}(v_{e}({\mathbb{P}}^{m}))=2m+1$ (i.e., non-defective).
-Then $\tilde{\pi}$ must be a generically finite map since general two tangent spaces of $\sigma_{2}(v_{e}({\mathbb{P}}^{m}))$ do not meet each other, and we reach a contradiction since $v_{e}(x_{l}^{\prime})$ is a general point.
+In this case, by (14), $\dim\tilde{\pi}(v_{e}({\mathbb{P}}^{m}))=m$, i.e., the map $\tilde{\pi}$ must be generically finite.
+Then we reach a contradiction since $v_{e}(x_{l}^{\prime})$ is a general point.
 It implies that $\dim_{v_{e}(x_{1}^{\prime})}({R\cap v_{e}({\mathbb{P}}^{m})})=0$.
 Finally, we consider the case of $e=2$.
-Then it is well known that the tangential projection $\tilde{\pi}:v_{2}({\mathbb{P}}^{m})\dashrightarrow\tilde{\pi}(v_{2}({\mathbb{P}}^{m}))$ has fibers of dimension $1$.
-Moreover, as in Remark 8 (a) we know that $\pi_{\mathbb{T}_{v_{2}(x_{1}^{\prime})}v_{2}({\mathbb{P}}^{m})}:{\mathbb{P}}^{\beta_{2}=\frac{(m+2)(m+1)}{2}-1}\dashrightarrow{\mathbb{P}}^{\beta_{2}-m-1=\frac{(m+1)m}{2}-1}$ satisfies the commutative diagram:
+Then $\dim\sigma_{2}(v_{2}({\mathbb{P}}^{m}))=2m$ for $m\geq 2$ (i.e., defective), and by (14), $\dim\tilde{\pi}(v_{2}({\mathbb{P}}^{m}))=m-1$.
+It means that the tangential projection $\tilde{\pi}:v_{2}({\mathbb{P}}^{m})\dashrightarrow\tilde{\pi}(v_{2}({\mathbb{P}}^{m}))$ has fibers of dimension $1$.
+Moreover, as in Remark 8 (a) we know that
+$\pi_{\mathbb{T}_{v_{2}(x_{1}^{\prime})}v_{2}({\mathbb{P}}^{m})}:{\mathbb{P}}^{\beta_{2}=\frac{(m+2)(m+1)}{2}-1}\dashrightarrow{\mathbb{P}}^{\beta_{2}-m-1=\frac{(m+1)m}{2}-1}$
+satisfies the commutative diagram:
 ${{\mathbb{P}}^{m}}$${{\mathbb{P}}^{\frac{(m+2)(m+1)}{2}-1}}$${{\mathbb{P}}^{m-1}}$${{\mathbb{P}}^{\frac{(m+1)m}{2}-1}\quad,}$$\scriptstyle{v_{2}}$$\scriptstyle{\pi_{x_{1}^{\prime}}}$$\scriptstyle{\pi_{\mathbb{T}_{v_{2}(x_{1}^{\prime})}v_{2}({\mathbb{P}}^{m})}}$$\scriptstyle{v_{2}}$
 where $\pi_{x_{1}^{\prime}}:{\mathbb{P}}^{m}\dashrightarrow{\mathbb{P}}^{m-1}$ is the linear projection from $x_{1}^{\prime}$, and $v_{2}:{\mathbb{P}}^{m-1}\hookrightarrow{\mathbb{P}}^{\frac{(m+1)m}{2}-1}$ is the Veronese embedding of ${\mathbb{P}}^{m-1}$.
-Then $\dim(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})))=\dim(v_{2}({\mathbb{P}}^{m-1}))=m-1$.
-If $k-1\leq\beta_{2}-2m$, then $\operatorname{codim}(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{2}-m-1})-(k-1)\geq(\beta_{2}-2m)-(k-1)\geq 0$; hence the trisecant lemma implies $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$ for some $l=2,\dots,k$ as we discussed for (10).
-In the above diagram, for any $y^{\prime}\in\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle\subset{\mathbb{P}}^{m}$ with $y^{\prime}\neq x_{1}^{\prime}$, we have $\tilde{\pi}(v_{2}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(x_{l}^{\prime}))=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$; indeed, $\tilde{\pi}^{-1}(\tilde{\pi}(v_{2}(x_{l}^{\prime})))=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$.
+Then
+$\dim(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})))=\dim(v_{2}({\mathbb{P}}^{m-1}))=m-1.$
+If $k-1\leq\beta_{2}-2m$, then
+$\operatorname{codim}(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{2}-m-1})-(k-1)\geq(\beta_{2}-2m)-(k-1)\geq 0;$
+hence the trisecant lemma implies $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$ for some $l=2,\dots,k$ as we discussed for (15).
+In the diagram above, for any $y^{\prime}\in\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle\subset{\mathbb{P}}^{m}$ with $y^{\prime}\neq x_{1}^{\prime}$, we have
+$\tilde{\pi}(v_{2}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(x_{l}^{\prime}))=\tilde{\pi}(v_{2}(x_{l}^{\prime}));$
+indeed, $\tilde{\pi}^{-1}(\tilde{\pi}(v_{2}(x_{l}^{\prime})))=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$.
 Since $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$, we have $D\subset v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$.
 ∎
-Now we give a proof for Proposition 15. For simplicity, we set $i=1$. Then $x_{1}\in q_{1}(F)\subset Z=v_{d}({\mathbb{P}}^{m})$.
+We give one calculation before proving Proposition 15.
+Remark 20.
+For $d=3$ and $m\neq 2$, if $\mu_{0}=\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\lceil\mu_{0}\rceil$ (as under the conditions of Proposition 15 and Remark 16 (a)), then we have $(km+k-1)-\beta_{d}\neq 1$.
+The reason is as follows.
+First, we may write $\mu_{0}={(m+3)(m+2)}/6=M/3$ for some $M\in{\mathbb{N}}$ since $(m+3)(m+2)$ is divisible by $2$.
+In addition, dividing $M$ by $3$ with remainder, we have $M=3Q+R$ for a quotient $Q$ and a remainder $R$.
+Since $\mu_{0}=M/3\notin{\mathbb{N}}$, $R$ must be $1$ or $2$.
+In this setting, $k=\lceil M/3\rceil=Q+1$.
+It follows that $(km+k-1)-\beta_{d}$ is equal to
+$(km+k)-\tbinom{m+3}{3}=(m+1)\left({k-\tfrac{(m+3)(m+2)}{6}}\right)=(m+1)\left((Q+1)-(Q+\tfrac{R}{3})\right)=(m+1)\cdot\tfrac{3-R}{3}.$
+If $(km+k-1)-\beta_{d}=1$, then $3=(m+1)(3-R)$. Since $3-R$ is $1$ or $2$ and $m\in{\mathbb{N}}$, we get $m=2$.
 Proof of Proposition 15.
-(i) Set $s=\dim\sigma_{k}(Z)$. Then an irreducible component $F$ of $p^{-1}(a)$ is of dimension $\dim J-s=(km+k-1)-s$. From Lemma 13, we have $\dim q_{1}(F)=(km+k-1)-s$.
-Let $q_{1}(F)^{\prime}\subset{\mathbb{P}}^{m}$ be the preimage of $q_{1}(F)\subset Z$ under $v_{d}:{\mathbb{P}}^{m}\simeq Z$, and let $A=q_{1}(F)^{\prime}\cup\left\{x_{1}^{\prime},\dots,x_{k}^{\prime}\right\}\subset{\mathbb{P}}^{m}$.
-Let $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}}$ be the ($d-1$)-uple Veronese embedding.
-Then the ($k-1$)-plane $\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset{\mathbb{P}}^{\beta_{d-1}}$ is contained in the linear variety
+(i) For simplicity, we set $i=1$; then $x_{1}\in q_{1}(F)\subset Z=v_{d}({\mathbb{P}}^{m})$.
+For $s=\dim\sigma_{k}(Z)$, an irreducible component $F$ of $p^{-1}(a)$ is of dimension $\dim J-s=(km+k-1)-s$.
+From Lemma 13, we have $\dim q_{1}(F)=(km+k-1)-s$.
+Let $q_{1}(F)^{\prime}\subset{\mathbb{P}}^{m}$ be the preimage of $q_{1}(F)\subset Z$ under $v_{d}:{\mathbb{P}}^{m}\simeq Z$, and let
+$A=q_{1}(F)^{\prime}\cup\left\{x_{1}^{\prime},\dots,x_{k}^{\prime}\right\}\subset{\mathbb{P}}^{m}.$
+Let $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}}$ be the ($d-1$)-uple Veronese embedding. Then the ($k-1$)-plane
+$\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset{\mathbb{P}}^{\beta_{d-1}}$
+is contained in the linear variety
 $\langle\,v_{d-1}(A)\,\rangle=\langle\,v_{d-1}(q_{1}(F)^{\prime})\cup\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}\,\rangle,$
-and is of codimension $c=\dim\langle\,v_{d-1}(A)\,\rangle-(k-1)$.
-By Lemma 17 (i), $\beta_{d-1}-2m-k\geq 0$.
-So, by the generalized trisecant lemma, $v_{d-1}({\mathbb{P}}^{m})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle=\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$.
-In particular, $v_{d-1}(q_{1}(F)^{\prime})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$.
+and is of codimension $c=\dim\langle\,v_{d-1}(A)\,\rangle-(k-1)$. By Lemma 17 (i), $\beta_{d-1}-2m-k\geq 0$. So, by the generalized trisecant lemma,
+$v_{d-1}({\mathbb{P}}^{m})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle=\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}.$
+In particular,
+$v_{d-1}(q_{1}(F)^{\prime})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}.$
 Since $\dim q_{1}(F)^{\prime}\geq 1$, we may take a point $y^{\prime}\in q_{1}(F)^{\prime}$ such that
 $v_{d-1}(y^{\prime})\notin\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle.$
-Assume $d\geq 4$. Applying Lemma 19 (ii) to $R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle\subset\langle\,v_{d-1}(A)\,\rangle$, we have
+Assume $d\geq 4$. Applying Lemma 19 (ii) to
+$R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle\subset\langle\,v_{d-1}(A)\,\rangle,$
+we have
 $\dim_{v_{d-1}(x_{1}^{\prime})}({R\cap v_{d-1}({\mathbb{P}}^{m})})=0.$
 In particular, $\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}(q_{1}(F)^{\prime}))=0$.
-On the other hand, regarding it as an intersection of two irreducible subvarieties in $\langle\,v_{d-1}(A)\,\rangle$, we deduce that every irreducible component of $R\cap v_{d-1}(q_{1}(F)^{\prime})$ is of dimension $\geq\dim(v_{d-1}(q_{1}(F)^{\prime}))-(c-1)$.
-Hence $\dim(\langle\,v_{d-1}(A)\,\rangle)\geq k+\dim(v_{d-1}(q_{1}(F)^{\prime}))=k+(km+k-1)-s$.
+Regarding it as an intersection of two irreducible subvarieties in $\langle\,v_{d-1}(A)\,\rangle$, we deduce that every irreducible component of $R\cap v_{d-1}(q_{1}(F)^{\prime})$ is of dimension $\geq\dim(v_{d-1}(q_{1}(F)^{\prime}))-(c-1)$.
+Hence
+$\dim(\langle\,v_{d-1}(A)\,\rangle)\geq k+\dim(v_{d-1}(q_{1}(F)^{\prime}))=k+(km+k-1)-s.$
 Next, let us consider the case of $d=3$.
@@
 $\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),z\,\rangle$
-is a linear subvariety of dimension $\leq k$.
-Since $v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$ is empty or is a set of points, the intersection $v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$ is of dimension $\leq 1$.
-On the other hand, since $m\neq 2$ by Remark 18 (a), we have $\dim q_{1}(F)^{\prime}=(km+k-1)-\beta\geq 2$ as in Remark 20 below.
-For the union
+is a linear subvariety of dimension $\leq k$. Since $v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$ is empty or is a set of points, the intersection
+$v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$
+is of dimension $\leq 1$. On the other hand, since $m\neq 2$ by Remark 18 (a), we have
+$\dim q_{1}(F)^{\prime}=(km+k-1)-\beta\geq 2$
+as in Remark 20. For the union
 $W=\bigcup_{l=2,\dots,k}v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle\subset{\mathbb{P}}^{\beta_{2}},$
@@
 Taking $D=C$ and applying Lemma 19 (iii), we have $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$ for some $l>1$.
-If $\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k$, then $R=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$, contradicting to the definition of $W$ and our choice of $y^{\prime}$.
-Else if, $\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k-1$, then $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\subset v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$, also contradicting that the intersection is of dimension at most 00.
-Hence $\dim_{v_{2}(x_{1}^{\prime})}(R\cap v_{2}(q_{1}(F)^{\prime}))=0$. Then, in the same way as above, we have $\dim(\langle\,v_{2}(A)\,\rangle)\geq k+(km+k-1)-s$.
-(ii) In the case when $(k-1)m+k=\binom{m+d}{m}$, we have $km+k-1-s\geq m$.
+If
+$\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k,$
+then $R=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$, contradicting to the definition of $W$ and our choice of $y^{\prime}$.
+Else if
+$\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k-1,$
+then $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\subset v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$, also contradicting that the intersection is of dimension at most 00.
+Hence $\dim_{v_{2}(x_{1}^{\prime})}(R\cap v_{2}(q_{1}(F)^{\prime}))=0$. Then, in the same way as above, we have
+$\dim(\langle\,v_{2}(A)\,\rangle)\geq k+(km+k-1)-s.$
+(ii) In the case when
+$(k-1)m+k=\tbinom{m+d}{m},$
+we have $km+k-1-s\geq m$.
 It follows from Lemma 13 and ${\mathbb{P}}^{m}\simeq Z$ that $q_{i}(F)=Z$.
@@
 ∎
-Remark 20.
-For $d=3$, $m\neq 2$, if $\mu:=\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$ (as under the conditions of Proposition 15 and Remark 16 (a)), then we have $(km+k-1)-\beta_{d}\neq 1$.
-The reason is as follows.
-First, we may write $\mu={(m+3)(m+2)}/6=M/3$ for some $M\in{\mathbb{N}}$ since $(m+3)(m+2)$ is divided by $2$.
-In addition, dividing $M$ by $3$ with remainder, we have $M=3Q+R$ for a quotient $Q$ and a remainder $R$.
-Since $\mu=M/3\notin{\mathbb{N}}$, $R$ must be $1$ or $2$.
-In this setting, $k=\lceil M/3\rceil=Q+1$.
-It follows that $(km+k-1)-\beta_{d}$ is equal to
-$(km+k)-\tbinom{m+3}{3}=(m+1)\left({k-\tfrac{(m+3)(m+2)}{6}}\right)=(m+1)\left((Q+1)-(Q+\tfrac{R}{3})\right)=(m+1)\cdot\tfrac{3-R}{3}~{}.$
-If $(km+k-1)-\beta_{d}=1$, then $3=(m+1)(3-R)$. Since $3-R$ is $1$ or $2$ and $m\in{\mathbb{N}}$, we get $m=2$.
-We end this subsection by making one more important remark on the case of secant defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s, which do not satisfy the condition $\beta<km+k-1$ in Proposition 15.
+We end this subsection by making one more important remark on the case when $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta_{d}}$ is secant defective, which will be used in the proof of of Part (ii) of Theorem 3.
 Remark 21 (Estimate in defective cases).
-For four defective cases $(k,d,m)=(5,4,2)$,$(9,4,3)$,$(14,4,4)$ and $(7,3,4)$, similarly as in Proposition 15, we can have an estimation
-$\dim\langle\,v_{d-1}(A)\,\rangle\geq k+\delta~{},$
-where $A=v_{d}^{-1}(q_{1}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$, the preimage of the entry locus of $a$, and $\delta$ is the secant defect of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$; here $A$ is $\delta$-equidimensional, the $k$ general points $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ are contained in $A$, and it’s well known that $\delta=2$ when $(k,d,m)=(9,4,3)$ and $\delta=1$ in all the other defective cases.
-For 3 cases $(k,d,m)=(5,4,2),(9,4,3)$, and $(14,4,4)$, we see that $\beta_{d-1}-2m\geq k$ and $v_{d-1}(A)\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$ by the trisecant lemma so that we may take a point $y^{\prime}\in A$ such that $v_{d-1}(y^{\prime})\notin\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle$.
-By Lemma 19 (ii), we get $\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}({\mathbb{P}}^{m}))=0$ where $R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle$.
-Thus, by the same intersection argument in $\langle\,v_{d-1}(A)\,\rangle$ (as Proposition 15 (i)), we derive the estimation
+For four defective cases
+$(k,d,m)=(7,3,4),(5,4,2),(9,4,3),(14,4,4),$
+similarly as in Proposition 15, we can have an estimation
+$\dim\langle\,v_{d-1}(A)\,\rangle\geq k+\delta,$
+where $A=v_{d}^{-1}(q_{1}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$, the preimage of the entry locus of $a$, and $\delta$ is the secant defect of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$; here $A$ is $\delta$-equidimensional, the $k$ general points $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ are contained in $A$, and it is well known that $\delta=2$ when $(k,d,m)=(9,4,3)$ and $\delta=1$ in all the other defective cases.
+For three cases
+$(k,d,m)=(5,4,2),(9,4,3),(14,4,4),$
+we see that $\beta_{d-1}-2m\geq k$ and
+$v_{d-1}(A)\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$
+by the trisecant lemma so that we may take $y^{\prime}\in A$ such that $v_{d-1}(y^{\prime})\notin\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle$. By Lemma 19 (ii), we get
+$\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}({\mathbb{P}}^{m}))=0,$
+where $R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle$.
+Thus, by the intersection argument in $\langle\,v_{d-1}(A)\,\rangle$ (similar to Proposition 15 (i)), we derive the estimation
 $\dim\langle\,v_{d-1}(A)\,\rangle\geq\dim R+\dim v_{d-1}(A)=k+\delta~{}.$
-For the remaining case $(k,d,m)=(7,3,4)$, it holds $\beta_{d-1}-2m=k-1$, and we still can claim that $\dim(\langle\,v_{2}(A)\,\rangle)\geq k+\delta=7+1=8$ as follows.
-For the $6$-dimensional subspace $M:=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\,\rangle\subset\langle\,v_{2}(A)\,\rangle$, the trisecant lemma implies $M\cap v_{2}(A)\subset M\cap v_{2}({\mathbb{P}}^{4})=\left\{v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\right\}$, 00-dimensional intersection.
-Then $\dim(\langle\,v_{2}(A)\,\rangle)\geq 7$ (otherwise, we get $M=\langle\,v_{2}(A)\,\rangle$ so that $M\cap v_{2}(A)=v_{2}(A)$, a contradiction).
-Suppose that $\dim(\langle\,v_{2}(A)\,\rangle)=7$, and set $R=\langle\,v_{2}(A)\,\rangle$.
-We take the irreducible components of $1$-equidimensional closed set $A$ as $A=\bigcup_{j=1}^{s}A_{j}$.
+For the remaining case $(k,d,m)=(7,3,4)$, it holds $\beta_{d-1}-2m=k-1$, and we still can claim that
+$\dim(\langle\,v_{2}(A)\,\rangle)\geq k+\delta=7+1=8$
+as follows. For the $6$-dimensional subspace $M=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\,\rangle\subset\langle\,v_{2}(A)\,\rangle$, the trisecant lemma implies
+$M\cap v_{2}(A)\subset M\cap v_{2}({\mathbb{P}}^{4})=\left\{v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\right\},$
+00-dimensional intersection. Then $\dim(\langle\,v_{2}(A)\,\rangle)\geq 7$ (otherwise, we get $M=\langle\,v_{2}(A)\,\rangle$ so that $M\cap v_{2}(A)=v_{2}(A)$, a contradiction). Suppose that
+$\dim(\langle\,v_{2}(A)\,\rangle)=7,$
+and set $R=\langle\,v_{2}(A)\,\rangle$.
+We take the irreducible components of the $1$-equidimensional closed set $A$ as $A=\bigcup_{j=1}^{s}A_{j}$.
 Note that $v_{2}(A_{j})\subset R\cap v_{2}({\mathbb{P}}^{m})$.
@@
 2.3. Estimate for the linear span of tangents moving along a subsecant variety
-First, we give an explicit description of the embedded tangent space $\mathbb{T}_{x}v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$ at a point $x$ in $v_{d}({\mathbb{P}}^{n})$ or $v_{d}({\mathbb{P}}^{m})$ as follows (see also [14]).
+First, we give the following explicit description of the embedded tangent space $\mathbb{T}_{x}v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$ to $v_{d}({\mathbb{P}}^{n})$ at a point $x$ in $v_{d}({\mathbb{P}}^{n})$ or $v_{d}({\mathbb{P}}^{m})$.
+Note that it is related to computations of Gauss maps (see [FI17]).
 Recall that $\mathsf{mono}[t]_{\leq e}$ denotes the set of monomials $f\in{\mathbb{C}}[t_{1},\dots,t_{m}]$ with $\deg f\leq e$.
-Note that $1\in\mathsf{mono}[t]_{\leq e}$ as the monomial of degree 00.
-As mentioned in Remark 8, as changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}:=n-m$, we may assume that ${\mathbb{P}}^{m}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$.
+Then $1\in\mathsf{mono}[t]_{\leq e}$ as the monomial of degree 00.
+As mentioned in Remark 8, as changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}=n-m$, we may assume that ${\mathbb{P}}^{m}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$.
 On the affine open subset $\left\{t_{0}\neq 0\right\}$, the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized by monomials of ${\mathbb{C}}[t_{1},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}]$ of degree $\leq d$.
 So it is expressed as
-(11) $[\,\mathsf{mono}[t]_{\leq d}:u_{1}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad\cdots\quad:u_{m^{\prime}}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad*\quad\,]~{},$
-where $u_{i}\cdot\mathsf{mono}[t]_{\leq d-1}$ means $\left\{u_{i}f\mid f\in\mathsf{mono}[t]_{\leq d-1}\right\}=(u_{i}:u_{i}t_{1}:u_{i}t_{2}:\dots:u_{i}t_{m}^{d-1})$, and “*” means the remaining monomials.
+(16) $[\,\mathsf{mono}[t]_{\leq d}:u_{1}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad\cdots\quad:u_{m^{\prime}}\cdot\mathsf{mono}[t]_{\leq d-1}:\quad*\quad\,]\,,$
+where $u_{i}\cdot\mathsf{mono}[t]_{\leq d-1}$ means
+$\left\{u_{i}f\mid f\in\mathsf{mono}[t]_{\leq d-1}\right\}=(u_{i}:u_{i}t_{1}:u_{i}t_{2}:\dots:u_{i}t_{m}^{d-1}),$
+and “*” means the remaining monomials.
 Let $x=v_{d}(x^{\prime})$ with $x^{\prime}\in\left\{t_{0}\neq 0\right\}\subset{\mathbb{P}}^{n}$. Then $\mathbb{T}_{x}v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$ coincides with
 (17) $n$-plane spanned by the $(n+1)$ points corresponding to the row vectors of the matrix
-$\begin{bmatrix}v_{d}\\ \partial v_{d}/\partial t_{1}\\ \vdots\\ \partial v_{d}/\partial t_{m}\\ \partial v_{d}/\partial u_{1}\\ \vdots\\ \partial v_{d}/\partial u_{m^{\prime}}\end{bmatrix}(x^{\prime})$ =[$\mathsf{mono}[t]_{\leq d}&u_{1}\cdot\mathsf{mono}[t]_{\leq d-1}&\dots&u_{m^{\prime}}\cdot\mathsf{mono}[t]_{\leq d-1}&&*&\\ (\mathsf{mono}[t]_{\leq d})_{t_{1}}&u_{1}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{1}}&\dots&u_{m^{\prime}}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{1}}&&*&\\ \vdots&\vdots&&\vdots&&\vdots&\\ (\mathsf{mono}[t]_{\leq d})_{t_{m}}&u_{1}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{m}}&\dots&u_{m^{\prime}}\cdot(\mathsf{mono}[t]_{\leq d-1})_{t_{m}}&&*&\\ \mathsf{O}&\mathsf{mono}[t]_{\leq d-1}&\dots&\mathsf{O}&&*&\\ \vdots&&\ddots&&&\vdots&\\ \mathsf{O}&\mathsf{O}&\dots&\mathsf{mono}[t]_{\leq d-1}&&*&\end{array}\right](x^{\prime})~{}$
+[…]
 using (16), where $(\mathsf{mono}[t]_{\leq e})_{t_{i}}$ means $\left\{\partial f/\partial t_{i}\mid f\in\mathsf{mono}[t]_{\leq e}\right\}$ and $\mathsf{O}$ is a zero matrix with suitable size.
@@
 (18) [$\mathsf{mono}[t]_{\leq d}&\mathsf{O}&\dots&\mathsf{O}&&&&&\mathsf{O}&&&\\ (\mathsf{mono}[t]_{\leq d})_{t_{1}}&\mathsf{O}&\dots&\mathsf{O}&&&&&\mathsf{O}&&&\\ \vdots&\vdots&&\vdots&&&&&\vdots&&&\\ (\mathsf{mono}[t]_{\leq d})_{t_{m}}&\mathsf{O}&\dots&\mathsf{O}&&&&&\mathsf{O}&&&\\ \mathsf{O}&\mathsf{mono}[t]_{\leq d-1}&\dots&\mathsf{O}&&&&&\mathsf{O}&&&\\ \vdots&&\ddots&&&&&&\vdots&&&\\ \mathsf{O}&\mathsf{O}&\dots&\mathsf{mono}[t]_{\leq d-1}&&&&&\mathsf{O}&&&\end{array}\right](x^{\prime})~{}.$
-As a consequence, we settle a key proposition which estimates a lower bound of dimension of linear span of moving embedded tangent spaces along a subset of a given ${\mathbb{P}}^{m}$.
+As a consequence, we settle a key proposition which estimates a lower bound of the dimension of the linear span of moving embedded tangent spaces along a subset of a given ${\mathbb{P}}^{m}$.
 Proposition 22.
@@
 where $v_{e,m}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\binom{m+e}{m}-1}$ is the $e$-uple Veronese embedding of ${\mathbb{P}}^{m}$.
-Proof of Proposition 22.
+Proof.
 For a given $A\subset{\mathbb{P}}^{m}$, we consider
-$B_{0}=v_{d}(A),B_{1}=(\partial v_{d}/\partial u_{1})(A),~{}\dots~{},B_{m^{\prime}}=(\partial v_{d}/\partial u_{m^{\prime}})(A)$
+$B_{0}=v_{d}(A),B_{1}=(\partial v_{d}/\partial u_{1})(A),\,\dots\,,B_{m^{\prime}}=(\partial v_{d}/\partial u_{m^{\prime}})(A)$
 as subsets in ${\mathbb{P}}^{N}$, where $B_{i}$ is embedded by the parameterization of $(m+1+i)$-th row of the matrix of (17) for $1\leq i\leq m^{\prime}$.
@@
 Since $A\subset\left\{u_{1}=\dots=u_{m^{\prime}}=0\right\}$, it follows from (18) that the linear variety
-(15) $\langle\,\Lambda\cup B_{0},B_{1},~{}\dots~{},B_{m^{\prime}}\,\rangle\subset{\mathbb{P}}^{N}$
-is of dimension $\dim(\langle\,\Lambda\cup B_{0}\,\rangle)+\dim(\langle\,B_{1}\,\rangle)+\dots\dim(\langle\,B_{m^{\prime}}\,\rangle)+m^{\prime}$.
-By (13) again, we see that $B_{0}\simeq v_{d}(A)$ and $B_{i}\simeq v_{d-1,m}(A)$ for $1\leq i\leq m^{\prime}$.
+(20) $\langle\,\Lambda\cup B_{0},B_{1},\,\dots\,,B_{m^{\prime}}\,\rangle\subset{\mathbb{P}}^{N}$
+is of dimension $\dim(\langle\,\Lambda\cup B_{0}\,\rangle)+\dim(\langle\,B_{1}\,\rangle)+\dots+\dim(\langle\,B_{m^{\prime}}\,\rangle)+m^{\prime}$.
+Again, by (18) we see that $B_{0}\simeq v_{d}(A)$ and $B_{i}\simeq v_{d-1,m}(A)$ for $1\leq i\leq m^{\prime}$.
 Since the linear variety (20) is contained in $\langle\,\Lambda\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle$, we have the assertion.
@@
 3.1. Symmetric flattening and conormal space computation
-For the proof of Theorem 1, we begin with some preliminaries on equations for secant varieties and conormal space computation via known sets of equations, whereas we mainly adopt geometric viewpoint and techniques for $m\geq 2$ case in §4.
+For the proof of Theorem 1, we begin with some preliminaries on equations for secant varieties and conormal space computation via known sets of equations, whereas we mainly adopt the geometric viewpoint and techniques for the $m\geq 2$ case in §4.
 Let $V$ be a $(n+1)$-dimensional ${\mathbb{C}}$-vector space ${\mathbb{C}}\langle x_{0},x_{1},\dots,x_{n}\rangle$.
-For the Veronese variety $v_{d}({\mathbb{P}}V)$, we have a natural 1-1 correspondence between a point of the ambient space $\langle\,v_{d}({\mathbb{P}}V)\,\rangle$ and a homogeneous polynomial $f$ of degree $d$ (or a form) in $S:={\mathbb{C}}[x_{0},x_{1},\dots,x_{n}]$.
-First of all, let us recall some related notions to this correspondence.
-Given a form $f$ of degree $d$, the minimum number of linear forms $l_{i}$ needed to write $f$ as a sum of $d$-th powers is the so-called (Waring) rank of $f$ and let us denote it by $\operatorname{rank}(f)$.
+Let $f\in S^{d}V$ be a homogeneous polynomial of degree $d$ (or $d$-form) and $[f]$ be the corresponding point in ${\mathbb{P}}S^{d}V$.
+In this paper, we frequently abuse notations as denoting both a $d$-form in $S^{d}V$ and the point in ${\mathbb{P}}S^{d}V$ just by $f$.
+For the Veronese variety $v_{d}({\mathbb{P}}V)$, we have a natural 1-1 correspondence between points of the ambient space $\langle\,v_{d}({\mathbb{P}}V)\,\rangle$ and equivalent classes of degree $d$-forms in $S={\mathbb{C}}[x_{0},x_{1},\dots,x_{n}]$.
+First of all, let us recall some notions related to this correspondence.
+Given a form $f$ of degree $d$, the minimum number of linear forms $l_{i}$ needed to write $f$ as a sum of $d$-th powers is the so-called (Waring) rank of $f$ and we denote it by $\operatorname{rank}(f)$.
+Note that one can define $\operatorname{rank}([f])$ by $\operatorname{rank}(f)$, because this rank is invariant under non-zero scaling.
 The (Waring) border rank is given by the same notion in the limiting sense.
 In other words, if there is a family $\{f_{\epsilon}\mid\epsilon>0\}$ of polynomials with constant rank $r$ and $\lim_{\epsilon\to 0}f_{\epsilon}=f$, then we say that $f$ has border rank at most $r$.
-The minimum such $r$ is called the border rank of $f$ and let’s denote it again by $\underline{\mathrm{rank}}(f)$.
+The minimum such $r$ is called the border rank of $f$ and we denote it again by $\underline{\mathrm{rank}}(f)$.
 Note that by definition $\sigma_{k}(v_{d}({\mathbb{P}}V))$ is the variety of homogeneous polynomials $f$ of degree $d$ with border rank $\underline{\mathrm{rank}}(f)\leq k$.
@@
 Define the differential action of $T$ on $S$ as follows: for any $g\in T_{d-a},f\in S_{d}$, we set
-$g\cdot f=g(\partial_{0},\partial_{1},\ldots,\partial_{n})f\in S_{a}~{},$
+$g\cdot f=g(\partial_{0},\partial_{1},\ldots,\partial_{n})f\in S_{a},$
 where $\partial_{i}=\partial/\partial{x_{i}}$ Let us take bases for $S_{a}$ and $T_{d-a}$ as
-$\mathbf{X}^{I}=\frac{1}{i_{0}!\cdots i_{n}!}x_{0}^{i_{0}}\cdots x_{n}^{i_{n}}$ $\quad\textrm{and}\quad\mathbf{Y}^{J}=y_{0}^{j_{0}}\cdots y_{n}^{j_{n}}~{},$
+$\mathbf{X}^{I}=\frac{1}{i_{0}!\cdots i_{n}!}x_{0}^{i_{0}}\cdots x_{n}^{i_{n}}$ $\quad\textrm{and}\quad\mathbf{Y}^{J}=y_{0}^{j_{0}}\cdots y_{n}^{j_{n}},$
 with $|I|=i_{0}+\cdots+i_{n}=a$ and $|J|=j_{0}+\cdots+j_{n}=d-a$. For a given $f=\sum_{|I|=d}b_{I}\cdot\mathbf{X}^{I}$ in $S_{d}$, we have a linear map
@@
 for any $a$ with $11\leq a\leq d-$ 1, which can be represented by the following (${a+n\choose n}\times{d-a+n\choose n}$-matrix:
-($\ &b_{I,J}&\\ &&\end{array}\right)\quad\textrm{with $b_{I,J}=b_{I+J}$}~{},$
-in the bases defined above. We call this ‘$(d-a,a)$-symmetric flattening (or catalecticant) matrix’ of $f$. It is easy to see that the transpose $\phi_{d-a,a}(f)^{T}$ is equal to $\phi_{a,d-a}(f)$.
+($\ &b_{I,J}&\\ &&\end{array}\right)\quad\textrm{with $b_{I,J}=b_{I+J}$},$
+in the bases defined above.
+We call this ‘the $(d-a,a)$-symmetric flattening (or catalecticant) matrix’ of $f$.
+It is easy to see that the transpose $\phi_{d-a,a}(f)^{T}$ is equal to $\phi_{a,d-a}(f)$.
 It is obvious that if $f$ has (border) rank 1, then any symmetric flattening $\phi_{d-a,a}(f)$ has rank 1.
@@
 So, we obtain a set of defining equations coming from $(k+1)$-minors of the matrix $\phi_{d-a,a}(f)$ for $\sigma_{k}(v_{d}({\mathbb{P}}V))$.
-For some small values of $k$, it is known that these minors are sufficient to cut the variety $\sigma_{k}(v_{d}({\mathbb{P}}V))$ scheme-theoretically (see [23, theorem 3.2.1 (1)]).
-Let us recall another basic terms and facts.
+For some small values of $k$, it is known that these minors are sufficient to cut the variety $\sigma_{k}(v_{d}({\mathbb{P}}V))$ scheme-theoretically (see [LO13, Theorem 3.2.1]).
+Let us recall some more basic terms and facts.
 Let $Z\subset{\mathbb{P}}W$ be a (reduced and irreducible) variety and $\hat{Z}$ be its affine cone in $W$.
-Consider a (closed) point $p\in\hat{Z}$ and say $[p]$ the corresponding point in ${\mathbb{P}}W$.
-We denote the affine tangent space to $Z$ at $[p]$ in $W$ by $\hat{T}_{[p]}Z$ and we define the (affine) conormal space to $Z$ at $[p]$, $\hat{N}^{\ast}_{[p]}Z$ as the annihilator $(\hat{T}_{[p]}Z)^{\perp}\subset W^{\ast}$.
-Since $\dim\hat{N}^{\ast}_{[p]}Z+\dim\hat{T}_{[p]}Z=\dim W$ and $\dim Z\leq\dim\hat{T}_{[p]}Z-1$, we get that $\dim\hat{N}^{\ast}_{[p]}Z\leq\operatorname{codim}(Z,{\mathbb{P}}W)$ and the equality holds if and only if $Z$ is smooth at $[p]$.
-This conormal space is quite useful to study the (embedded) tangent space $\mathbb{T}_{[p]}Z$.
-For any given form $f\in S^{d}V$, we call $\partial\in T_{t}$ apolar to $f$ if the differentiation $\partial\cdot f$ gives zero (i.e.
-$\partial\in\ker\phi_{t,d-t}(f)$).
+Consider a (closed) point $\hat{p}\in\hat{Z}$ and say $p$ the corresponding point in ${\mathbb{P}}W$.
+We denote the affine tangent space to $Z$ at $p$ in $W$ by $\hat{T}_{p}Z$ and we define the (affine) conormal space to $Z$ at $p$, $\hat{N}^{\ast}_{p}Z$ as the annihilator $(\hat{T}_{p}Z)^{\perp}\subset W^{\ast}$.
+Since $\dim\hat{N}^{\ast}_{p}Z+\dim\hat{T}_{p}Z=\dim W$ and $\dim Z\leq\dim\hat{T}_{p}Z-1$, we get that
+(21) $\dim\hat{N}^{\ast}_{p}Z\leq\operatorname{codim}(Z,{\mathbb{P}}W)$
+and the equality holds if and only if $Z$ is smooth at $p$. This conormal space is quite useful to study the (embedded) tangent space $\mathbb{T}_{p}Z$.
+For any given form $f\in S^{d}V$, we call $\partial\in T_{t}$ apolar to $f$ if the differentiation $\partial\cdot f$ gives zero (i.e., $\partial\in\ker\phi_{t,d-t}(f)$).
 And we define the apolar ideal $f^{\perp}\subset T$ as
-$f^{\perp}:=\{\partial\in T~{}|~{}\partial\cdot f=0\}~{}.$
+$f^{\perp}=\{\partial\in T~{}|~{}\partial\cdot f=0\}.$
 It is straightforward to see that $f^{\perp}$ is indeed an ideal of $T$.
-Moreover, it is well-known that the quotient ring $T_{f}:=T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ (see e.g.
-[19, chapter 1]).
+Moreover, it is well-known that the quotient ring $T_{f}=T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ (see e.g.
+[IK99, Chapter 1]).
 In terms of this apolar ideal, we have a useful description of (a part of) conormal space as follows:
 Proposition 23.
-Suppose that $f\in S^{d}V$ corresponds to a (closed) point of $\sigma_{k}(v_{d}({\mathbb{P}}V))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}V))$.
+Suppose that $f\in S^{d}V$ corresponds to a (closed) point $[f]$ of $\sigma_{k}(v_{d}({\mathbb{P}}V))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}V))$.
 Then, for any $a$ with $1\leq a\leq{\lfloor\frac{d+1}{2}\rfloor}$ with $\operatorname{rank}\phi_{d-a,a}(f)=k$ we have
@@
 Let $X\subset{\mathbb{P}}W$ be any variety.
-For any linear embedding $W\hookrightarrow A\otimes B$ and the induced embedding $X\subset{\mathbb{P}}W\hookrightarrow{\mathbb{P}}(A\otimes B)$, it is well-known that for any $f\in\sigma_{k}(X)\subset{\mathbb{P}}(A\otimes B)$, considering $\sigma_{k}(X)$ as a subvariety of ${\mathbb{P}}(A\otimes B)$, we have
-$\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\supseteq\ker(f)\otimes\operatorname{im}(f)^{\perp}=\hat{N}^{\ast}_{[f]}\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))\quad\textrm{in $A^{\ast}\otimes B^{\ast}$}$
-provided that $X\subseteq\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$, $X\nsubseteq\sigma_{p-1}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$ and $f$ has rank $k\cdot p$ as a linear map in $\operatorname{Hom}(A^{\ast},B)$ (see e.g.
-[23, section 2.5]).
+For any linear embedding $W\hookrightarrow A\otimes B$ and the induced embedding $X\subset{\mathbb{P}}W\hookrightarrow{\mathbb{P}}(A\otimes B)$, it is well-known that for any $[f]\in\sigma_{k}(X)\subset{\mathbb{P}}(A\otimes B)$, considering $\sigma_{k}(X)$ as a subvariety of ${\mathbb{P}}(A\otimes B)$, we have
+$\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\supseteq\ker(f)\otimes\operatorname{im}(f)^{\perp}=\hat{N}^{\ast}_{[f]}\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$
+in $A^{\ast}\otimes B^{\ast}$ provided that $X\subseteq\sigma_{p}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$, $X\nsubseteq\sigma_{p-1}(Seg({\mathbb{P}}A\times{\mathbb{P}}B))$ and $f$ has rank $k\cdot p$ as a linear map in $\operatorname{Hom}(A^{\ast},B)$ (see e.g.
+[LO13, section 2.5]).
 Here, $Seg({\mathbb{P}}A\times{\mathbb{P}}B)$ means the Segre variety in ${\mathbb{P}}(A\otimes B)$.
@@
 where $\pi:A^{\ast}\otimes B^{\ast}\to W^{\ast}$ is the dual map of the given inclusion $W\hookrightarrow A\otimes B$.
-The assertion is immediate when we apply this fact to a partial polarization $S^{d}V\hookrightarrow S^{a}V\otimes S^{d-a}V$, because $X=v_{d}({\mathbb{P}}V)$ is contained in $Seg({\mathbb{P}}S^{a}V\times{\mathbb{P}}S^{d-a}V)$ in ${\mathbb{P}}(S^{a}V\otimes S^{d-a}V)$ (i.e.
-$p=1$ case) and $\operatorname{rank}\phi_{d-a,a}(f)=k$, $\ker\phi_{d-a,a}(f)=(f^{\perp})_{d-a}$, $\operatorname{im}(\phi_{d-a,a}(f))^{\perp}=(f^{\perp})_{a}$.
+The assertion is immediate when we apply this fact to a partial polarization $S^{d}V\hookrightarrow S^{a}V\otimes S^{d-a}V$, because $X=v_{d}({\mathbb{P}}V)$ is contained in $Seg({\mathbb{P}}S^{a}V\times{\mathbb{P}}S^{d-a}V)\subset{\mathbb{P}}(S^{a}V\otimes S^{d-a}V)$ (i.e., $p=1$ case) and $\operatorname{rank}\phi_{d-a,a}(f)=k$, $\ker\phi_{d-a,a}(f)=(f^{\perp})_{d-a}$, $\operatorname{im}(\phi_{d-a,a}(f))^{\perp}=(f^{\perp})_{a}$.
 ∎
@@
 Consider $f$ as a polynomial in ${\mathbb{C}}[x_{0},x_{1}]$ as in §3.1.
-Then, by [19, theorem 1.44] we know that $T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ and that $f^{\perp}$ is a complete intersection of two homogeneous polynomials $F,G$ of each degree $a$ and $b$ $(a\leq b)$ with $a+b=d+2$ as an ideal of ${\mathbb{C}}[y_{0},y_{1}]$, where the Hilbert function of $T/f^{\perp}$ is $(1,2,\cdots,a-1,a,\cdots,a,a-1,\cdots,2,1)$.
-We claim that $\operatorname{rank}~{}\phi_{k,d-k}(f)=k$ (i.e.
-$a=k$).
-If $a<k$, then $\operatorname{rank}~{}\phi_{t,d-t}(f)<k$ for all $t$.
-Since it is well-known that the $k$-minors of catalecticant $\phi_{t,d-t}$ for any $k-1\leq t\leq d-(k-1)$ gives the ideal of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))$, we get $f\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ which is contradiction.
+Then, by [IK99, Theorem 1.44] we know that $T/f^{\perp}$ is an Artinian Gorenstein algebra with socle degree $d$ and that $f^{\perp}$ is a complete intersection of two homogeneous polynomials $F,G$ of each degree $a$ and $b$ $(a\leq b)$ with $a+b=d+2$ as an ideal of ${\mathbb{C}}[y_{0},y_{1}]$, where the Hilbert function of $T/f^{\perp}$ is
+(22) $(1,2,\ldots,a-1,a,\ldots,a,a-1,\ldots,2,1).$
+We claim that $\operatorname{rank}~{}\phi_{k,d-k}(f)=k$ (i.e., $a=k$).
+If $a<k$, then by the shape (22) we see that $\operatorname{rank}~{}\phi_{t,d-t}(f)<k$ for all $t$.
+In particular, all $k$-minors of $\phi_{t,d-t}(f)$ vanish for any $t$.
+Since the $k$-minors of catalecticant $\phi_{t,d-t}$ for each $k-1\leq t\leq d-(k-1)$ give the ideal of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))$ (e.g.
+[Ibid., Theorem 1.45]), this implies $f\in\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, which is a contradiction.
 Hence, we have that $f^{\perp}=\big{(}F,G,y_{2},\ldots,y_{n}\big{)}$ as an ideal in $T={\mathbb{C}}[y_{0},y_{1},\ldots,y_{n}]$ for some polynomial $F$ of degree $k$ and $G$ of degree $(d-k+2)$ in ${\mathbb{C}}[y_{0},y_{1}]$.
-Now, let us show that $\sigma_{k}(X)$ is smooth at $f$ by computing the dimension of conormal space. In general, we have
-(16) (${n+d\choose d}-kn-k\geq\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X)~{},$
-where the left hand side is given by the expected codimension of the conormal space. By Proposition 23, we also have
-(17) $\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\geq\dim_{\mathbb{C}}(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}~{}.$
+Now, let us show that $\sigma_{k}(X)$ is smooth at $f$ by computing the dimension of conormal space. In general, by (21) we have
+(23) (${n+d\choose d}-kn-k\geq\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X),$
+where the left hand side is given by the expected codimension of the $k$-th secant variety. By Proposition 23, we also have
+(24) $\dim_{\mathbb{C}}\hat{N}^{\ast}_{[f]}\sigma_{k}(X)\geq\dim_{\mathbb{C}}(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}\,.$
 Thus, $f$ is a smooth point of $\sigma_{k}(X)$ if the lower bound for the dimension of conormal space in (24) is equal to the expected codimension in (23).
-Since $k\leq\frac{d+1}{2}$ by the assumption, note that $d-k\geq k$ unless $d$ is odd and $k=\frac{d+1}{2}$, in which $d-k=\frac{d-1}{2}<k$.
-a) $d$: odd and $k=\frac{d+1}{2}$ case : In this case, we have
+Since $k\leq\frac{d+1}{2}$ by the assumption, note that $d-k\geq k$ unless $d$ is odd and $k=\frac{d+1}{2}$, where $d-k=\frac{d-1}{2}<k$.
+a) If $d$ is odd and $k=\frac{d+1}{2}$, then we have
 $(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}$ $=\big{(}F,y_{2},\ldots,y_{n}\big{)}_{k}\cdot\big{(}y_{2},\ldots,y_{n}\big{)}_{d-k}$
@@
 =($={n+d\choose d}-(d+1)-d(n-1)+(n-1)(d-k)\quad(\textrm{note that $k=\frac{d+1}{2}$})$
-=($={n+d\choose d}-kn-k~{},$
+=($={n+d\choose d}-kn-k,$
 which tells us that $\sigma_{k}(X)$ is smooth at $f$.
-b) $d$: even or $d$: odd and $k<\frac{d+1}{2}$ case : In the second case, we have $k\leq d-k$ and
+b) When $d$ is odd and $k<\frac{d+1}{2}$ or $d$ is even, we have $k\leq d-k$ and
 $(f^{\perp})_{k}\cdot(f^{\perp})_{d-k}$ $=\big{(}F,y_{2},\ldots,y_{n}\big{)}_{k}\cdot\big{(}F,y_{2},\ldots,y_{n}\big{)}_{d-k}$
@@
 $\dim\hat{N}^{\ast}_{[f]}\sigma_{k}(X)$ ($\geq{n+d\choose d}-(d+1)-d(n-1)+(n-1)(d-k)+(d-2k+1)$
-=($={n+d\choose d}-kn-k~{},$
+=($={n+d\choose d}-kn-k,$
 which coincides with the expected codimension as desired. Thus, $f$ is a smooth point of $\sigma_{k}(X)$.
 (ii) First note that $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-1,d\right\}$ and the incidence $I$ has dimension $12k-$ 1.
-In the case $d\leq 2k-2$, each fiber of $p:I\rightarrow{\mathbb{P}}^{d}$ is of dimension $\geq 1$, and then $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$ with $1\leq i\leq k$ in the incidence (5) in §2.2.
+In the case $d\leq 2k-2$, each fiber of $p:I\rightarrow{\mathbb{P}}^{d}$ is of dimension $\geq 1$, so for a general $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$, it holds $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$ with $1\leq i\leq k$ in the incidence (9) in §2.2.
 Now, let $n\geq 3,k=3$ or $n\geq 2,k\geq 4$ and $d=2k-2$.
@@
 Then a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{1}))={\mathbb{P}}^{d}$ is a smooth point of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
-Since $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$, it follows from Lemma 12 that
-(18) $\langle\,\bigcup_{x\in v_{d}({\mathbb{P}}^{1})}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
-Taking $m=1$ and $A={\mathbb{P}}^{1}$ in Proposition 22, the number (14), a lower bound for dimension of left hand side of (18), is equal to $dn$. Thus we have
-$(2k-2)n=dn\leq k(n+1)-1,$
+Since $q_{i}(p^{-1}(a))=v_{d}({\mathbb{P}}^{1})$ for some $i$, it follows from Lemma 12 that for $M=\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, we have the inclusion $\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\subset M$ for a general $x\in v_{d}({\mathbb{P}}^{1})$, and then the inclusion holds for any $x\in v_{d}({\mathbb{P}}^{1})$.
+This is because, for the Gauss map $\gamma:v_{d}({\mathbb{P}}^{n})\rightarrow{\mathbb{G}}(n,{\mathbb{P}}^{N})$ sending $\gamma(z)=\mathbb{T}_{z}(v_{d}({\mathbb{P}}^{n}))$ (it is a morphism since $v_{d}({\mathbb{P}}^{n})$ is smooth), considering the closed set $G_{M}=\left\{W\in{\mathbb{G}}(n,{\mathbb{P}}^{N})\mid W\subset M\right\}$, we have $\gamma(U)\subset G_{M}$ for a certain non-empty open subset $U\subset v_{d}({\mathbb{P}}^{1})$, and then $\gamma(v_{d}({\mathbb{P}}^{1}))\subset G_{M}$.
+Therefore,
+(25) $\langle\,\bigcup_{x\in v_{d}({\mathbb{P}}^{1})}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\,\rangle\subset\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{n})).$
+Taking $m=1$, $\Lambda=\emptyset$, and $A={\mathbb{P}}^{1}$ in Proposition 22, the number (19), a lower bound for dimension of left hand side of (25), is equal to $dn$. Thus we have
+$(2k-2)n=dn\leq k(n+1)-1\;\big{(}\!=\dim\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\big{)},$
 which is equivalent to the formula $n\leq(k-1)/(k-2)$. It follows that $n\leq 2$ if $k=3$, and $n=1$ if $k\geq 4$, contrary to our assumption.
-Finally, since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subsetneq\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ when $d\geq 2k-2$ (note that $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=2k-3<d$), the $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ is a non-trivial singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which means $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, by Lemma 10.
-(iii) By assumption, $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-3,d\right\}=d$, that is to say, $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle={\mathbb{P}}^{d}$; hence the assertion follows.
-(iv) For $(n,k)=(2,3)$, smoothness of all points in $\sigma_{3}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{2}(v_{d}({\mathbb{P}}^{2}))$ for $d\geq 4$ was already proved in [16, theorem 2.14].
-This is included for completeness of the statement.
+Finally, since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))\subsetneq\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ when $d\geq 2k-2$ (note that $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=2k-3<d$), the $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ is a non-trivial singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which means that $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, by Lemma 10.
+(iii) By assumption, $\dim\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\min\left\{2k-3,d\right\}=d$, that is to say,
+$\sigma_{k-1}(v_{d}({\mathbb{P}}^{1}))=\sigma_{k}(v_{d}({\mathbb{P}}^{1}))=\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle={\mathbb{P}}^{d};$
+hence the assertion follows.
+(iv) For $(n,k)=(2,3)$, smoothness of all points in $\sigma_{3}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{2}(v_{d}({\mathbb{P}}^{2}))$ for $d\geq 4$ was already proved in [Han18, Theorem 2.14].
+This is included for completeness to the statement.
 ∎
@@
 (iv) is the exception to the trichotomy in Theorem 1.
-Under the condition $(k,d,m,n)=(3,4,1,2)$ of (iv), the arithmetic deduced from the containment assumption (18) of moving tangents in the proof does not make any contradiction.
+Under the condition $(k,d,m,n)=(3,4,1,2)$ of (iv), the arithmetic deduced from the inclusion assumption (25) of moving tangents in the proof does not make any contradiction.
 The situation is similar in the other exceptional case to the trichotomy, $(k,d,m,n)=(4,3,2,3)$ (Part (iv) of Theorem 2).
 4. Proof of main results
-In this section, we prove Theorem 2 and Theorem 3. We will discuss the non-singularity result first and results for the singular loci next.
+In this section, we prove Theorem 2 and Theorem 3. We will first discuss the non-singularity result and then the results for the singular loci.
 4.1. Generic smoothness
@@
 Lemma 25.
-Assume $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$, $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={mk+k-1}$, and $k\leq\tbinom{m+d-1}{m}$.
+Assume $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$ and $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={mk+k-1}$ (i.e., $v_{d}({\mathbb{P}}^{n})$ and $v_{d}({\mathbb{P}}^{m})$ are non-defective).
+Let $k\leq\tbinom{m+d-1}{m}$.
 Fix $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ to be an $m$-plane, and take $a\in\sigma_{k}(v_{d}(L))$ to be a general point.
@@
 Proof.
-Consider any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in p^{-1}(a)\subset I$.
-Let $I_{(m)}\subset I\subset{\mathbb{P}}^{N}\times(L)^{k}$ be the another incidence as in Lemma 9.
+(i) Consider any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in p^{-1}(a)\subset I$.
+Let $I_{(m)}\subset{\mathbb{P}}^{N}\times(L)^{k}$ be the another incidence as in Lemma 9.
 Since $a\in\sigma_{k}(v_{d}(L))$ is general, it follows $a\notin\sigma_{k-1}(v_{d}(L))$ and $a\notin p(I_{(m)}\setminus I_{(m)}^{0})$ by Remark 7 (b).
@@
 From Lemma 10, it is also true that $a\notin\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
-We may write $a=\sum_{i=1}^{k}c_{i}x_{i}$ for some $c_{i}\in{\mathbb{C}}$ as regarding $a$ and $x_{i}=v_{d}(x_{i}^{\prime})$ as vectors in affine space ${\mathbb{C}}^{N+1}$.
-As in Remark 8 (a), we set $y_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:0:\dots:0]$.
-Then $a=\sum_{i=1}^{k}c_{i}y_{i}$, where it holds that $c_{i}\neq 0$ and $y_{i}^{\prime}\neq 0$ for all $1\leq i\leq k$.
-We may assume $x_{i}^{\prime}\in\left\{t_{0}\neq 0\right\}$ for all $i$.
-Since $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized by $\mathsf{mono}[t,u]_{\leq d}$, and $a\in\langle\,v_{d}(L)\,\rangle$, it holds that
-$0=\sum_{i=1}^{k}c_{i}\cdot\{(u_{1}\cdot\mathsf{mono}[t]_{\leq d-1})(x_{i}^{\prime})\}=\sum_{i=1}^{k}c_{i}\cdot x_{i,m+1}^{\prime}\cdot\{\mathsf{mono}[t]_{\leq d-1}(y_{i}^{\prime})\}~{},$
-where $\{Mono(pt)\}$ means the vector obtained by evaluating monomials in $Mono$ at the value of $pt$.
-Applying Remark 7 (b) to $\sigma_{k}(v_{d}(L))$, we have
-$\dim(\langle\,v_{d-1}(y_{1}^{\prime}),\dots,v_{d-1}(y_{k}^{\prime})\,\rangle)=k-1.$
-It means that $c_{i}\cdot x_{i,m+1}^{\prime}=0$, thus $x_{i,m+1}^{\prime}=0$ for all $1\leq i\leq k$.
-Similarly, we can obtain $x_{i,j}^{\prime}=0$ for all $1\leq i\leq k$ and $j>m$, which give the linear defining equations for $(L)^{k}$ in $({\mathbb{P}}^{n})^{k}$.
-Hence, scheme-theoretically, it follows that $x_{1}^{\prime},\dots,x_{k}^{\prime}\in L$.
+Thus, we may write $a=\sum_{i=1}^{k}c_{i}x_{i}$ for some $c_{i}\in{\mathbb{C}}$ as regarding $a$ and $x_{i}=v_{d}(x_{i}^{\prime})$ as vectors in the affine space ${\mathbb{C}}^{N+1}$, where $c_{i}\neq 0$ for all $1\leq i\leq k$.
+As in Remark 8, let us set $y_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:0:\dots:0]$.
+For $y_{i}=v_{d}(y_{i}^{\prime})$, the diagram (5) implies $a=\sum_{i=1}^{k}c_{i}y_{i}$, where $y_{i}^{\prime}\neq 0$ for $1\leq i\leq k$.
+For the affine open subset $V_{0}=\left\{t_{0}\neq 0\right\}\subset{\mathbb{P}}^{n}$, we may assume $x_{i}^{\prime}\in V_{0}$ for all $i$.
+Since $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ is parameterized on $V_{0}$ by $\mathsf{mono}[t,u]_{\leq d}$, and $a\in\langle\,v_{d}(L)\,\rangle$, it holds that
+$0=\sum_{i=1}^{k}c_{i}\cdot\{(u_{1}\cdot\mathsf{mono}[t]_{\leq d-1})(x_{i}^{\prime})\}=\sum_{i=1}^{k}c_{i}\cdot x_{i,m+1}^{\prime}\cdot\{\mathsf{mono}[t]_{\leq d-1}(y_{i}^{\prime})\},$
+where for $Mono=u_{1}\cdot\mathsf{mono}[t]_{\leq d-1},\mathsf{mono}[t]_{\leq d-1}$ and $pt=x_{i}^{\prime},y_{i}^{\prime}$, the symbol $\{Mono(pt)\}$ means the vector obtained by evaluating monomials in $Mono$ at the value of $pt$.
+Since $k\leq\tbinom{m+d-1}{m}$, applying Remark 7 (b) to $\sigma_{k}(v_{d}(L))$, we may assume
+(26) $\dim\langle\,v_{d-1}(y_{1}^{\prime}),\dots,v_{d-1}(y_{k}^{\prime})\,\rangle=k-1,$
+which gives $c_{i}\cdot x_{i,m+1}^{\prime}=0$, thus $x_{i,m+1}^{\prime}=0$ for all $1\leq i\leq k$ (more precisely, the linear independence of (26) means a $k$-minor of the corresponding matrix is non-zero, and $c_{i}\cdot x_{i,m+1}^{\prime}=0$ is obtained by multiplying the inverse of the $k\times k$ submatrix).
+Similarly, we can obtain $x_{i,j}^{\prime}=0$ for each $j>m$ and for all $1\leq i\leq k$, which gives the linear defining equations for $(L)^{k}$ in $({\mathbb{P}}^{n})^{k}$.
+Hence, $x_{1}^{\prime},\dots,x_{k}^{\prime}\in L$.
+(ii) Let $U\subset({\mathbb{P}}^{n})^{k}$ be the open subset used in Remark 7, where $I^{0}$ is the ${\mathbb{P}}^{k-1}$-bundle over $U$.
+We define a morphism $\Phi:{\mathbb{P}}^{k-1}\times U\rightarrow{\mathbb{P}}^{N}\times U$ by
+$\Phi((c_{1}:\dots:c_{k}),(x_{1}^{\prime},\dots,x_{k}^{\prime}))=\Bigl{(}\,\sum_{i=1}^{k}c_{i}v_{d}(x_{i}^{\prime}),\,(x_{1}^{\prime},\dots,x_{k}^{\prime})\Bigr{)}.$
+Note that, by the linear independence of $v_{d}(x_{1}^{\prime}),\dots,v_{d}(x_{k}^{\prime})$ for $(x_{1}^{\prime},\dots,x_{k}^{\prime})\in U$, $\sum_{i=1}^{k}c_{i}v_{d}(x_{i}^{\prime})=\sum_{i=1}^{k}\tilde{c}_{i}v_{d}(x_{i}^{\prime})\in{\mathbb{P}}^{N}$ if and only if $(c_{1}:\dots:c_{k})=(\tilde{c}_{1}:\dots:\tilde{c}_{k})\in{\mathbb{P}}^{k-1}$.
+Then $\Phi({\mathbb{P}}^{k-1}\times U)=I^{0}$, and moreover, we have the isomorphism ${\mathbb{P}}^{k-1}\times U\simeq I^{0}$ under $\Phi$.
+Let $U_{0}=U\cap(V_{0})^{k}\subset({\mathbb{P}}^{n})^{k}$, where $(V_{0})^{k}$ is an affine variety and its affine coordinates ring is $A={\mathbb{C}}[\{x_{i,j}^{\prime}\}]$.
+In addition, for each $k$-minor $\xi$ of the matrix whose $i$-th column consists of monomials of $m$ variables $x_{i,1}^{\prime},\dots,x_{i,m}^{\prime}$ of degrees $\leq d-1$, we set $(V_{0})^{k}_{\xi}=\left\{\xi\neq 0\right\}$, an open subset of $(V_{0})^{k}$ whose coordinates ring is $A_{\xi}$.
+Let $W\subset\left\{c_{1}\neq 0\right\}\subset{\mathbb{P}}^{k-1}$ be the affine open subset such that all the coordinates $c_{1},\dots,c_{k}$ are nonzero, where the coordinates ring of $W$ is ${\mathbb{C}}[c_{2},\dots,c_{k}]_{c_{2}\dotsm c_{k}}$ by regarding $c_{1}=1$.
+We may assume $p^{-1}(a)\subset I^{0}\cap(\left\{a\right\}\times(V_{0})^{k})$.
+To consider the scheme-theoretic structure of $p^{-1}(a)$, for the composite morphism $\Phi_{1}=p\circ\Phi:{\mathbb{P}}^{k-1}\times U_{0}\rightarrow{\mathbb{P}}^{N}$, we take the fiber
+$\Phi_{1}^{-1}(a)\subset{\mathbb{P}}^{k-1}\times U_{0}\subset{\mathbb{P}}^{k-1}\times(V_{0})^{k}.$
+Since $a\notin\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, $\Phi_{1}^{-1}(a)\subset W\times(V_{0})^{k}$.
+Since $a$ is general in $\sigma_{k}(v_{d}(L))$, and by (26), $\Phi_{1}^{-1}(a)$ is contained in the union of affine open subsets $W\times(V_{0})^{k}_{\xi}$ with all $k$-minors $\xi$.
+We take $F_{\xi}=\Phi_{1}^{-1}(a)\cap(W\times(V_{0})^{k}_{\xi})$ for each $\xi$, and consider the ideal $I(F_{\xi})$ in $A_{\xi}[c_{2},\dots,c_{k}]_{c_{2}\dotsm c_{k}}$, the affine coordinates ring of $W\times(V_{0})^{k}_{\xi}$.
+For $\beta=\binom{m+d}{m}-1$, we may write
+$a=(1:a^{(1)}:\dots:a^{(\beta)}:0:\dots:0)\in\langle\,v_{d}(L)\,\rangle\subset{\mathbb{P}}^{N}$
+with $a^{(1)},\dots,a^{(\beta)}\in{\mathbb{C}}$ and $a^{(\ell)}=0$ if $\ell>\beta$.
+Then the expression $a=\sum_{i=1}^{k}c_{i}\cdot v_{d}(x_{i}^{\prime})$ means that $a^{(\ell)}\sum_{i=1}^{k}c_{i}\cdot v_{d}(x_{i}^{\prime})^{(0)}-\sum_{i=1}^{k}c_{i}\cdot v_{d}(x_{i}^{\prime})^{(\ell)}\in{I(F_{\xi})}$ for $1\leq\ell\leq N$, where $v_{d}(x_{i}^{\prime})^{(\ell)}$ is the $\ell$-th coordinate of $v_{d}(x_{i}^{\prime})\in{\mathbb{P}}^{N}$.
+In particular, $\sum_{i=1}^{k}c_{i}\cdot v_{d}(x_{i}^{\prime})^{(\ell)}\in{I(F_{\xi})}$ for $\ell>\beta$.
+Using the discussion of (i), we have $x_{i,j}^{\prime}\in I(F_{\xi})$ for all $1\leq i\leq k$ and $j>m$, which means that $I(F_{\xi})$ contains the defining ideal of $({\mathbb{P}}^{k-1}\times(L)^{k})\cap(W\times(V_{0})^{k}_{\xi})$.
+Thus, scheme-theoretically, it follows $F_{\xi}\subset{\mathbb{P}}^{k-1}\times(U_{0}\cap(L)^{k})$ for any $\xi$, and hence $\Phi_{1}^{-1}(a)\subset{\mathbb{P}}^{k-1}\times(U_{0}\cap(L)^{k})$.
+Therefore, $p^{-1}(a)\subset\left\{a\right\}\times(L)^{k}$.
 ∎
 Remark 26.
-We recall some known results on $k$-the secant variety and its incidence in terms of $k$-fold symmetric product of ${\mathbb{P}}^{n}$.
+We recall some known results on the $k$-the secant variety and its incidence in terms of $k$-fold symmetric product of ${\mathbb{P}}^{n}$.
   1. (a)
-    It is known that $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ is non-singular at $(x_{1}^{\prime},x_{2}^{\prime},\dots,x_{k}^{\prime})$ if $x_{i}^{\prime}\neq x_{j}^{\prime}$ whenever $i\neq j$.
+    It is known that $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ is non-singular at $(x_{1}^{\prime},\dots,x_{k}^{\prime})$ if $x_{i}^{\prime}\neq x_{j}^{\prime}$ whenever $i\neq j$.
 Thus, the subset of all distinct $k$-points of ${\mathbb{P}}^{n}$ is a smooth open subscheme of $\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ (see e.g.
-[6, lemma 7.1.4]).
+[BK05, Lemma 7.1.4]).
 Then we also consider the incidence variety in this setting as
-    $\widetilde{I}\subset{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$
-    , where $\widetilde{I}$ corresponds to $I$ in (3) under the natural map ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ and $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\subset{\mathbb{P}}^{N}$ is the first projection.
+    $\widetilde{I}\subset{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n}),$
+    where $\widetilde{I}$ corresponds to $I$ in (4) under the natural map ${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$ and $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\subset{\mathbb{P}}^{N}$ is the first projection.
   2. (b)
     Assume $k(n+1)<\binom{n+d}{n}$.
-Then, from [9, theorem 1.1] we know that the projection $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is birational except $(k,d,n)=(9,6,2),(8,4,3),(9,3,5)$, because it is a dominant and generically injective morphism.
+Then, we know from [COV17, Theorem 1.1] that the projection $\widetilde{p}:\widetilde{I}\rightarrow\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is birational except $(k,d,n)=(9,6,2),(8,4,3),(9,3,5)$, because it is a dominant and generically injective morphism.
 Now, we are ready to prove Part (i) of Theorem 2 and Theorem 3.
 Proof of Part (i) of Theorem 2 and Part (i) of Theorem 3.
-For an $m$-plane $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$, we take the $m$-subsecant variety $Z=\sigma_{k}(v_{d}(L))$ of $Y=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
-By assumptions of Theorems 2 and 3, we know that $\dim Y={nk+k-1}$ and $\dim Z={mk+k-1}\leq\dim\langle Z\rangle=\tbinom{m+d}{m}-1$ (i.e.
-$Z$ is non-defective) by Alexander-Hirschowitz theorem and $Z$ is also identifiable at its general point by [9, theorem 1.1].
-Further, $(k,d,m)\neq(9,6,2),(8,4,3),(9,3,5)$ and note that $k\leq\tbinom{m+d-1}{m}$ for each $k$ of this range.
+For an $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ with $m\geq 2$, we take the $m$-subsecant variety $Z=\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ of $Y=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+From [AH95], for $d\geq 3$, $Z$ does not fill $\langle\,Z\,\rangle$ and is secant defective if and only if $(k,d,m)=(7,3,4),(5,4,2),(9,4,3),(14,4,4)$.
+Thus, by the assumptions of Theorems 2 and 3, we know that
+$\dim Y={nk+k-1},\quad\dim Z={mk+k-1}\leq\dim\langle\,Z\,\rangle=\tbinom{m+d}{m}-1,$
+that is, $Y,Z$ are non-defective.
+In this case, $Z=\langle\,Z\,\rangle$ if and only if $k=\frac{\binom{m+d}{m}}{m+1}\in{\mathbb{N}}$.
+In particular, under the assumption $k<\mu$ of Theorem 2 (i), we have $Z\subsetneq\langle\,Z\,\rangle$.
+If $Z\subsetneq\langle\,Z\,\rangle$, then since $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$ are excluded from Theorem 2 and (i) of Table 1, it follows from [COV17, Theorem 1.1] that $Z$ is generically identifiable.
+If $Z=\langle\,Z\,\rangle$, then $(k,d,m)=(5,3,3),(7,5,2)$ of (i) of Table 1 only occur, and in these cases, it follows from [GM17, Theorem 1] that $Z$ is generically identifiable.
 Let $a\in Z$ be a general point and consider $\widetilde{p}:\widetilde{I}\rightarrow Y$.
-By Lemma 25 and Remark 26 (b), we may assume that the scheme-theoretic fiber $\widetilde{p}^{-1}(a)$ is one point $\mathbf{x}=(a,x_{1}^{\prime},\dots,x_{k}^{\prime})$ in $\widetilde{I}\cap({\mathbb{P}}^{N}\times\operatorname{Sym}^{k}(L))$ and $\mathbf{x}$ is a non-singular point in $\widetilde{I}$, because $\mathbf{x}$ is contained in a smooth Zariski open subset of $\widetilde{I}$ (i.e.
-a projective bundle over a smooth open base, see Remarks 7 (a) and 26 (a)).
-Now, we restrict the projective morphism $\widetilde{p}:\widetilde{I}\rightarrow Y$ onto a non-empty affine open neighborhood $a\in W=\operatorname{Spec}A\subset Y$ and another open subset $\mathbf{x}\in U=\operatorname{Spec}B\subset\widetilde{I}$ and denote the injective ring homomorphism corresponding to $\widetilde{p}|_{U}:U\to W$ by $f:A\hookrightarrow B$ (let also $m_{a}$ (resp.
+Note that $k\leq\tbinom{m+d-1}{m}$ for each $(k,d,m)$ of our range, which is an assumption of Lemma 25.
+Applying Lemma 25, Remark 26 (b), and the generic identifiability of $Z$, we may assume that the scheme-theoretic fiber $\widetilde{p}^{-1}(a)$ is one point $\mathbf{x}=(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in\widetilde{I}\cap({\mathbb{P}}^{N}\times\operatorname{Sym}^{k}(L))$ and $\mathbf{x}$ is a non-singular point in $\widetilde{I}$, because $\mathbf{x}$ is contained in a smooth Zariski open subset of $\widetilde{I}$ (i.e., a projective bundle over a smooth open base, see Remarks 7 (a) and 26 (a)).
+Now, we restrict the projective morphism $\widetilde{p}:\widetilde{I}\rightarrow Y$ onto a non-empty affine open neighborhood $a\in W=\operatorname{Spec}A\subset Y$ and another open subset $\mathbf{x}\in U=\operatorname{Spec}B\subset\widetilde{I}$, and take the injective ring homomorphism $A\hookrightarrow B$ corresponding to $\widetilde{p}|_{U}:U\to W$.
+Also, let $m_{a}$ (resp.
 $m_{\mathbf{x}}$) be the maximal ideal of $a$ in $A$ (resp.
-of $\mathbf{x}$ in $B$)).
+of $\mathbf{x}$ in $B$).
 Note that we may take $U$ so that $\widetilde{p}|_{U}:U\to W$ is a finite morphism (cf.
-[18, chap.
-II, ex.3.22 (d)]).
-Since $A/m_{a}\simeq B/m_{\mathbf{x}}\simeq{\mathbb{C}}$ and $\tilde{p}^{-1}(a)\simeq\operatorname{Spec}(B\otimes_{A}A/m_{a})\simeq\operatorname{Spec}(B/f(m_{a})B)$ is isomorphic to $\operatorname{Spec}{A/m_{a}}$, we have $f(m_{a})=m_{\mathbf{x}}$ in $B$ and $m_{a}B_{m_{a}}=m_{\mathbf{x}}B_{m_{a}}$ in the local ring $B_{m_{a}}$.
-We also know $A_{m_{a}}+m_{a}B_{m_{a}}/m_{a}B_{m_{a}}\simeq A_{m_{a}}/m_{a}A_{m_{a}}\simeq B_{m_{a}}/m_{a}B_{m_{a}}$, thus as $A_{m_{a}}$-module $A_{m_{a}}=B_{m_{a}}$ holds by Nakayama lemma.
-Note that $A_{m_{a}}=B_{m_{a}}$ is also true as a local ring.
-Finally, since we have $m_{a}A_{m_{a}}=m_{a}B_{m_{a}}=m_{\mathbf{x}}B_{m_{a}}$, we obtain the desired local ring isomorphism
-$A_{m_{a}}=B_{m_{a}}\simeq(B_{m_{a}})_{m_{a}B_{m_{a}}}=(B_{m_{a}})_{m_{\mathbf{x}}B_{m_{a}}}\simeq B_{m_{\mathbf{x}}},$
+[H77, Chapter II, Ex.3.22 (d)]).
+Since $A/m_{a}=B/m_{\mathbf{x}}\simeq{\mathbb{C}}$ and
+$\widetilde{p}^{-1}(a)\simeq\operatorname{Spec}(B\otimes_{A}A/m_{a})\simeq\operatorname{Spec}(B/m_{a}B)$
+is isomorphic to one simple point $\operatorname{Spec}{B/m_{\mathbf{x}}}$, we have $m_{a}B=m_{\mathbf{x}}$ in $B$.
+Let $B_{m_{a}}=B\otimes_{A}A_{m_{a}}$, whose member can be expressed as $b/s$ with $b\in B$ and $s\in A\setminus m_{a}$.
+We have $m_{a}B_{m_{a}}=m_{\mathbf{x}}B_{m_{a}}$ in $B_{m_{a}}$, and then,
+$(A_{m_{a}}+m_{a}B_{m_{a}})/m_{a}B_{m_{a}}\simeq A_{m_{a}}/m_{a}A_{m_{a}}\simeq B_{m_{a}}/m_{a}B_{m_{a}},$
+which implies $A_{m_{a}}+m_{a}B_{m_{a}}=B_{m_{a}}+m_{a}B_{m_{a}}=B_{m_{a}}$ as $A_{m_{a}}$-module.
+By Nakayama lemma (see e.g.
+[Matsu, Corollary of Theorem 2.2]), it follows $A_{m_{a}}=B_{m_{a}}$.
+In particular, $B_{m_{a}}$ is a local ring, whose maximal ideal is $m_{\mathbf{x}}B_{m_{a}}$.
+Thus, we have
+$A_{m_{a}}=B_{m_{a}}=(B_{m_{a}})_{m_{\mathbf{x}}B_{m_{a}}}=B_{m_{\mathbf{x}}},$
 which implies that $a$ is a smooth point in $Y$. ∎
 We present an example which shows that one can not extend this generic smoothness result to arbitrary point in the locus $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
-Example 27 (Singularity can happen at a special point in Part (i) of Theorem 3).
-Let $V={\mathbb{C}}\langle x,y,z,w\rangle\supset W={\mathbb{C}}\langle x,y,z\rangle$ and $f=x^{2}y^{2}+z^{4}$ be a form of degree $4$.
+Example 27 (Singularity can occur at a special point in Part (i) of Theorem 3).
+Let
+$V={\mathbb{C}}\langle x,y,z,w\rangle\supset W={\mathbb{C}}\langle x,y,z\rangle$
+and $f=x^{2}y^{2}+z^{4}$ be a form of degree $4$.
 Then, $f$ represents a point in $\sigma_{4}(v_{4}({\mathbb{P}}W))\setminus\sigma_{3}(v_{4}({\mathbb{P}}V))$.
@@
 We know that the form $f_{D}=x^{2}y^{2}$ has Waring rank 3 so that $f_{D}=\ell_{1}^{4}+\ell_{2}^{4}+\ell_{3}^{4}$ for some $\ell_{i}\in{\mathbb{C}}[x,y]_{1}$.
-By [16, theorem 2.1] $f_{D}$ is also a singular point of $\sigma_{3}(v_{4}({\mathbb{P}}V))$.
+By [Han18, Theorem 2.1] $f_{D}$ is also a singular point of $\sigma_{3}(v_{4}({\mathbb{P}}V))$.
 Since $f\in\langle f_{D},z^{4}\rangle$, by Terracini’s lemma, we see that $\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\subset\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))$ and $\mathbb{T}_{\ell_{i}}v_{4}({\mathbb{P}}V)\subset\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))$ for any $i$.
@@
 Thus, we have
-(19) $\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))\supseteq\langle\bigcup_{\ell_{i}\in{\mathbb{P}}^{1}}\mathbb{T}_{\ell_{i}^{4}}v_{4}({\mathbb{P}}V),\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\rangle~{}.$
+(27) $\mathbb{T}_{f}\sigma_{4}(v_{4}({\mathbb{P}}V))\supseteq\langle\bigcup_{\ell_{i}\in{\mathbb{P}}^{1}}\mathbb{T}_{\ell_{i}^{4}}v_{4}({\mathbb{P}}V),\mathbb{T}_{z^{4}}v_{4}({\mathbb{P}}V)\rangle.$
 Note that using the parameterization (16) we can estimate the dimension of right hand side of (27).
-Take an affine open subset $\{[1:t:u_{1}:u_{2}]\}$ of ${\mathbb{P}}^{3}$ and (with a change of coordinates) let $z^{4}$ be $[1:0:1:1]$ and $\ell_{i}\in{\mathbb{P}}^{1}$ be reprsented by $[1:t:0:0]$ for $t\in{\mathbb{C}}$.
+Take an affine open subset $\{[1:t:u_{1}:u_{2}]\}$ of ${\mathbb{P}}^{3}$ and (with a change of coordinates) let $z^{4}$ be $[1:0:1:1]$ and $\ell_{i}\in{\mathbb{P}}^{1}$ be represented by $[1:t:0:0]$ for $t\in{\mathbb{C}}$.
 Then, by (17) the embedded tangent space to $v_{4}({\mathbb{P}}V)$ at $[1:t:u_{1}:u_{2}]$ is given as the row span of
@@
 4.2. Singularity
-In this subsection we will prove Part (ii) and (iii) both in Theorem 2 and Theorem 3, which show singularity of the $m$-subsecant loci $\Sigma_{k,d}(m)$ in the $k$-secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
-Since $\Sigma_{k,d}(m)$ is the union of all the $m$-subsecant varieties $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ as (2), it is enough to prove the statements for $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+In this subsection we will prove parts (ii) and (iii) both in Theorem 2 and Theorem 3, which show the singularity of the $m$-subsecant loci $\Sigma_{k,d}(m)$ in the $k$-th secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+Since $\Sigma_{k,d}(m)$ is the union of all the $m$-subsecant varieties $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ as (3), it is enough to prove the statements for any $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
 Proof of Part (ii) of Theorem 2 and Part (ii) of Theorem 3.
-As we noted above, here it is enough to show that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for each $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
-We will first prove $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ under the condition in Theorem 2 with $(k,d,m)\neq(9,6,2),(9,3,5)$, next for Theorem 3 with $(k,d,m)\neq(8,4,3)$, and finally for the remaining $3$ cases $(k,d,m)=(9,6,2),(9,3,5),(8,4,3)$.
-Note that in the remaining $3$ cases, $(k,d,m)$ satisfy that $\beta=\binom{m+d}{m}-1\not<km+k-1$ and $k=\bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}-1\bigl{\rceil}$.
-Basically, we use the same idea for the proof, but a detailed way of estimation will be slightly different according to each case (due to defectivity and identifiability).
-The non-triviality of the singular locus, i.e.
-$\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ can be directly obtained at the end by Lemma 10.
+As we noted above, it is enough here to show that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for each $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+We will first prove that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ under the condition in Theorem 2 with $\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$, next for Theorem 3 with $(k,d,m)\neq(9,3,5),(8,4,3),(9,6,2)$, and finally for $\frac{\binom{m+d}{m}}{m+1}\in{\mathbb{N}}$ or $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$.
+Basically, we use the same idea for the proof, but a detailed way of estimation will be slightly different according to each case (due to secant defectivity and non-identifiability).
+The non-triviality of the singular locus, i.e., $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ can be directly obtained at the end by Lemma 10.
 Let us take a general point $(a,x_{1},\dots,x_{k})$ in the incidence $J$ as (9) for $Z=v_{d}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta}$ and take an irreducible component $F$ of $p^{-1}(a)$ containing $(a,x_{1},\dots,x_{k})$.
@@
 In particular, we have
-$\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))~{}.$
+$\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
 Terracini’s lemma implies $\mathbb{T}_{x_{i}}v_{d}({\mathbb{P}}^{n})\subset\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ for $i=1,\dots,k$, and Lemma 12 implies $\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ for a general point $x\in q_{i}(F)$.
 Thus, we have
-(20) $\langle\,\;\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\cup\bigcup_{x\in q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\}}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\;\,\rangle\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
-First of all, let us consider (ii) of Theorem 2 with $(k,d,m)\neq(9,6,2),(9,3,5)$.
+(28) $\bigl{\langle}\;\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\,\cup\bigcup_{x\in q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\}}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\;\bigr{\rangle}\,\subset\,\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
+First of all, let us consider (ii) of Theorem 2 with $\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$.
+Set $k=\Bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}\Bigl{\rceil}$.
 Then $\beta=\binom{m+d}{m}-1<km+k-1$ and $(k-1)m+k\leq\binom{m+d}{m}$ as in Remark 16.
-We have ${\mathbb{P}}^{\beta}=\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))$, since $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ fills the whole ${\mathbb{P}}^{\beta}$.
-It is enough to discuss the following cases: (a1) $(k-1)m+k<\binom{m+d}{m}$; (a2) $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)\neq(3,2)$; (a3) $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)=(3,2)$.
-For the case (a1) (i.e., $(k-1)m+k<\binom{m+d}{m}$), by applying Proposition 15 (i) and Proposition 22 to $A=v_{d}^{-1}(q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\})$ in ${\mathbb{P}}^{m}$ and $\Lambda={\mathbb{P}}^{\beta}$, the number (14), which is a lower bound for the dimension of the left hand side in (20), is greater than or equal to $\beta+(n-m)(1+k+(km+k-1)-\beta)$.
-Thus, from the containment (20), we obtain
-$\beta+(n-m)(1+k+(km+k-1)-\beta)\leq kn+k-1~{},$
+We have ${\mathbb{P}}^{\beta}=\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))$ since $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ fills up the whole ${\mathbb{P}}^{\beta}$.
+It is enough to discuss the following three cases
+  1.  (a1)
+    $(k-1)m+k<\binom{m+d}{m}$,
+  2.  (a2)
+    $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)\neq(3,2)$,
+  3.  (a3)
+    $(k-1)m+k=\binom{m+d}{m}$ and $(d,m)=(3,2)$.
+For the case (a1) (i.e., $(k-1)m+k<\binom{m+d}{m}$), we take $A=v_{d}^{-1}(q_{i}(F)\cup\left\{x_{1},\dots,x_{k}\right\})$ in ${\mathbb{P}}^{m}$ and $\Lambda={\mathbb{P}}^{\beta}$.
+From Proposition 15 (i), it follows $\dim\langle\,v_{d-1,m}(A)\,\rangle\geq k+(km+k-1)-\beta$ for the ($d-1$)-uple Veronese embedding $v_{d-1,m}$ of ${\mathbb{P}}^{m}$.
+From Proposition 22, the dimension of the left hand side in (28) is greater than or equal to the number (19), which is
+$\dim\langle\,\Lambda\cup v_{d}(A)\,\rangle+(n-m)\big{\{}1+\dim\langle\,v_{d-1,m}(A)\,\rangle\big{\}}\geq\beta+(n-m)(1+k+(km+k-1)-\beta).$
+From the inclusion (28), we obtain
+$\beta+(n-m)(1+k+(km+k-1)-\beta)\leq kn+k-1,$
 which implies $(n-m)(1+(km+k-1)-\beta)\leq(km+k-1)-\beta$. It is a contradiction, because $n>m$ and $(km+k-1)-\beta>0$.
 Now, assume $(k-1)m+k=\binom{m+d}{m}$ (equivalently, $km+k-1-\beta=m$).
-Then, in the same way as above, using Propositions 15 (ii) and taking $A={\mathbb{P}}^{m}=v_{d}^{-1}(Z)$ and $\Lambda=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle={\mathbb{P}}^{\beta}$, we have
+Then, in the same way as above, using Proposition 15 (ii) and taking $A={\mathbb{P}}^{m}=v_{d}^{-1}(Z)$ and $\Lambda=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle={\mathbb{P}}^{\beta}$, we have
 (29) $\beta+(n-m)(1+\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle)\leq kn+k-1.$
-For $(d,m)\neq(3,2)$ (i.e., the case of (a2)), by Propositions 15 (ii), it holds $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle\geq k+m$.
-Then $(n-m)(m+1)\leq km+k-1-\beta=m$, contrary to $n>m$.
-For $(d,m)=(3,2)$ (i.e., the case (a3)), we get $\beta=9$, $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle=5$, and $k=\lceil\frac{\binom{m+d}{m}}{m+1}\rceil=4$.
+For $(d,m)\neq(3,2)$ (i.e., the case of (a2)), Proposition 15 implies that $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle\geq k+m$. Then $(n-m)(m+1)\leq km+k-1-\beta=m$, contrary to $n>m$.
+For $(d,m)=(3,2)$ (i.e., the case (a3)), we get $\beta=9$, $\dim\langle\,v_{d-1,m}({\mathbb{P}}^{m})\,\rangle=5$, and $k=\lceil{\binom{m+d}{m}}/({m+1})\rceil=4$.
 The condition $(k,d,n)\neq(4,3,3)$ implies $n\geq 4$.
@@
 Secondly, let us regard (ii) of Theorem 3.
-Except $(k,d,m)=(8,4,3)$, this part consists of (b1) $(k,d,m)=(7,3,4),(5,4,2),(9,4,3)$, and $(14,4,4)$ (i.e., the case of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ being defective); (b2) $(k,d,m)=(8,3,4),(6,4,2),(10,4,3)$, and $(15,4,4)$ (i.e., just after defective case).
-By the same reason, we also have the containment (20) for these cases provided that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
-For the defective cases (i.e., the case (b1)), it is known that all the $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s are hypersurfaces in ${\mathbb{P}}^{\beta}$.
-So, as taking $A=v_{d}^{-1}(q_{i}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$ corresponding to the entry locus of $a$, by Proposition 22 a containment of the same kind as (20) implies
-(22) $\beta-1+(n-m)(1+\dim\langle\,v_{d-1}(A)\,\rangle)\leq kn+k-1~{},$
+For $(k,d,m)=(10,3,5),(10,6,2)$, we have the same result as Theorem 2 since $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ and it satisfies (a1),(a2) respectively.
+Then, except $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$, the remaining part of (ii) of Theorem 3 consists of the following two cases
+  1.  (b1)
+    $(k,d,m)=(7,3,4),(5,4,2),(9,4,3),(14,4,4)$ (i.e., the case of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ being defective),
+  2.  (b2)
+    $(k,d,m)=(8,3,4),(6,4,2),(10,4,3),(15,4,4)$ (i.e., just after the defective case).
+By the same reason, we also have the inclusion (28) for these cases provided that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
+For the case (b1), i.e., the defective case, it is known that all the $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s are hypersurfaces in ${\mathbb{P}}^{\beta}$ (see [AH95]).
+So, as taking $A=v_{d}^{-1}(q_{i}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$ corresponding to the entry locus of $a$, by Proposition 22 a inclusion of the same kind as (28) implies
+(30) $\beta-1+(n-m)(1+\dim\langle\,v_{d-1}(A)\,\rangle)\leq kn+k-1,$
 where $\beta$ is equal to $km+k-\delta$ and $\delta$ is the secant defect of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$. (30) is equivalent to
-$n-m\leq\frac{\delta}{1+\dim\langle\,v_{d-1}(A)\,\rangle-k}~{}\bigg{(}\leq\frac{\delta}{1+\delta}<1\bigg{)}~{},$
+$n-m\leq\frac{\delta}{1+\dim\langle\,v_{d-1}(A)\,\rangle-k}\leq\frac{\delta}{1+\delta}<1,$
 which contradicts $n-m\geq 1$, because $\dim\langle\,v_{d-1}(A)\,\rangle\geq k+\delta$ by Remark 21.
-For the case (b2) i.e., just after defective cases $(k,d,m)=(8,3,4),(6,4,2),(10,4,3)$, and $(15,4,4)$, we can also get a contradiction in a similar way.
-Since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))$ is a hypersurface in ${\mathbb{P}}^{\beta}$ in all these cases, by Lemma 14 we have $q_{i}(F)=v_{d}({\mathbb{P}}^{m})$ for any maximal irreducible component $F$ of $p^{-1}(a)$ so that we can take $A={\mathbb{P}}^{m}$.
-Further we see that $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is non-defective for any $n>m$ in each case and $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ so that $\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))={\mathbb{P}}^{\beta}$.
-By Proposition 22, the containment (20) implies
-$\beta+(n-m)(1+\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle)=\binom{m+d}{m}-1+(n-m)\binom{m+d-1}{m}\leq kn+k-1~{},$
-which fails to hold in each 4 just after defective case above, because $n\geq m+1$.
-Thus, we obtain $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
-For the remaining $(k,d,m)=(9,6,2),(9,3,5),(8,4,3)$, we first note that these are the only generic non-identifiability cases of subgeneric rank in which a generic secant fiber $p^{-1}(a)$ consists of not one, but two points (see [9, theorem 1.1]).
-Even though they do not satisfy ($\frac{{m+d\choose m}}{m+1}<k<\frac{{m+d\choose m}}{m+1}+1$ as above (so, not having a positive dimensional secant fibers), we can still get a proof by contradiction using another estimation.
-Similarly, suppose that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and take a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ so that $a$ is a smooth point both in $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
-In each 3 cases of $(k,d,m)$, let $x_{1},\dots,x_{k}$ and $y_{1},\dots,y_{k}$ be the two secant fibers of $a$ (i.e., $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$ and $a\in\langle\,y_{1},\dots,y_{k}\,\rangle$) and let $x_{i}^{\prime}\in{\mathbb{P}}^{m}$ (resp.
-$y_{j}^{\prime}$) be the preimage of $x_{i}$, that is $v_{d}(x_{i}^{\prime})=x_{i}$, (resp.
+For the case (b2), i.e., just after the defective case (b1), the $k$-th secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ fills up ${\mathbb{P}}^{\beta}$, and hence $\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))={\mathbb{P}}^{\beta}$.
+Then, we can also get a contradiction in a similar way, as follows.
+Since the $(k-1)$-secant variety $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))$ is a hypersurface in ${\mathbb{P}}^{\beta}$, by Lemma 14 we have $q_{i}(F)=v_{d}({\mathbb{P}}^{m})$ for an irreducible component $F$ of $p^{-1}(a)$ for general $a\in{\mathbb{P}}^{\beta}$ so that we can take $A={\mathbb{P}}^{m}$.
+By Proposition 22, the inclusion (28) implies
+$\beta+(n-m)(1+\dim\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle)=\binom{m+d}{m}-1+(n-m)\binom{m+d-1}{m}\leq kn+k-1,$
+which fails to hold in (b2); more precisely, for $(k,d,m)=(8,3,4),(6,4,2),(10,4,3),(15,4,4)$, the value $\binom{m+d}{m}-1+(n-m)\binom{m+d-1}{m}-(kn+k-1)$ is equal to
+$7n-33,4n-11,10n-35,20n-85,$
+respectively, which must be greater than 00 because of the condition $n\geq m+1$. Thus, we obtain $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$.
+Now, we discuss the following two cases
+  1.  (c1)
+    $k=\frac{\binom{m+d}{m}}{m+1}\in{\mathbb{N}}$ of Theorem 2; then, since we exclude $(k,d,m)=(5,3,3),(7,5,2)$, a general point $a\in{\mathbb{P}}^{\beta}=\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is not $k$-identifiable and the secant fiber $p^{-1}(a)$ consists of two or more points (see [GM17, Theorem 1]),
+  2.  (c2)
+    $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$ of Theorem 3; then, a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}$ is not $k$-identifiable and $p^{-1}(a)$ consists of two points (see [COV17, Theorem 1.1]).
+In these cases, even though they do not have positive dimensional secant fibers, we can still get a proof by contradiction using a different estimate, as follows.
+Similarly, suppose that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ and take a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ so that $a$ is a smooth point in both $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ and $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+We take general $k$ points $x_{1},\dots,x_{k}\in v_{d}({\mathbb{P}}^{m})$ with $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$.
+By the non-identifiability, we have another set of $k$ points $y_{1},\dots,y_{k}\in v_{d}({\mathbb{P}}^{m})$ with $a\in\langle\,y_{1},\dots,y_{k}\,\rangle$ such that $(a,x_{1},\dots,x_{k})$ and $(a,y_{1},\dots,y_{k})$ are distinct in the secant fiber $p^{-1}(a)\subset I\subset{\mathbb{P}}^{\beta}\times(v_{d}({\mathbb{P}}^{m}))^{k}$ (modulo permutation on $(v_{d}({\mathbb{P}}^{m}))^{k}$).
+Let $x_{i}^{\prime}\in{\mathbb{P}}^{m}$ (resp.
+$y_{j}^{\prime}$) be the preimage of $x_{i}$, that is $v_{d}(x_{i}^{\prime})=x_{i}$ (resp.
 of $y_{j}$ with $v_{d}(y_{j}^{\prime})=y_{j}$).
-Now, let us set $A=\{x_{1}^{\prime},\dots,x_{k}^{\prime},y_{1}^{\prime},\dots,y_{k}^{\prime}\}$.
-Then, similarly we have a containment as before
-(23) $\big{\langle}\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\big{\rangle}\subset\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
-Since $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is general, we may assume that at least one $k$ points set of the secant fiber (say $x_{1}^{\prime},\dots,x_{k}^{\prime}$) are general points in ${\mathbb{P}}^{m}$.
-Then $\langle\,v_{d-1,m}(A)\,\rangle$ contains the $(k-1)$-plane $\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$.
-It follows that $\dim\langle\,v_{d-1,m}(A)\,\rangle\geq k$; otherwise, $\dim\langle\,v_{d-1,m}(A)\,\rangle\leq k-1$ implies $\langle\,v_{d-1,m}(A)\,\rangle=\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$ but by the generalized trisecant lemma ([28, proposition 1.4.3]) $\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle\cap v_{d-1}({\mathbb{P}}^{m})=\{v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\}$ as a scheme, which gives a contradiction.
+Setting $A=\{x_{1}^{\prime},\dots,x_{k}^{\prime},y_{1}^{\prime},\dots,y_{k}^{\prime}\}$, we have an inclusion, similar to (28),
+(31) $\bigl{\langle}\;\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{m})))\,\cup\bigcup_{x\in v_{d}(A)}\mathbb{T}_{x}(v_{d}({\mathbb{P}}^{n}))\;\bigr{\rangle}\,\subset\,\mathbb{T}_{a}(\sigma_{k}(v_{d}({\mathbb{P}}^{n}))).$
+For the $(d-1)$-uple Veronese embedding $v_{d-1}=v_{d-1,m}:{\mathbb{P}}^{m}\hookrightarrow{\mathbb{P}}^{\beta_{d-1}}$ with $\beta_{d-1}=\binom{m+d-1}{m}-1$, we have $\dim\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle=k-1$ since $x_{1}^{\prime},\dots,x_{k}^{\prime}$ are general in ${\mathbb{P}}^{m}$.
+The $(k-1)$-plane $\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$ is contained in $\langle\,v_{d-1,m}(A)\,\rangle$.
+On the other hand, the codimension of $v_{d-1}({\mathbb{P}}^{m})\subset{\mathbb{P}}^{\beta_{d-1}}$, that is $\binom{m+d-1}{m}-1-m$, is greater than or equal to $k$; it follows from (iii) of Lemma 17 in the case (c1), and from explicit calculations in the case (c2).
+Then, we have $\dim\langle\,v_{d-1,m}(A)\,\rangle\geq k$, as follows.
+Otherwise, $\dim\langle\,v_{d-1,m}(A)\,\rangle\leq k-1$ implies $\langle\,v_{d-1,m}(A)\,\rangle=\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle$.
+Since $y_{1}^{\prime},\dots,y_{k}^{\prime}\in A\subset{\mathbb{P}}^{m}$, it follows
+$v_{d-1}(y_{1}^{\prime}),\dots,v_{d-1}(y_{k}^{\prime})\in\langle\,v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\,\rangle\cap v_{d-1}({\mathbb{P}}^{m}),$
+where the right hand side must be $\{v_{d-1}(x_{1}^{\prime}),\ldots,v_{d-1}(x_{k}^{\prime})\}$ because of the generalized trisecant lemma ([Rus16, Proposition 1.4.3]), which gives a contradiction.
 Again by (31) and Proposition 22, we get
@@
 $\geq\dim\mathbb{T}_{a}\sigma_{k}(v_{d}({\mathbb{P}}^{m}))+(n-m)(1+k)=km+k-1+(n-m)(1+k)$
-$=kn+k+(n-m-1)~{},$
+$=kn+k+(n-m-1),$
 which is a contradiction since $n-m-1\geq 0$.
-Thus, it also holds that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ in these 3 generic non-identifiable cases.
+Thus, it also holds that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ in these generic non-identifiable cases.
+Note that for $(k,d,m)=(10,3,5),(9,4,3),(10,6,2)$, i.e., just after the non-identifiable case (c2), the singularity is already shown in the second part of this proof, where $(k,d,m)=(9,4,3)$ is also in the defective case (b1).
+The case $(k,d,m)=(5,3,3),(7,5,2)$, which is excluded from (c1), belongs to (i) of Theorem 3; in this sense, the non-trivial singularity does not appear for $(d,m)=(3,3),(5,2)$.
 Finally, since $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ for the $k$’s of the range in this Part (ii), the $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is a non-trivial singular locus, which means $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$, by Lemma 10.
@@
 Proof of Part (iii) of Theorem 2 and Part (iii) of Theorem 3.
-By the conditions in the Part (iii) of these two Theorems, we see that $k-1\geq\lceil\frac{{m+d\choose m}}{m+1}\rceil$ if $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is never defective or $k-1\geq\lceil\frac{{m+d\choose m}}{m+1}\rceil+1$ else if $(d,m)$ appears in the defective list of Alexander-Hiroschowitz.
+By the conditions in the Part (iii) of these two theorems, we see that
+$k-1\geq\Bigl{\lceil}\tfrac{{m+d\choose m}}{m+1}\Bigr{\rceil}$
+if $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is never defective, or
+$k-1\geq\Bigl{\lceil}\tfrac{{m+d\choose m}}{m+1}\Bigr{\rceil}+1$
+else if $(d,m)\in\{(3,4),(4,2),(4,3),(4,4)\}$, the defective list of Alexander-Hiroschowitz.
 In any case, we have $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle$.
-Hence, $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ and the assertion follows.
-∎
+Hence,
+$\sigma_{k}(v_{d}({\mathbb{P}}^{m}))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$
+and the assertion follows. ∎
 Proof of Part (iv) of Theorem 2.
-It is shown in [15] by giving the defining equations of $\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ explicitly. ∎
+It is shown in [FH24] by giving the defining equations of $\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ explicitly. ∎
 5. Case of 4-th secant variety of Veronese embedding
-In this section we aim to prove Theorem 5 as an investigation of the singular loci of the fourth secant variety (i.e.
-$k=4$) of any Veronese variety.
-This theorem consists of one part dealing with (non-)singularity of points in full-secant loci (i.e.
-$m=3$) and the other part for points in the maximum subsecant loci $\Sigma_{4,d}(\min\{k-1,n\}-1)$.
+In this section we aim to prove Theorem 5 as an investigation of the singular loci of the fourth secant variety (i.e., $k=4$) of any Veronese variety.
+This theorem consists of one part dealing with the (non-)singularity of points in full-secant loci (i.e., $m=3$) and the other part for points in the maximum subsecant loci $\Sigma_{4,d}(\min\{k-1,n\}-1)$.
 So, we will obtain Theorem 5 by proving Theorem 29 (Part (i) of Theorem 5) and Corollary 30 (Part (ii) and (iii) of Theorem 5).
 5.1. Equations by Young flattening
-In [23] another source of equations for secant varieties of Veronese was introduced via so-called Young flattening.
-Here we briefly review construction of a certain type of Young flattening and exploit it for computation of conormal space of a given form.
+In [LO13] another source of equations for secant varieties of Veronese varieties was introduced via the so-called Young flattening.
+Here we briefly review the construction of a certain type of Young flattening and use it to compute the conormal space of a given form.
 Let $V={\mathbb{C}}^{n+1}$ and $d=d_{1}+d_{2}+1$. For $1\leq a\leq n$, we consider a map
@@
 Further, for $k\leq\binom{n+d^{\prime}}{d^{\prime}}$ with $d^{\prime}=min\left\{d_{1},d_{2}\right\}$ it is also immediate to see that $\operatorname{rank}(\mathrm{YF}^{a}_{d_{1},d_{2},n}(f))=k\binom{n}{a}$ for a general $k$ sum of $d$-th power $f=\sum_{i=1}^{k}w_{i}^{d}$.
-Thus, from $k\binom{n}{a}+1$ minors of the matrix $\mathrm{YF}^{a}_{d_{1},d_{2},n}(f)$ we obtain a set of equations for $\sigma_{k}(v_{d}({\mathbb{P}}V))$ for this range of $k$ (for some values of $k,d,d^{\prime},a$, it is known that these minors cut $\sigma_{k}(v_{d}({\mathbb{P}}V))$ as an irreducible component (see [23, theorem 1.2.3])).
+Thus, from $k\binom{n}{a}+1$ minors of the matrix $\mathrm{YF}^{a}_{d_{1},d_{2},n}(f)$ we obtain a set of equations for $\sigma_{k}(v_{d}({\mathbb{P}}V))$ for this range of $k$ (for some values of $k,d,d^{\prime},a$, it is known that these minors cut $\sigma_{k}(v_{d}({\mathbb{P}}V))$ as an irreducible component (see [LO13, Theorem 1.2.3])).
 We can also use this Young flattening to compute conormal space of secant varieties of Veronese.
@@
 Then, we have
-$\hat{N}^{\ast}_{f}\sigma_{k}(v_{d}({\mathbb{P}}V))\supseteq(\ker\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))\cdot(\operatorname{im}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))^{\perp}~{},$
-where the right hand side is to be understood as the image of the multiplication $S^{d_{1}}V^{\ast}\otimes\bigwedge^{a}V\otimes S^{d_{2}}V^{\ast}\otimes\bigwedge^{a+1}V^{\ast}\to S^{d}V^{\ast}$.
+$\hat{N}^{\ast}_{[f]}\sigma_{k}(v_{d}({\mathbb{P}}V))\supseteq(\ker\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))\cdot(\operatorname{im}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f))^{\perp}~{},$
+where the right hand side is to be understood as the image of the multiplication
+$S^{d_{1}}V^{\ast}\otimes\bigwedge^{a}V\otimes S^{d_{2}}V^{\ast}\otimes\bigwedge^{a+1}V^{\ast}\to S^{d}V^{\ast}.$
 Proof.
-This proposition comes directly from the same idea as Proposition 23 by applying it to a linear embedding $S^{d}V\hookrightarrow S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V$.
-Since $\operatorname{rank}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f)=k{n\choose a}$ and, as observed before, $v_{d}({\mathbb{P}}V)$ is contained in $\sigma_{n\choose a}(Seg({\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast})\times{\mathbb{P}}(S^{d_{2}}V\otimes\bigwedge^{a+1}V)))$ in ${\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V)$ and not in the previous secants of the same Segre variety, it is straightforward from the proof of Proposition 23 (i.e.
-$p={n\choose a}$ case).
-∎
+This proposition follows directly from the same idea as Proposition 23 by applying it to a linear embedding
+$S^{d}V\hookrightarrow S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V.$
+Since $\operatorname{rank}\mathrm{YF}_{d_{1},d_{2},n}^{a}(f)=k{n\choose a}$ and, as observed before, $v_{d}({\mathbb{P}}V)$ is contained in
+$\sigma_{n\choose a}(Seg({\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast})\times{\mathbb{P}}(S^{d_{2}}V\otimes\bigwedge^{a+1}V)))\subset{\mathbb{P}}(S^{d_{1}}V\otimes\bigwedge^{a}V^{\ast}\otimes S^{d_{2}}V\otimes\bigwedge^{a+1}V)$
+and not in the previous secants of the same Segre variety, it is straightforward from the proof of Proposition 23 (i.e., the case $p={n\choose a}$). ∎
 5.2. Singularity and non-singularity
@@
 $\pi:\sigma_{4}(v_{d}({\mathbb{P}}^{n}))\setminus\big{(}\Sigma_{4,d}(2;{{\mathbb{P}}^{n}})\cup\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\big{)}\to\operatorname{Gr}(3,{\mathbb{P}}^{n})~{}$
-whose fibers $\pi^{-1}({\mathbb{P}}U)$’s are all isomorphic to $\sigma_{4}(v_{d}({\mathbb{P}}U))\setminus(\Sigma_{4,d}(2;{{\mathbb{P}}U})\cup\sigma_{3}(v_{d}({\mathbb{P}}U)))$, where recall that $\Sigma_{4,d}(2;{{\mathbb{P}}^{n}})\subset\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is the maximum subsecant locus, i.e.
-the union of all $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$’s and $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))$’s in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$.
+whose fibers $\pi^{-1}({\mathbb{P}}U)$’s are all isomorphic to $\sigma_{4}(v_{d}({\mathbb{P}}U))\setminus(\Sigma_{4,d}(2;{{\mathbb{P}}U})\cup\sigma_{3}(v_{d}({\mathbb{P}}U)))$, recalling that $\Sigma_{4,d}(2;{{\mathbb{P}}^{n}})\subset\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is the maximum subsecant locus, i.e., the union of all $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$’s and $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))$’s in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$.
 So, we can reduce the proof of theorem to the case of $n=3$.
-In case of $n=3$, there is a list of normal forms in $\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\setminus(\Sigma_{4,d}(\min\{k-1,n\}-1)\cup\sigma_{3}(v_{d}({\mathbb{P}}^{3})))$ due to Landsberg-Teitler (see [21, theorem 10.9.3.1]) such as
-$f_{1}={x_{0}}^{d}+{x_{1}}^{d}+{x_{2}}^{d}+{x_{3}}^{d}$
-$f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$
-$f_{3}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d-1}{x_{3}}$
-$f_{4}={x_{0}}^{d-2}{x_{1}}^{2}+{x_{0}}^{d-1}x_{2}+{x_{3}}^{d}$
+In case of $n=3$, there is a list of normal forms in $\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\setminus(\Sigma_{4,d}(\min\{k-1,n\}-1)\cup\sigma_{3}(v_{d}({\mathbb{P}}^{3})))$ due to Landsberg-Teitler (see [Land12, Theorem 10.9.3.1]) such as
+  1. (i)
+    $f_{1}={x_{0}}^{d}+{x_{1}}^{d}+{x_{2}}^{d}+{x_{3}}^{d}$,
+  2. (ii)
+    $f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$,
+  3. (iii)
+    $f_{3}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d-1}{x_{3}}$ ,
+  4. (iv)
+    $f_{4}={x_{0}}^{d-2}{x_{1}}^{2}+{x_{0}}^{d-1}x_{2}+{x_{3}}^{d}$,
+  5. (v)
     $f_{5}=x_{0}^{d-3}x_{1}^{3}+x_{0}^{d-2}x_{1}x_{2}+x_{0}^{d-1}x_{3}$
@@
 Hence, $f_{1}$ is a smooth point of $\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$.
-Case (ii) $f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$.
-Say $U={\mathbb{C}}\langle x_{0},x_{1},x_{2},x_{3}\rangle$.
-Consider the Young flattening $\mathrm{YF}_{d-2,1,3}^{1}(f_{2})\in S^{d-2}U\otimes U\otimes U^{\ast}\otimes\wedge^{2}U\simeq\operatorname{Hom}(S^{d-2}U^{\ast}\otimes U,U\otimes\wedge^{2}U)$ defined in (24).
-For simplicity, call this type of Young flattening by $\phi$ throughout the proof.
-Then, $\phi(f_{2})$ is
+Case (ii) $f_{2}={x_{0}}^{d-1}x_{1}+{x_{2}}^{d}+{x_{3}}^{d}$. Say
+$U={\mathbb{C}}\langle x_{0},x_{1},x_{2},x_{3}\rangle.$
+Consider the Young flattening
+$\mathrm{YF}_{d-2,1,3}^{1}(f_{2})\in S^{d-2}U\otimes U\otimes U^{\ast}\otimes\wedge^{2}U\simeq\operatorname{Hom}(S^{d-2}U^{\ast}\otimes U,U\otimes\wedge^{2}U)$
+defined in (32). For simplicity, we will denote this type of Young flattening by $\phi$ throughout the proof. Then, $\phi(f_{2})$ is
 $\alpha x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum_{j=0}^{3}y_{j}\otimes x_{j}\wedge x_{1}\big{)}+\beta x_{0}^{d-3}x_{1}\otimes x_{0}\otimes\big{(}\sum_{j=0}^{3}y_{j}\otimes x_{j}\wedge x_{0}\big{)}+\gamma x_{0}^{d-2}\otimes x_{1}\otimes\big{(}\sum_{j=0}^{3}y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
@@
 Note that as a linear map $S^{d-2}U^{\ast}\otimes U\to U\otimes\wedge^{2}U$, $\operatorname{rank}\phi(x_{0}^{5})=3$ and $\operatorname{rank}\phi(f_{2})=4\cdot 3=12$.
-Thus, by Proposition 28 $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ produces a subspace of $\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$.
-For $d=3$, the expected dimension of $\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ for the smoothness is (${3+3\choose 3}-16=4$ and the corresponding four can be chosen as $y_{0}y_{2}y_{3},y_{1}^{2}y_{2},y_{1}y_{2}y_{3},y_{1}^{2}y_{3}$ in ${S}^{3}U^{\ast}$ which are given by the product of $\{y_{1}\otimes x_{0},y_{2}\otimes x_{2},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{2})\subset U^{\ast}\otimes U$ and
+Thus, by Proposition 28 $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ produces a subspace of $\hat{N}^{\ast}_{[f_{2}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$.
+For $d=3$, the expected dimension of $\hat{N}^{\ast}_{[f_{2}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ for the smoothness is (${3+3\choose 3}-16=4$ and the corresponding four can be chosen as $y_{0}y_{2}y_{3},y_{1}^{2}y_{2},y_{1}y_{2}y_{3},y_{1}^{2}y_{3}$ in ${S}^{3}U^{\ast}$ which are given by the product of $\{y_{1}\otimes x_{0},y_{2}\otimes x_{2},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{2})\subset U^{\ast}\otimes U$ and
 (33) $\{~{}y_{0}\otimes y_{2}\wedge y_{3},~{}y_{1}\otimes y_{1}\wedge y_{2},~{}y_{1}\otimes y_{1}\wedge y_{3},~{}y_{1}\otimes y_{2}\wedge y_{3},$
@@
 in $\operatorname{im}\phi(f_{2})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$. So, $\sigma_{4}$ is non-singular at $f_{2}$.
-For any $d\geq 4$, in $\ker\phi(f_{2})\subset S^{d-2}U^{\ast}\otimes U$, one can find a subspace generated by $\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\}$, where $J=\langle y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{2}y_{3}\rangle$ is an ideal in ${S}^{\bullet}U^{\ast}$.
+For any $d\geq 4$, in $\ker\phi(f_{2})\subset S^{d-2}U^{\ast}\otimes U$, one can find a subspace generated by
+$\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\},$
+where $J=\langle y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{2}y_{3}\rangle$ is an ideal in ${S}^{\bullet}U^{\ast}$.
 Also, in $\operatorname{im}\phi(f_{2})^{\perp}\subset U^{\ast}\otimes\wedge^{2}U^{\ast}$ there exists the same subspace as in (33).
 In this case, our $(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$ contains the subspace of ${S}^{d}U^{\ast}$ generated by
+$\{y_{0}^{2}y_{2}^{2},y_{0}^{2}y_{2}y_{3},y_{0}^{2}y_{3}^{2}\}\cup\{y_{0}y_{1}^{2}y_{2},\ldots,y_{0}y_{1}y_{3}^{2}\}\cup\{y_{0}y_{2}^{2}y_{3},y_{0}y_{2}y_{3}^{2}\}\cup\{y_{1}^{4},\ldots,y_{1}^{2}y_{3}^{2}\}\cup\{y_{1}y_{2}^{2}y_{3},y_{1}y_{2}y_{3}^{2},y_{2}^{2}y_{3}^{2}\}$
+for $d=4$ and by
 $\{y_{0}^{d-2}y_{2}^{2},y_{0}^{d-2}y_{2}y_{3},y_{0}^{d-2}y_{3}^{2}\}\cup\{y_{0}^{d-3}y_{1}^{2}y_{2},y_{0}^{d-3}y_{1}^{2}y_{3},\ldots,y_{0}^{d-3}y_{3}^{3}\}$
@@
 $\cup\{y_{0}y_{1}^{d-1},y_{0}y_{1}^{d-2}y_{2},\ldots,y_{0}y_{1}y_{2}^{d-2}\}\cup\{y_{0}y_{2}^{d-2}y_{3},\ldots,y_{0}y_{2}y_{3}^{d-2}\}$
-$\cup\{y_{1}^{d},\ldots,y_{1}^{2}y_{2}^{d-2},\ldots,y_{1}^{2}y_{3}^{d-2}\}\cup\{y_{1}y_{2}^{d-2}y_{3},\ldots,y_{1}y_{2}y_{3}^{d-2}\}\cup\{y_{2}^{d-2}y_{3},\ldots,y_{2}^{2}y_{3}^{d-2}\}$
-(note that the terms above are listed in the lexicographical order). These monomial generators can be also represented as
-($\left(\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{3}^{d-1}\}\right)$
-($\cup\left(\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{3}^{d-1},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}y_{3}^{d-1},y_{3}^{d}\}\right)~{},$
+$\cup\{y_{1}^{d},\ldots,y_{1}^{2}y_{2}^{d-2},\ldots,y_{1}^{2}y_{3}^{d-2}\}\cup\{y_{1}y_{2}^{d-2}y_{3},\ldots,y_{1}y_{2}y_{3}^{d-2}\}\cup\{y_{2}^{d-2}y_{3}^{2},\ldots,y_{2}^{2}y_{3}^{d-2}\}$
+for any $d>4$ (note that the terms above are listed in the lexicographical order). In both cases, these monomial generators can be also represented as
+($\bigl{(}\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{3}^{d-1}\}\bigr{)}$
+($\cup\bigl{(}\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{3}^{d-1},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}y_{3}^{d-1},y_{3}^{d}\}\bigr{)}~{},$
 which implies that by Proposition 28
-$\dim\hat{N}^{\ast}_{f_{2}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$ $\geq\dim(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}~{},$
-{($\geq\bigg{\{}{d-1+3\choose 3}-10+{d+2\choose 2}-6\bigg{\}}=\binom{d+3}{3}-16~{}.$
+$\dim\hat{N}^{\ast}_{[f_{2}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))$ $\geq\dim(\ker\phi(f_{2}))\cdot(\operatorname{im}\phi(f_{2}))^{\perp}$
+{($\geq\bigg{\{}{d-1+3\choose 3}-10+{d+2\choose 2}-6\bigg{\}}=\binom{d+3}{3}-16.$
 Hence, $f_{2}$ is a smooth point of $\sigma_{4}$.
@@
 (34) $\langle~{}y_{0}\otimes y_{2}\wedge y_{3},~{}y_{1}\otimes y_{1}\wedge y_{2},~{}y_{1}\otimes y_{1}\wedge y_{3},~{}y_{1}\otimes y_{2}\wedge y_{3},$
-$~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{3},~{}y_{3}\otimes y_{1}\wedge y_{3}~{}\rangle~{}$
+$~{}y_{2}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{3},~{}y_{3}\otimes y_{1}\wedge y_{3}~{}\rangle~{}$
 produces a desired 4-dimensional subspace $\langle~{}y_{0}y_{3}^{2},y_{1}^{2}y_{2},y_{1}^{2}y_{3},y_{1}y_{3}^{2}~{}\rangle$ in ${S}^{3}U^{\ast}$, which says that $\sigma_{4}$ is non-singular at $f_{3}$.
-Similarly, for the case of $d\geq 4$, $(\ker\phi(f_{3}))\cdot(\operatorname{im}\phi(f_{3}))^{\perp}$ contains a subspace of $\hat{N}^{\ast}_{f_{3}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\subset{S}^{d}U^{\ast}$ which is generated by
-($\left(\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{2}^{d-2}y_{3}\}\right)$
-($\cup\left(\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{2}^{d-2}y_{3},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}^{d-2}y_{3}^{2},y_{2}^{d-3}y_{3}^{3}\}\right)~{},$
+Similarly, for the case of $d\geq 4$, $(\ker\phi(f_{3}))\cdot(\operatorname{im}\phi(f_{3}))^{\perp}$ contains a subspace of $\hat{N}^{\ast}_{[f_{3}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\subset{S}^{d}U^{\ast}$ which is generated by
+($\bigl{(}\{y_{0}^{d},y_{0}^{d-1}y_{1},\cdots,y_{0}y_{3}^{d-1}\}\setminus\{y_{0}^{d},y_{0}^{d-1}y_{1},y_{0}^{d-1}y_{2},y_{0}^{d-1}y_{3},y_{0}^{d-2}y_{1}^{2},y_{0}^{d-2}y_{1}y_{2},y_{0}^{d-2}y_{1}y_{3},y_{0}^{d-3}y_{1}^{3},y_{0}y_{2}^{d-1},y_{0}y_{2}^{d-2}y_{3}\}\bigr{)}$
+($\cup\bigl{(}\{y_{1}^{d},y_{1}^{d-1}y_{2},\cdots,y_{3}^{d}\}\setminus\{y_{1}y_{2}^{d-1},y_{1}y_{2}^{d-2}y_{3},y_{2}^{d},y_{2}^{d-1}y_{3},y_{2}^{d-2}y_{3}^{2},y_{2}^{d-3}y_{3}^{3}\}\bigr{)},$
 using a subspace $\langle~{}\{F\otimes x_{i}~{}|~{}F\in J_{d-2},~{}i=0,\ldots,3\}~{}\rangle$ in $\ker\phi(f_{3})$, where $J$ is an ideal generated by $\{y_{0}y_{2},y_{0}y_{3},y_{1}^{2},y_{1}y_{2},y_{1}y_{3},y_{3}^{2}\}$ in ${S}^{\bullet}U^{\ast}$, and the same subspace in $\operatorname{im}\phi(f_{3})^{\perp}$ as (34).
-Thus, $\dim\hat{N}^{\ast}_{f_{3}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq{d+3\choose 3}-16$, which means that $f_{3}$ is also smooth.
+Thus, $\dim\hat{N}^{\ast}_{[f_{3}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq{d+3\choose 3}-16$, which means that $f_{3}$ is also smooth.
 Case (iv) $f_{4}={x_{0}}^{d-2}{x_{1}}^{2}+{x_{0}}^{d-1}x_{2}+{x_{3}}^{d}$. For $d=3$, we have
@@
 and $\operatorname{rank}\phi(f_{4})=12$.
-$\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{2}y_{3}+y_{1}^{2}y_{3},y_{1}y_{2}y_{3},y_{2}^{3},y_{2}^{2}y_{3}\rangle$ which can be spanned by $\{y_{2}\otimes x_{0},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{4})\subset U^{\ast}\otimes U$ and
+$\hat{N}^{\ast}_{[f_{4}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{2}y_{3}+y_{1}^{2}y_{3},y_{1}y_{2}y_{3},y_{2}^{3},y_{2}^{2}y_{3}\rangle$ which can be spanned by $\{y_{2}\otimes x_{0},y_{3}\otimes x_{3}\}$ in $\ker\phi(f_{4})\subset U^{\ast}\otimes U$ and
 $\{~{}y_{3}\otimes y_{0}\wedge y_{1},y_{3}\otimes y_{0}\wedge y_{2},-y_{1}\otimes y_{1}\wedge y_{2}+y_{2}\otimes y_{0}\wedge y_{2},-y_{0}\otimes y_{2}\wedge y_{3}+y_{1}\otimes y_{1}\wedge y_{3}\}$
@@
 $+(d-1)x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{2}\big{)}+(d-1)x_{0}^{d-2}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
-$+(d-1)(d-2)x_{0}^{d-3}x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+d(d-1)x_{3}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}~{}.$
+$+(d-1)(d-2)x_{0}^{d-3}x_{2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}+d(d-1)x_{3}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}.$
 In this case, $\operatorname{rank}\phi(f_{4})$ is also $12$ and $\ker\phi(f_{4})$ has a subspace $A_{1}$ which is generated by
@@
 $\{~{}y_{2}\otimes y_{1}\wedge y_{2},~{}y_{2}\otimes y_{1}\wedge y_{3},~{}y_{2}\otimes y_{2}\wedge y_{3},~{}y_{3}\otimes y_{0}\wedge y_{1},~{}y_{3}\otimes y_{0}\wedge y_{2},~{}y_{3}\otimes y_{1}\wedge y_{2},$
-$~{}-(d-1)y_{1}\otimes y_{1}\wedge y_{2}+2y_{2}\otimes y_{0}\wedge y_{2},~{}-2y_{0}\otimes y_{2}\wedge y_{3}+(d-1)y_{1}\otimes y_{1}\wedge y_{3}\}~{}.$
-Then, one can check that $A_{1}\cdot B_{1}$ produces a subspace of $\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{1}$ generated by 19 quartics
-$\{~{}-4\underline{y_{0}^{2}y_{2}^{2}}+(2d+4)y_{0}y_{1}^{2}y_{2},~{}-2\underline{y_{0}^{2}y_{2}y_{3}}+3y_{0}y_{1}^{2}y_{3},~{}\underline{y_{0}^{2}y_{3}^{2}},~{}-2\underline{y_{0}y_{1}y_{2}^{2}}+(d-1)y_{1}^{3}y_{2},~{}-2\underline{y_{0}y_{1}y_{2}y_{3}}+3y_{1}^{3}y_{3},$
-$~{}\underline{y_{0}y_{1}y_{3}^{2}},~{}-2\underline{y_{0}y_{2}^{3}}+(d-1)y_{1}^{2}y_{2}^{2},~{}-2\underline{y_{0}y_{2}^{2}y_{3}}+(d-1)y_{1}^{2}y_{2}y_{3},~{}-2\underline{y_{0}y_{2}y_{3}^{2}}+(d-1)y_{1}^{2}y_{3}^{2},~{}\underline{y_{1}^{3}y_{3}},$
-$~{}\underline{y_{1}^{2}y_{2}^{2}},~{}\underline{y_{1}^{2}y_{2}y_{3}},~{}\underline{y_{1}^{2}y_{3}^{2}},~{}\underline{y_{1}y_{2}^{3}},~{}\underline{y_{1}y_{2}^{2}y_{3}},~{}\underline{y_{1}y_{2}y_{3}^{2}},~{}\underline{y_{2}^{4}},~{}\underline{y_{2}^{3}y_{3}},~{}\underline{y_{2}^{2}y_{3}^{2}}\}$
+$~{}-(d-1)y_{1}\otimes y_{1}\wedge y_{2}+2y_{2}\otimes y_{0}\wedge y_{2},~{}-2y_{0}\otimes y_{2}\wedge y_{3}+(d-1)y_{1}\otimes y_{1}\wedge y_{3}\}.$
+Then, one can check that $A_{1}\cdot B_{1}$ produces a subspace of $\hat{N}^{\ast}_{[f_{4}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{1}$ generated by 19 quartics
+[…]
+$\,-2\underline{y_{0}y_{1}y_{2}y_{3}}+3y_{1}^{3}y_{3},~{}\underline{y_{0}y_{1}y_{3}^{2}},~{}-2\underline{y_{0}y_{2}^{3}}+(d-1)y_{1}^{2}y_{2}^{2},~{}-2\underline{y_{0}y_{2}^{2}y_{3}}+(d-1)y_{1}^{2}y_{2}y_{3},~{}-2\underline{y_{0}y_{2}y_{3}^{2}}+(d-1)y_{1}^{2}y_{3}^{2},$
+$\,~{}\underline{y_{1}^{3}y_{3}},~{}\underline{y_{1}^{2}y_{2}^{2}},~{}\underline{y_{1}^{2}y_{2}y_{3}},~{}\underline{y_{1}^{2}y_{3}^{2}},~{}\underline{y_{1}y_{2}^{3}},~{}\underline{y_{1}y_{2}^{2}y_{3}},~{}\underline{y_{1}y_{2}y_{3}^{2}},~{}\underline{y_{2}^{4}},~{}\underline{y_{2}^{3}y_{3}},~{}\underline{y_{2}^{2}y_{3}^{2}}\ \}$
 (here, the underline means the leading term with respect to the lexicographic order). Say $T={S}^{\bullet}U^{\ast}$. $I_{1}$ has a minimal free resolution as
@@
 $H(I,d)$ $=19{d-4+3\choose 3}-36{d-5+3\choose 3}+22{d-6+3\choose 3}-4{d-7+3\choose 3}$
-=($={d+3\choose 3}-16\quad(d\geq 4)~{}.$
-This implies that (${d+3\choose 3}-16\geq\dim\hat{N}^{\ast}_{f_{4}}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq H(I,d)={d+3\choose 3}-16$, which means that $\sigma_{4}$ is also smooth at $f_{4}$.
+=($={d+3\choose 3}-16\quad(d\geq 4).$
+This implies that
+(${d+3\choose 3}-16\geq\dim\hat{N}^{\ast}_{[f_{4}]}\sigma_{4}(v_{d}({\mathbb{P}}^{3}))\geq H(I,d)={d+3\choose 3}-16,$
+which means that $\sigma_{4}$ is also smooth at $f_{4}$.
 Case (v) The final form $f_{5}=x_{0}^{d-3}x_{1}^{3}+x_{0}^{d-2}x_{1}x_{2}+x_{0}^{d-1}x_{3}$. We begin with $d=3$. We have
@@
 and $\operatorname{rank}\phi(f_{5})=12$.
-The conormal space $\hat{N}^{\ast}_{f_{5}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{3}^{2}+4y_{1}y_{2}y_{3}-24y_{2}^{3},-y_{1}y_{3}^{2}+12y_{2}^{2}y_{3},y_{2}y_{3}^{2},y_{3}^{3}\rangle$ which can be spanned by $\{y_{3}\otimes x_{0},2y_{1}\otimes x_{0}+12y_{2}\otimes x_{1}+y_{3}\otimes x_{2}\}$ in $\ker\phi(f_{5})\subset U^{\ast}\otimes U$ and
+The conormal space $\hat{N}^{\ast}_{[f_{5}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ contains a 4-dimensional subspace corresponding to $\langle-y_{0}y_{3}^{2}+4y_{1}y_{2}y_{3}-24y_{2}^{3},-y_{1}y_{3}^{2}+12y_{2}^{2}y_{3},y_{2}y_{3}^{2},y_{3}^{3}\rangle$ which can be spanned by $\{y_{3}\otimes x_{0},2y_{1}\otimes x_{0}+12y_{2}\otimes x_{1}+y_{3}\otimes x_{2}\}$ in $\ker\phi(f_{5})\subset U^{\ast}\otimes U$ and
 (36) $\{~{}y_{3}\otimes y_{1}\wedge y_{2},-2y_{2}\otimes y_{1}\wedge y_{2}+y_{3}\otimes y_{0}\wedge y_{2},-2y_{1}\otimes y_{2}\wedge y_{3}+y_{3}\otimes y_{0}\wedge y_{3}\}$
@@
 $+x_{0}^{d-2}\otimes x_{2}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{1}\big{)}+(d-1)(d-2)x_{0}^{d-3}x_{3}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
-$+(d-1)x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}+(d-1)x_{0}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}~{}$
+$+(d-1)x_{0}^{d-2}\otimes x_{0}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{3}\big{)}+(d-1)x_{0}^{d-2}\otimes x_{3}\otimes\big{(}\sum y_{j}\otimes x_{j}\wedge x_{0}\big{)}$
 and $\operatorname{rank}\phi(f_{5})$ is also $12$. Now, $\ker\phi(f_{5})$ contains a subspace $A_{2}$ which is generated by
@@
 $~{}-6y_{0}\otimes y_{2}\wedge y_{3}+(d-2)y_{1}\otimes y_{1}\wedge y_{3}-6(d-1)y_{2}\otimes y_{1}\wedge y_{2}~{}\}~{}.$
-Then, one can check that $A_{2}\cdot B_{2}$ produces a subspace of $\hat{N}^{\ast}_{f_{5}}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{2}$ generated by 19 quartics
-$\{~{}36y_{0}^{2}y_{2}^{2}-12(d-2)y_{0}y_{1}^{2}y_{2}+(d-2)^{2}y_{1}^{4},~{}6y_{0}^{2}y_{2}y_{3}-(d-2)y_{0}y_{1}^{2}y_{3}-6(d-1)y_{0}y_{1}y_{2}^{2}+(d-1)(d-2)y_{1}^{3}y_{2},$
-$~{}-y_{0}^{2}y_{3}^{2}+2(d-1)y_{0}y_{1}y_{2}y_{3}-(d-1)^{2}y_{1}^{2}y_{2}^{2},~{}-6y_{0}y_{1}y_{2}y_{3}+(d-2)y_{1}^{3}y_{3},~{}-y_{0}y_{1}y_{3}^{2}+(d-1)y_{1}^{2}y_{2}y_{3},$
-$~{}-6y_{0}y_{2}^{3}+(d-2)y_{1}^{2}y_{2}^{2},~{}y_{0}y_{2}^{2}y_{3}-(d-1)y_{1}y_{2}^{3},~{}y_{0}y_{2}y_{3}^{2}-(d-1)y_{1}y_{2}^{2}y_{3},~{}y_{0}y_{3}^{3}-(d-1)y_{1}y_{2}y_{3}^{2},$
-$~{}(d-2)y_{1}^{2}y_{2}y_{3}-6(d-1)y_{1}y_{2}^{3},~{}y_{1}^{2}y_{3}^{2},~{}y_{1}y_{2}^{2}y_{3},~{}y_{1}y_{2}y_{3}^{2},~{}y_{1}y_{3}^{3},~{}y_{2}^{4},~{}y_{2}^{3}y_{3},y_{2}^{2}y_{3}^{2},~{}y_{2}y_{3}^{3},~{}y_{3}^{4}\}~{}.$
+Then, one can check that $A_{2}\cdot B_{2}$ produces a subspace of $\hat{N}^{\ast}_{[f_{5}]}\sigma_{4}(v_{3}({\mathbb{P}}^{3}))$ in ${S}^{d}U^{\ast}$ which is the degree $d$-part of an ideal $I_{2}$ generated by 19 quartics
+[…]
+$-y_{0}^{2}y_{3}^{2}+2(d-1)y_{0}y_{1}y_{2}y_{3}-(d-1)^{2}y_{1}^{2}y_{2}^{2},~{}-6y_{0}y_{1}y_{2}y_{3}+(d-2)y_{1}^{3}y_{3},~{}-y_{0}y_{1}y_{3}^{2}+(d-1)y_{1}^{2}y_{2}y_{3},$
+$-6y_{0}y_{2}^{3}+(d-2)y_{1}^{2}y_{2}^{2},~{}y_{0}y_{2}^{2}y_{3}-(d-1)y_{1}y_{2}^{3},~{}y_{0}y_{2}y_{3}^{2}-(d-1)y_{1}y_{2}^{2}y_{3},~{}y_{0}y_{3}^{3}-(d-1)y_{1}y_{2}y_{3}^{2},$
+$\,(d-2)y_{1}^{2}y_{2}y_{3}-6(d-1)y_{1}y_{2}^{3},~{}y_{1}^{2}y_{3}^{2},~{}y_{1}y_{2}^{2}y_{3},~{}y_{1}y_{2}y_{3}^{2},~{}y_{1}y_{3}^{3},~{}y_{2}^{4},~{}y_{2}^{3}y_{3},y_{2}^{2}y_{3}^{2},~{}y_{2}y_{3}^{3},~{}y_{3}^{4}\ \}.$
 Note that $I_{2}$ has the same minimal free resolution as $I_{1}$ in (35). Therefore, by the same argument, we conclude that $f_{5}$ is also a smooth point when $d\geq 4$. ∎
-As a direct consequence of main results in the paper, we also obtain the following corollary on (non-)singularity of subsecant loci in the 4-th secant variety.
+As a direct consequence of the main results in the paper, we also obtain the following corollary on the (non-)singularity of subsecant loci in the 4-th secant variety.
 Corollary 30 (from subsecant loci).
@@
   2. (ii)
-    $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\subset\operatorname{Sm}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$ if $d\geq 7$.
-$\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$ is non-trivial singular locus when $d=6$ and $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ when $d\leq 5$.
+    $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is smooth at each point in $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ if $d\geq 7$.
+$\sigma_{4}(v_{d}({\mathbb{P}}^{1}))$ is non-trivial singular locus when $d=6$ and $\sigma_{4}(v_{d}({\mathbb{P}}^{1}))\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ in case of $d\leq 5$.
 Proof.
 Since $k=4~{}\&~{}n\geq 3$, the relevant range for an $m$-subsecant locus in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ is $1\leq m\leq 2$.
-(i) For $m=2$, Theorem 2 (ii) tells $\sigma_{4}(v_{d}({\mathbb{P}}^{m}))$ is non-trivial singular locus in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ if $d=3,n\geq 4$.
+(i) For $m=2$, Theorem 2 (ii) says that $\sigma_{4}(v_{d}({\mathbb{P}}^{m}))$ is a non-trivial singular locus in $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ if $d=3,n\geq 4$.
 The case $(d,n)=(3,3)$ is also discussed in Theorem 2 (iv).
-When $d=4$, we can say that a general point in $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is smooth by Theorem 3 (i).
-For any $d\geq 5$, the same conclusion comes from Theorem 2 (i).
+When $d=4,5$ and $6$, we can say that a general point in $\sigma_{4}(v_{d}({\mathbb{P}}^{2}))\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is smooth by Theorem 3 (i).
+For any $d\geq 7$, the same conclusion follows from Theorem 2 (i).
 (ii) This is given by Theorem 1 for the case $k=4,m=1$. ∎
@@
 Corollary 32.
-Let $V$ be a $(n+1)$-dimensional complex vector space $(n\geq 1)$ and $v_{d}({\mathbb{P}}V)\subset{\mathbb{P}}^{N}$ be the image of the $d$-uple $(d\geq 2)$ Veronese embedding of ${\mathbb{P}}V$.
+Let $V$ be an $(n+1)$-dimensional complex vector space $(n\geq 1)$ and $v_{d}({\mathbb{P}}V)\subset{\mathbb{P}}^{N}$ be the image of the $d$-uple $(d\geq 2)$ Veronese embedding of ${\mathbb{P}}V$.
 Assume that $(k,d,n)$ satisfies one of the following conditions:
   1. (i)
-    $d=2$, $\forall~{}n\geq k-1$,
+    $d=2$ and $n\geq k-1$,
   2. (ii)
-    $k=2$, $\forall~{}d,n$,
+    $k=2$, $d\geq 2$ and $n\geq 1$,
   3. (iii)
-    $k=3$, $d=3$ and $\forall~{}n\geq 2$ or $k=3$, $d=4$ and $\forall~{}n\geq 3$,
+    $k=3$, $d=3$ and $n\geq 2$, or $k=3$, $d=4$ and $n\geq 3$,
   4. (iv)
-    $k=4$, $d=3$ and $\forall~{}n\geq 4$.
+    $k=4$, $d=3$ and $n\geq 4$
 Then the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}V))$ is given exactly as
-(29) $\big{\{}f\in{\mathbb{P}}S^{d}V~{}|~{}\textrm{$f$ is any form which can be expressed using at most $k-1$ variables}\big{\}}~{},$
-which is an irreducible locus of dimension $(k-1)(n-k+2)+{d+k-2\choose d}-1$ and is equal to the maximum subsecant locus $\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$.
+(37) $\big{\{}\,f\in{\mathbb{P}}S^{d}V~{}|~{}\textrm{$f$ is any form which can be expressed using at most $k-1$ variables}\,\big{\}},$
+which is an irreducible locus of dimension
+$(k-1)(n-k+2)+{d+k-2\choose d}-1$
+and is equal to the maximum subsecant locus $\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$.
 Proof.
 For case (i), the assertion is immediate since it corresponds to symmetric matrices.
-In case (ii), we draw the conclusion from the fact $\operatorname{Sing}(\sigma_{2}(v_{d}({\mathbb{P}}V)))=v_{d}({\mathbb{P}}V)$ for every $d,n$ [20].
+In case (ii), we draw the conclusion from the fact that $\operatorname{Sing}(\sigma_{2}(v_{d}({\mathbb{P}}V)))=v_{d}({\mathbb{P}}V)$ for every $d,n$ [Kan99].
 For the remaining cases, we first claim that for any $13\leq k\leq n+$ 1 it holds
-(30) $\sigma_{k-1}(v_{d}({\mathbb{P}}V))\subset\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle~{}.$
-We note that the right hand side of (30) is an irreducible and closed subvariety of $\sigma_{k}(v_{d}({\mathbb{P}}V))$, since it coincides with a subvariety $\bigcup_{\Lambda\in\mathrm{Im}\Phi}\Lambda$ where a map $\Phi:\mathbb{G}(k-2,n)\to\mathbb{G}({d+k-2\choose d}-1,N)$ sending each subspace $L$ of dimension $k-2$ to the linear span $\langle\,v_{d}(L)\,\rangle$ in ${\mathbb{P}}^{N}$ is regular (see e.g.
-[17, example 6.10, proposition 6.13]).
+(38) $\sigma_{k-1}(v_{d}({\mathbb{P}}V))\subset\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle.$
+We note that the right hand side of (38) is an irreducible and closed subvariety of $\sigma_{k}(v_{d}({\mathbb{P}}V))$, since it coincides with a subvariety $\bigcup_{\Lambda\in\mathrm{Im}\Phi}\Lambda$, where a map
+$\Phi:\mathbb{G}(k-2,n)\to\mathbb{G}\left(\binom{d+k-2}{d}-1,N\right)$
+sending each subspace $L$ of dimension $k-2$ to the linear span $\langle\,v_{d}(L)\,\rangle$ in ${\mathbb{P}}^{N}$ is regular (see e.g.
+[Harr92, Example 6.10, Proposition 6.13]).
 Then, because a general element of the left hand side is of the form $\ell_{1}^{d}+\cdots+\ell_{k-1}^{d}$ for some linear forms $\ell_{i}$’s, it belongs to $\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$ for some ${\mathbb{P}}^{k-2}\subset{\mathbb{P}}V$ so that the closure is also contained in the subvariety $\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$.
-For case (iii), by [16, theorem 2.1, remark 2.4(a), and corollary 2.11] and Theorem 1, and for case (iv), by Theorem 5, we know that
-$\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}V)))=\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))\big{\}}~{},$
-which can be also written as $\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$.
-In both case (iii) and (iv), we have $\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))=\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$.
-Thus, by the claim above, the singular locus is equal to $\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle=\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V)$, which is irreducible and can be described as written in the statement.
-The formula for dimension is immediate from dimension counting.
-∎
+For case (iii), by [Han18, Theorem 2.1, Remark 2.4(a), and Corollary 2.11] and Theorem 1, and for case (iv), by Theorem 5, we know that
+$\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}V)))=\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))\big{\}},$
+which can also be written as
+$\sigma_{k-1}(v_{d}({\mathbb{P}}V))\cup\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V).$
+In both cases (iii) and (iv), we have $\sigma_{k}(v_{d}({\mathbb{P}}^{k-2}))=\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle$. Thus, by the above claim, the singular locus is equal to
+$\bigcup_{{\mathbb{P}}^{k-2}\subset{\mathbb{P}}V}\langle\,v_{d}({\mathbb{P}}^{k-2})\,\rangle=\Sigma_{k,d}(\min\{k-1,n\}-1;{\mathbb{P}}V),$
+which is irreducible and can be described as written in the statement. The formula for the dimension is immediate from dimension counting. ∎
 6. Concluding remark
-So far, we reported results on singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ coming from the subsecant loci.
-To the best of our knowledge, there has been no general idea or clear consensus on the singular locus of arbitrary higher secant variety of any Veronese variety yet.
-From this point of view, the present paper contributes to it as making a more visible picture on the singular locus via showing a generic smoothness of the subsecant loci for relatively low $k$’s and confirming singularity of the same loci for other $k$’s.
+So far, we have reported results on singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ coming from the subsecant loci.
+To the best of our knowledge, there is no general idea or clear consensus on the singular locus of an arbitrary higher secant variety of any Veronese variety yet.
+From this point of view, the present paper contributes by providing a more visible picture on the singular locus via showing a generic smoothness of the subsecant loci for relatively low $k$’s and confirming the singularity of the same loci for other $k$’s.
 As we mentioned in the introduction, each point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ is located in $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ for some $1\leq m\leq\min\{k-1,n\}$.
-To make the picture more complete, we have two future issues: (i) On the subsecant loci (i.e.
-$m<\min\{k-1,n\}$), one needs to check (non-)singularity not only at a general point but also at every point and (ii) Points in the full-secant locus (i.e.
-$m=\min\{k-1,n\}$) should be treated.
-Issue (i) is expected to be very complicated, because at some special point a singularity can also happen even for a low $k$ as shown in Example 27 (in fact, we can generate more examples using a similar idea).
+To make the picture more complete, we have two future issues: (i) On the subsecant loci (i.e., $m<\min\{k-1,n\}$), one needs to check the (non-)singularity not only at a general point but also at every point and (ii) Points in the full-secant locus (i.e., $m=\min\{k-1,n\}$) should be treated.
+Issue (i) is expected to be very complicated, because at some special point a singularity can also occur even for a low $k$ as shown in Example 27 (in fact, we can generate more examples using a similar idea).
 For the points in the subsecant loci, in general, one could not hope to find some nice ‘normal forms’ and the situation is expected to be wild (in other words, the subsecant loci may not be covered with finitely many nice families of $\mathbf{SL}$-orbits).
 But, still we can push on our viewpoint a bit further and, along the same spirit, we can refine a main result of this paper in the following manner.
-Based on the singularity results in Theorems 2 and 3 and using a similar estimation as §2.3, more generally we have:
+Based on the singularity results in Theorems 1, 2, and 3 and using a similar estimation as in §2.3, more generally we have:
 Theorem 33.
-Suppose that $k>\mu$ in case of Theorem 2 or $k\geq\lfloor\mu\rfloor$ in case of Theorem 3 (in other words, the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is a singular locus in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$).
-Let $r\leq n-m$.
-Then, the following does hold
-(31) $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))\subset\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n})))~{},$
-where $J(X,Y)$ denotes the ‘(embedded) join’ of two subvarieties $X,Y$ in some ${\mathbb{P}}^{N}$.
+Suppose that $m=1$ and $k,d$ satisfy the condition of (ii) or (iii) of Theorem 1, or suppose that $k,d,m$ satisfy the condition of (ii) or (iii) of Theorem 2 or Theorem 3; in other words, the $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is a singular locus in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$.
+Let $11\leq m\leq n-$ 1 and $r\leq n-m$.
+Then, unless $\sigma_{k+r}(v_{d}(\mathbb{P}^{n}))$ fills up the ambient space $\mathbb{P}^{N}$, the following holds
+(39) $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))\subset\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))),$
+where $J(X,Y)$ denotes the ‘(embedded) join’ of two subvarieties $X,Y$ in their ambient space.
 Proof.
@@
 Then, taking $x_{1},\dots,x_{k}$ to be general points of $v_{d}({\mathbb{P}}^{m})$ and $x_{k+1},\dots,x_{k+r}$ to be general points of $v_{d}({\mathbb{P}}^{n})$, we may assume that $x\notin\operatorname{Sing}(\sigma_{k+r}(v_{d}({\mathbb{P}}^{n})))$ for a general $x\in\langle\,x_{1},\dots,x_{k+r}\,\rangle$.
-By Terracini lemma, we have
-(32) $L_{1}:=\langle\,\mathbb{T}_{x_{1}}v_{d}({\mathbb{P}}^{n}),\dots,\mathbb{T}_{x_{k}}v_{d}({\mathbb{P}}^{n})\,\rangle\subset\mathbb{T}_{x}\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$
-and by the assumption on $k$ we know $\dim L_{1}>kn+k-1$.
+By Terracini’s lemma, we have
+(40) $L_{1}=\langle\,\mathbb{T}_{x_{1}}v_{d}({\mathbb{P}}^{n}),\dots,\mathbb{T}_{x_{k}}v_{d}({\mathbb{P}}^{n})\,\rangle\subset\mathbb{T}_{x}\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$
+and by the assumption on $k$ we know that $\dim L_{1}>kn+k-1$.
 On the other hand, since $x_{k+1},\dots,x_{k+r}$ are general points of $v_{d}({\mathbb{P}}^{n})$,
@@
 Taking ${\mathbb{P}}^{n-m-1}\subset{\mathbb{P}}^{n}$ such that $x_{k+1},\dots,x_{k+r}\in{\mathbb{P}}^{r-1}\subset{\mathbb{P}}^{n-m-1}$ and ${\mathbb{P}}^{m}\cap{\mathbb{P}}^{n-m-1}=\emptyset$, and changing coordinates $t_{0},\dots,t_{m},u_{1},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ as in §2.3, we may say that ${\mathbb{P}}^{n-m-1}$ is the zero set of $t_{0}=\dots=t_{m}=0$ and ${\mathbb{P}}^{m}$ is the zero set of $u_{1}=\dots=u_{m^{\prime}}=0$.
-For a point $x^{\prime}\in{\mathbb{P}}^{m}$, using the parametrization (11), the tangent space $\mathbb{T}_{v_{d}(x^{\prime})}v_{d}({\mathbb{P}}^{n})$ is spanned by the rows of the matrix of the form $[\ast:\mathsf{O}]$ as (13).
+For a point $x^{\prime}\in{\mathbb{P}}^{m}$, using the parameterization (16), the tangent space $\mathbb{T}_{v_{d}(x^{\prime})}v_{d}({\mathbb{P}}^{n})$ is spanned by the rows of the matrix of the form $[\ast:\mathsf{O}]$ as (18).
 On the other hand, for a point $x^{\prime\prime}\in{\mathbb{P}}^{n-m-1}$ and for an affine open set containing $x^{\prime\prime}$ we may take $u_{m^{\prime}}=1$ instead of $t_{0}=1$.
-Then, the only part in the parametrization of $v_{d}$ which contributes $\mathbb{T}_{v_{d}(x^{\prime\prime})}v_{d}({\mathbb{P}}^{n})$ is
+Then, the only part on the parameterization of $v_{d}$ which contributes $\mathbb{T}_{v_{d}(x^{\prime\prime})}v_{d}({\mathbb{P}}^{n})$ is
 $\ t_{0}\cdot\mathsf{mono}[u]_{\leq d-1},\ \dots,\ t_{m}\cdot\mathsf{mono}[u]_{\leq d-1},\mathsf{mono}[u]_{\leq d}$
@@
 This new singular locus $J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{n})))$ in (39) can be seen as a ‘partial version’ of subsecant locus in this paper.
-In particular, it contains $m$-subsecant variety $\sigma_{k+r}(v_{d}({\mathbb{P}}^{m}))=J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{m})))$.
+In particular, it contains the $m$-subsecant variety $\sigma_{k+r}(v_{d}({\mathbb{P}}^{m}))=J(\sigma_{k}(v_{d}({\mathbb{P}}^{m})),\sigma_{r}(v_{d}({\mathbb{P}}^{m})))$.
 So, let us call such a locus a partial subsecant locus of $\sigma_{k+r}(v_{d}({\mathbb{P}}^{n}))$.
@@
 Are the points of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\big{(}\mathcal{D}\cup\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))\big{)}$ all smooth in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$?
-Note that the answer to Question 35 is affirmative in cases of $k=2$ classically and $k=3$ (by [16]) and $k=4$ (by Theorem 29).
+Note that the answer to Question 35 is affirmative in cases of $k=2$ classically and $k=3$ (by [Han18]) and $k=4$ (by Theorem 29).
 For a large value $k$ compared to $n$ (e.g.
@@
 Take general 14 points on $v_{8}({\mathbb{P}}^{2})$.
-In [3, remark 4.10], the authors presented a concrete point in the linear span of the 14 points which is a non-normal point to $\sigma_{14}(v_{8}({\mathbb{P}}^{2}))$.
+In [AC20, remark 4.10], the authors presented a concrete point in the linear span of the 14 points which is a non-normal point to $\sigma_{14}(v_{8}({\mathbb{P}}^{2}))$.
 Note that one can also check this singular point does not belong to $\mathcal{D}$, the locus of all partial subsecants.
@@
 For instance, we can have a conjectural result like
-“ For $\vec{\mathbf{n}}:=(n_{1},n_{2},\ldots,n_{r})$, let $X$ be the Segre embedding ${\mathbb{P}}^{n_{1}}\times{\mathbb{P}}^{n_{2}}\times\cdots\times{\mathbb{P}}^{n_{r}}\subset{\mathbb{P}}^{\prod(n_{i}+1)-1}=:{\mathbb{P}}^{\beta(\vec{\mathbf{n}})}$ and denote $\sigma_{k}(X)$ by $\sigma_{k}(\vec{\mathbf{n}})$, $\dim\sigma_{k}(X)$ by $s_{k}(\vec{\mathbf{n}})$.
+“ For $\vec{\mathbf{n}}=(n_{1},n_{2},\ldots,n_{r})$, let $X$ be the Segre embedding ${\mathbb{P}}^{n_{1}}\times{\mathbb{P}}^{n_{2}}\times\cdots\times{\mathbb{P}}^{n_{r}}\subset{\mathbb{P}}^{\prod(n_{i}+1)-1}=:{\mathbb{P}}^{\beta(\vec{\mathbf{n}})}$ and denote $\sigma_{k}(X)$ by $\sigma_{k}(\vec{\mathbf{n}})$, the expected dimension of $\sigma_{k}(\vec{\mathbf{n}})$ by $s_{k}(\vec{\mathbf{n}})$.
 For every $\vec{\mathbf{m}}=(m_{1},m_{2},\ldots,m_{r})$ with $0\leq m_{i}\leq\min\{k-1,n_{i}-1\}$, besides a few exceptional cases, we have that the following holds
@@
     $\sigma_{k}(\vec{\mathbf{m}})\subset\sigma_{k-1}(\vec{\mathbf{n}})$ if $\beta(\vec{\mathbf{m}})\leq s_{k-1}(\vec{\mathbf{m}})$ ” ,
-which can recover the result on the singular locus of the secant varieties of Segre embeddings ([27, corollary 7.17]) for $k=2$.
+which can recover the result on the singular locus of the secant varieties of Segre embeddings ([MOZ15, Corollary 7.17]) for $k=2$.
 Note that, if we assume that everything is non-defective, then the ranges above can be computed as